### v1

# Linear Discriminant Regularized Regression

Abstract
Linear Discriminant Analysis (LDA) is an important classification approach. Its simple linear form makes it easy to interpret and it is capable to handle multi-class responses. It is closely related to other classical multivariate statistical techniques, such as Fisher’s discriminant analysis, canonical correlation analysis and linear regression. In this paper we strengthen its connection to multivariate response regression by characterizing the explicit relationship between the discriminant directions and the regression coefficient matrix. This key characterization leads to a new regression-based multi-class classification procedure that is flexible enough to deploy any existing structured, regularized, and even non-parametric, regression methods. Moreover, our new formulation is generically easy to analyze compared to existing regression-based LDA procedures. In particular, we provide complete theoretical guarantees for using the widely used $\ell_{1}$-regularization that has not yet been fully analyzed in the LDA context. Our theoretical findings are corroborated by extensive simulation studies and real data analysis.
Keywords: Multi-class classification, discriminant analysis, high-dimensional data, regularization, multivariate response regression, dimension reduction.
1 Introduction
This paper introduces a new regression-based approach to multi-class, high-dimensional linear discriminant analysis (LDA). The main premise that leads to LDA is that we have a sample of $n$ i.i.d. copies $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ of $(X,Y)$, with the label $Y\in\{{\bm{e}}_{1},\ldots,{\bm{e}}_{L}\}$ taking values in $L\geq 2$ classes and the conditional distributions of the feature vector $X\in\mathbb{R}^{p}$ satisfy
$X\mid Y={\bm{e}}_{\ell}~{}\sim~{}\mathcal{N}_{p}\left(\mu_{\ell},\Sigma_{w}\right),\qquad\text{for all }\ell\in[L]:=\{1,\ldots,L\}.$ (1.1)
Here $\{{\bm{e}}_{1},\ldots,{\bm{e}}_{L}\}$ are the canonical basis vectors in $\mathbb{R}^{L}$ and the within-class covariance matrix $\Sigma_{w}$ is assumed to be strictly positive definite. For any $i\in[n]$, we say that observation $X_{i}$ belongs to class $\ell\in[L]$ if its label $Y_{i}={\bm{e}}_{\ell}$, which one may simply write as $Y_{i}=\ell$. Our insistence on the perhaps unusual notation ${\bm{e}}_{\ell}$ will become transparent later when we formulate the estimation of the optimal hyperplane as a multivariate response regression problem. Since we can always subtract the marginal mean of $X$, we assume
$\mathbb{E}(X)=0_{p}$ (1.2)
for simplicity, and furthermore that the class probabilities
$\pi_{\ell}=\mathbb{P}\{Y={\bm{e}}_{\ell}\}$ (1.3)
are strictly positive for all $\ell\in[L]$. The Bayes rule under model (1.1) and (1.3) is to classify a new feature $x\in\mathbb{R}^{p}$ as
$g^{*}(x):=\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (1.4)
See, for instance, Izenman (2008). Estimation of this classifier based on $(X_{i},Y_{i})$ with $i\in[n]$ becomes challenging in (i) the high-dimensional setting $p>n$ and (ii) the multi-class setting $L>2$ that allows for $L\to\infty$ as $n\to\infty$.
Few approaches consider (ii) but notable exceptions are Abramovich and Pensky (2019); Nibbering and Hastie (2022). We recall various approaches to deal with (i). Write
$\pi=(\pi_{1},\ldots,\pi_{L})^{\top},\qquad D_{\pi}={\rm diag}(\pi),\qquad{\mathbb{M}}=(\mu_{1},\ldots,\mu_{L})\in\mathbb{R}^{p\times L}$ (1.5)
and denote the unconditional covariance matrix of $X$ by
$\Sigma={\rm Cov}(X)=\Sigma_{w}+{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}.$
We call $\Sigma_{w}$ the within-class covariance matrix and $\Sigma_{b}:=\Sigma-\Sigma_{w}={\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}$ the between-class covariance matrix. The latter due to (1.2) has rank less than $L$.
In high-dimensional settings $p>n$, Cai and Liu (2011) assume that the direction $\Sigma_{w}^{-1}(\mu_{2}-\mu_{1})$ is sparse in the binary case $L=2$, and propose a Dantzig-type procedure to estimate this direction (see, also, Cai and Zhang (2019) for extension to $L>2$). It isn’t always reasonable to assume that such directions are sparse and the procedure is less attractive compared to regression-based approaches discussed later. Witten and Tibshirani (2011) use the fact that the Bayes rule (1.4) is equivalent with Fisher’s discriminant rule
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}F{F^{\top}}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (1.6)
Here the columns of the matrix $F=(F_{1},\ldots,F_{K})\in\mathbb{R}^{p\times K}$, for some $K<L$, are implicitly defined via, for $1\leq k\leq K$,
$\begin{split}F_{k}:=\operatorname*{arg\,max}_{\beta\in\mathbb{R}^{p}}&~{}\beta^{\top}\Sigma_{b}\beta\\ \text{such that}&\quad\beta^{\top}\Sigma_{w}\beta=1,\quad\beta^{\top}\Sigma_{w}F_{i}=0,~{}~{}\forall~{}i<k.\end{split}$ (1.7)
Witten and Tibshirani (2011) develop methodology tailored to the high-dimensional scenarios with a diagonal matrix $\Sigma_{w}$ and sparse columns of $F$. However, the implicit formulation (1.7) of the matrix $F$ makes it extremely difficult to obtain the global solution and to analyze the theoretical properties of the regularized estimator, hence the resulting classifier. More closely related to our approach, introduced in the next paragraph, is optimal scoring, which dates back to De Leeuw et al. (1976); Young et al. (1978) and regains attention in the LDA setting, see Hastie et al. (1995, 1994); Clemmensen et al. (2011); Gaynanova (2020) and the references therein.
By writing the response matrix and feature matrix as
${\mathbb{Y}}=(Y_{1},\ldots,Y_{n})^{\top}\in\{0,1\}^{n\times L},\quad{\mathbb{X}}=(X_{1},\ldots,X_{n})^{\top}\in\mathbb{R}^{n\times p},$ (1.8)
optimal scoring is a regression-based approach that aims to solve a sequence of optimization problems: for $1\leq k\leq K<L$,
$\begin{split}\min_{\ \theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbb{Y}}\theta-{\mathbb{X}}\beta\|_{2}^{2}\\ \text{such that }&~{}\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\theta=n,\quad\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k.\end{split}$ (1.9)
Here $\widehat{\theta}_{i}$ is the solution to (1.9) in the $i$th iteration. The minimization problem (1.9) can be done via solving a generalized eigenvalue problem, see Hastie et al. (1995). Moreover, Hastie et al. (1995) also show that the solutions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$ from (1.9) can be computed from canonical correlation analysis (CCA), and are further related, via CCA, with the matrix $F$ obtained from the sample analogue of (1.7) in the low-dimensional case ($p<n$). As a result, classification can be based on the dimension reduction directions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$.
Since optimal scoring is a regression-based approach, as pointed out in Hastie et al. (1995, 1994), it is often easier to compute and more amenable to extensions to regularized and non-parametric regression as well as model selection than canonical correlation analysis, Fisher’s discriminant rule and the procedure of Cai and Zhang (2019).
In the high-dimensional setting, Hastie et al. (1995) show that the global solution to a penalized version of (1.9), that is, for $1\leq k\leq K<L$,
$\begin{split}\min_{\theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbb{Y}}\theta-{\mathbb{X}}\beta\|_{2}^{2}+\beta^{\top}\Omega\beta\\ \text{such that}&~{}\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\theta=n,\quad\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k,\end{split}$ (1.10)
can be computed for any $p\times p$ symmetric, positive definite matrix $\Omega$. A possible choice of $\Omega=\lambda\mathbb{I}_{p}$ for some $\lambda>0$ leads to the ridge penalty (Campbell, 1980; Friedman, 1989). In a more recent work (Gaynanova, 2020), the matrix formulation of (1.10)
$\begin{split}\min_{\Theta,~{}B}&~{}\|{\mathbb{Y}}\Theta-{\mathbb{X}}B\|_{F}^{2}+\text{pen}(B)\\ \text{such that }&~{}\Theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\Theta=n\mathbb{I}_{K},\quad\Theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}1_{n}=0_{K}.\end{split}$ (1.11)
is considered, and it is proved that the global solution $\widehat{B}\in\mathbb{R}^{p\times K}$ to (1.11) can be computed for any penalty $\text{pen}(B)$ that satisfies $\text{pen}(BR)=\text{pen}(B)$ for any orthogonal $K\times K$ matrix $R$. This allows for the penalty $\|B\|_{1,2}=\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ used in the group-lasso, but excludes the familiar $\|B\|_{1}=\sum_{j,k}|B_{jk}|$ used in the lasso regression as well as its extensions such as the elastic net penalty. That being said, for penalties that are not orthogonal invariant, there is no guarantee of computing the global solution to (1.11), let alone any theoretical property of the resulting estimator.
Key to our approach is the following reformulation of the problem. We show in Lemma 1 of Section 2 that the discriminant directional matrix under model (1.1) and (1.3)
$B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}\in\mathbb{R}^{p\times L}$ (1.12)
and the regression matrix
$B:=\Sigma^{-1}\Sigma_{XY}:=[{\rm Cov}(X)]^{-1}{\rm Cov}(X,Y)$ (1.13)
not only share the same column space, but have a closed-form connection, to wit,
$B^{*}$ $=BH^{-1}$
for some invertible $L\times L$ matrix $H$. This connection is novel and distinguishes our approach from the current literature. Although the fact that $B$ and $B^{*}$ share the same span is relatively well-known (see, for instance, Hastie et al. (1995); Ye (2007); Lee and Kim (2015); Nie et al. (2022) and the references therein), their closed-form connection has not been discovered and a thorough theoretical investigation on the corresponding classifier has not been conducted, to the best of our knowledge. Specifically, we propose to estimate the discriminant matrix $B^{*}$ by estimating the regression matrix $B$ first
$\widehat{B}=\operatorname*{arg\,min}_{B\in\mathbb{R}^{p\times L}}~{}{1\over n}\|{\mathbb{Y}}-{\mathbb{X}}B\|_{F}^{2}+\text{pen}(B),$ (1.14)
and then estimating $H^{-1}$ based on $\widehat{B}$. The advantage of our approach is three-fold:
  1. (a)
    We can easily tap into the large literature on estimation of the high-dimensional regression matrix $B\in\mathbb{R}^{p\times L}$.
  2. (b)
    After estimating $B$, we only need to estimate the inverse of an $L\times L$ matrix rather than that of a $p\times p$ matrix $\Sigma_{w}$ as in the classical LDA context.
  3. (c)
    We are able to quantify the estimation error of both $B$ and $B^{*}$ as well as to provide theoretical guarantees for the resulting classifier.
While various penalties are feasible in our approach, more specifically in (1.14), we focus on a lasso-type estimator of $B$ in this paper. This also contrasts our approach with Cai and Zhang (2019) and Witten and Tibshirani (2011) which assume sparsity of $B^{*}$ and $F$, respectively. Our earlier work Bing and Wegkamp (2023) considers a multi-class classification setting with high-dimensional features $X\in\mathbb{R}^{p}$ that follow a factor model $X=AZ+W$ with unobserved (latent) features $Z\in\mathbb{R}^{r}$ and unobserved noise $W$, that is independent of both $Z$ and the label $Y$ with $r<n<p$. The procedure is entirely different than the one proposed here. Future work will be devoted on how to adapt the procedure in this paper to this low-rank setting and its analysis.
Under sparsity of the regression matrix $B$, we carefully analyze the properties of $\widehat{B}$ in Section 3.1, the estimation of $B^{*}$ via estimating $H^{-1}$ in Section 3.2 as well as the rate of convergence for the excess misclassification risk of the corresponding classification rule in Section 3.3. Computationally, after computing $\widehat{B}$ from (1.14) we offer an alternative way in Section 2.3 to estimate the Bayes rule that performs Fisher’s discriminant analysis on low-dimensional transformations $\widehat{B}^{\top}x\in\mathbb{R}^{L}$ of the original features $x\in\mathbb{R}^{p}$. This leads to a practically useful method, which extracts fewer discriminant directions than those in $B^{*}$, for both downstream analysis and visualization. Finally, we provide a thorough simulation study in Section 4 and we report our findings in several real data studies of Section 5.
2 Methodology
We discuss our classification procedure in this section. Our goal is to estimate the Bayes rule in (1.4) which, for a given $x\in\mathbb{R}^{p}$, requires estimation of the discriminant functions
$G_{\ell}(x):=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2x^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}),\qquad\text{for all }\ell\in[L].$ (2.1)
This, in turn, requires that we estimate the discriminant directional matrix $B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}$. The following lemma establishes the connection between $B^{*}$ and the regression matrix $B$ in (1.13). Recall the definitions of $\pi$, $D_{\pi}$ and ${\mathbb{M}}$ from (1.5) as well as $\Sigma={\rm Cov}(X)$.
Lemma 1.
Under model (1.1) and (1.3), we have
$B^{*}=BH^{-1}$ (2.2)
for a matrix $H\in\mathbb{R}^{L\times L}$ that can be inverted and satisfies
$H=D_{\pi}-B^{\top}\Sigma B=D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B=D_{\pi}-B^{\top}{\mathbb{M}}D_{\pi}.$ (2.3)
Proof.
From the identity
$\Sigma_{XY}=\mathbb{E}(XY^{\top})-\mathbb{E}(X)(\mathbb{E}(Y))^{\top}={\mathbb{M}}D_{\pi}$ (2.4)
we see that
$B=\Sigma^{-1}{\mathbb{M}}D_{\pi}.$ (2.5)
Next, using
$\Sigma=\Sigma_{w}+\Sigma_{b}=\Sigma_{w}+{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top},$ (2.6)
the Woodbury matrix identity yields
$\Sigma_{w}^{-1}=\Sigma^{-1}+\Sigma^{-1}{\mathbb{M}}\left(\mathbb{I}_{L}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}\right)^{-1}D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}.$
It follows that
$B^{*}$ $=\Sigma_{w}^{-1}{\mathbb{M}}$
$=\Sigma^{-1}{\mathbb{M}}+\Sigma^{-1}{\mathbb{M}}\left(\mathbb{I}_{L}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}\right)^{-1}D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}$
$=\Sigma^{-1}{\mathbb{M}}\left(\mathbb{I}_{L}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}\right)^{-1}$
$=\Sigma^{-1}{\mathbb{M}}D_{\pi}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}D_{\pi}\right)^{-1}$
$=B\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)^{-1}$
which completes the proof. ∎
Lemma 1 suggests that estimation of $B^{*}$ could be done by first estimating the matrix $B$ and then estimating $H^{-1}$ according to (2.3). We discuss these two steps in details in the next two sections.
2.1 Estimation of the regression matrix
In view of (1.13), it is natural to estimate $B$ via regularized regression methods in (1.14). Choices of the penalty term $\text{pen}(B)$ depend on the concrete problem at hand. For example, if we believe that only a subset of the original features are important for predicting the label, then sparsity structures of the regression matrix $B$ are reasonable assumptions. If the subset of features, that are pertinent for predicting the label, varies across different levels, then the entry-wise lasso penalty is appropriate. To fix ideas and due to technical complications, we focus our analysis to the matrix estimator $\widehat{B}$ in (1.14) obtained with the following concrete lasso penalty
$\text{pen}(B)$ $=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{1}$ (2.7)
for some tuning parameter $\lambda>0$. It is fairly straightforward to modify our analysis to other choices of the penalty term discussed below.
Remark 1 (Alternative penalty choices).
If there exist many features that are not predictive for all levels of the response label, then the group-lasso penalty $\text{pen}(B)=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ is a more suitable choice. In this case, one can retain a much smaller subset of important features and achieve feature selection. Furthermore, one can also consider either a row-wise or a column-wise elastic net penalty if the features tend to (highly) correlate with each other. Another example that is commonly used in the literature is when entries of $B$ exhibit certain smoothness structure, such as some imaging or audio applications in Hastie et al. (1995). In this case, one could choose $\text{pen}(B)={\rm tr}(B^{\top}\Omega B)$ for some pre-specified “roughness” penalty matrix $\Omega$. On the other hand, if the goal is to construct a low-dimensional subset of linear combinations of the original feature for classification purpose, then a low-rank penalty in (1.14) could be more appropriate. For example, if the matrix of conditional means ${\mathbb{M}}\in\mathbb{R}^{p\times L}$ has a reduced rank $r\ll\min\{p,L\}$ such as under the model in Bing and Wegkamp (2023), then Lemma 1 implies that ${\rm rank}(B)\leq r$. Estimation of $B\in\mathbb{R}^{p\times L}$ in (1.14) could be done by using either the reduced rank penalty $\text{pen}(B)=\lambda~{}{\rm rank}(B)$ or the nuclear norm penalty $\text{pen}(B)=\lambda\|B\|_{*}$. We will investigate this approach in another paper.
2.2 Estimation of the discriminant matrix and the Bayes rule
The discriminant functions in (2.1) that form the basis of the Bayes classification rule (1.4) necessitate that we estimate the matrix of the conditional means ${\mathbb{M}}$, the vector of prior probabilities $\pi$ as well as the discriminant directional matrix $B^{*}$. We adopt the following standard estimators
$\widehat{\mathbb{M}}={\mathbb{X}}^{\top}{\mathbb{Y}}({\mathbb{Y}}^{\top}{\mathbb{Y}})^{-1},\qquad D_{\widehat{\pi}}={\rm diag}(\widehat{\pi})={1\over n}{\mathbb{Y}}^{\top}{\mathbb{Y}}$ (2.8)
to estimate ${\mathbb{M}}$ and $D_{\pi}$, respectively. Recall from Lemma 1 that $B^{*}=BH^{-1}$. Since the estimator $\widehat{B}$ of $B$ is computed in (1.14), in view of the closed-form of $H$ in (2.3), we propose to estimate $H$ by the plug-in estimator
$\widehat{H}=D_{\widehat{\pi}}-{1\over n}\widehat{B}^{\top}{\mathbb{X}}^{\top}{\mathbb{X}}\widehat{B}={1\over n}\left({\mathbb{Y}}^{\top}{\mathbb{Y}}-\widehat{B}^{\top}{\mathbb{X}}^{\top}{\mathbb{X}}\widehat{B}\right).$ (2.9)
Note that ${\mathbb{X}}\widehat{B}$ is simply the in-sample fit of the regression in (1.14). Computation of ${\mathbb{X}}\widehat{B}$ thus can be done efficiently even in non-parametric regression settings, for instance, when the columns of ${\mathbb{X}}$ consist of basis expansions. Now, using $B^{*}=BH^{-1}$, we can further estimate $B^{*}$ by
$\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}.$ (2.10)
Here $M^{+}$ denotes the moore-penrose inverse of any matrix $M$. Finally, in view of (2.1), we estimate the discriminant functions by
$\widehat{G}_{\ell}(x)=\widehat{\mu}_{\ell}^{\top}\widehat{B}^{*}_{\ell}-2x^{\top}\widehat{B}^{*}_{\ell}-2\log(\widehat{\pi}_{\ell})\quad\text{for all $\ell\in[L]$. }$ (2.11)
The corresponding classifier is
$\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad\text{ for any $x\in\mathbb{R}^{p}$}.$ (2.12)
Theoretically, we will show that $\widehat{H}$ is invertible with overwhelming probability for the lasso-penalty (see, Remark 3 as well as the proof of Theorem 3 in Section A.3). Moreover, it is worth mentioning that the inverse of $\widehat{H}$ always exists when any $\ell_{2}$-norm related penalty is deployed in (1.14).
2.3 Classification based on reduced dimensions
Since in practice it is often desirable to extract fewer discriminant directions than those in $B^{*}$, we discuss in this section an alternative procedure for this purpose, which estimates the Bayes rule by also using $\widehat{B}$ from (1.14) but in a different way.
We start with the fact (its proof is stated in Section A.1 for completeness) that the Bayes rule in (1.4) is equivalent to
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}(x-\mu_{\ell})-2\log(\pi_{\ell})$ (2.13)
which, by Lemma 1, further equals to
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B(B^{\top}\Sigma_{w}B)^{+}B^{\top}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (2.14)
The formulation in (2.14) has the following two-step interpretation:
  1. (1)
    Transform the original feature $x\in\mathbb{R}^{p}$ to $z:=B^{\top}x$, with the latter belonging to a subspace of $\mathbb{R}^{L}$;
  2. (2)
    Perform linear discriminant analysis on the space of $z$.
The second point follows by noticing that
$C_{w}:=B^{\top}\Sigma_{w}B\in\mathbb{R}^{L\times L}$
is indeed the within-class covariance matrix of $Z=B^{\top}X$ under model (1.1).
Due to the fact that $C_{w}$ is rank deficient with ${\rm rank}(C_{w})<L$ and for the purpose of extracting fewer discriminant directions, we adopt the equivalent formulation of (2.14) from the Fisher’s perspective in the sequel. Write the between-class covariance matrix of $Z=B^{\top}X$ as
$C_{b}:=B^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}B.$
For a given reduced dimension $1\leq K<L$, we can find $K$ discriminant directions in the space of $Z$ by solving, for each $k\in[K]$, the minimization problem
$\begin{split}\alpha_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~{}\alpha^{\top}C_{b}~{}\alpha\\ \text{such that }&~{}\alpha^{\top}C_{w}~{}\alpha=1,~{}\alpha^{\top}C_{w}~{}\alpha_{i}=0,~{}~{}\text{ for all }~{}i<k.\end{split}$
The resulting classification rule based on these $K$ discriminant directions is
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B\sum_{k=1}^{K}\alpha_{k}\alpha_{k}^{\top}B^{\top}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (2.15)
In practice, after computing $\widehat{B}$ from (1.14), we propose to estimate $C_{b}$ and $C_{w}$ by
$\widehat{C}_{b}=\widehat{B}^{\top}\widehat{\mathbb{M}}D_{\widehat{\pi}}\widehat{\mathbb{M}}^{\top}\widehat{B}={1\over n}\widehat{B}^{\top}{\mathbb{X}}^{\top}P_{{\mathbb{Y}}}{\mathbb{X}}\widehat{B},\qquad\widehat{C}_{w}={1\over n}\widehat{B}^{\top}{\mathbb{X}}^{\top}(\mathbb{I}_{n}-P_{{\mathbb{Y}}}){\mathbb{X}}\widehat{B},$
with $P_{{\mathbb{Y}}}={\mathbb{Y}}({\mathbb{Y}}^{\top}{\mathbb{Y}})^{-1}{\mathbb{Y}}^{\top}$. Subsequently, the estimated discriminant directions $\widehat{\alpha}_{k}$, for $1\leq k\leq K$, are computed by solving
$\begin{split}\widehat{\alpha}_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~{}\alpha^{\top}\widehat{C}_{b}~{}\alpha\\ \text{such that}&~{}\alpha^{\top}\widehat{C}_{w}~{}\alpha=1,\quad\alpha^{\top}\widehat{C}_{w}~{}\widehat{\alpha}_{i}=0,~{}~{}\forall~{}i<k,\end{split}$ (2.16)
By similar reasoning for solving the Fisher’s discriminant analysis in (1.7), we propose to solve the optimization problem (2.16) via eigen-decomposing
$[\widehat{C}_{w}^{+}]^{1\over 2}\widehat{C}_{b}[\widehat{C}_{w}^{+}]^{1\over 2}.$
The eigenvalues of the above matrix can also be used to determine the upper bound for $K$. For instance, $K$ should be chosen no greater than the rank of the above matrix.
Remark 2 (Numerical stability of computing (2.16)).
The procedure in (2.16) computes Fisher’s discriminant directions $(\widehat{\alpha}_{1},\ldots,\widehat{\alpha}_{K})$ in
$\text{span}({\mathbb{X}}\widehat{B})\subset\text{span}({\mathbb{X}})$
of dimension less than $L$. Consequently, as long as
${\rm rank}((\mathbb{I}_{n}-P_{{\mathbb{Y}}}){\mathbb{X}}\widehat{B})={\rm rank}({\mathbb{X}}\widehat{B}),$ (2.17)
we do not suffer numerically in solving (2.16), in stark contrast with the algorithmic issue of solving (1.7) (or a penalized version thereof) in the original $p$-dimensional space. As pointed out by Wu et al. (2015), the latter leads to meaningless solutions with the objective function equal to infinity. Finally, we remark that the requirement in (2.17) is easy to satisfy as ${\rm rank}({\mathbb{X}}\widehat{B})\leq L$ and ${\rm rank}(\mathbb{I}_{n}-P_{{\mathbb{Y}}})=n-L$ when we observe at least one data point per class.
3 Theoretical guarantees
In this section we provide theoretical guarantees for our procedure. We start by listing the main assumptions used in our analysis. The following condition on the priors $\pi=(\pi_{1},\ldots,\pi_{L})$ is assumed.
Assumption 1.
There exists some absolute constants $0<c\leq C<\infty$ such that
${c\over L}\leq\min_{k\in[L]}\pi_{k}\leq\max_{k\in[L]}\pi_{k}\leq{C\over L}.$ (3.1)
This is a common regularity assumption even in binary classification problems. We may relax Assumption 1 to $\min_{k\in[L]}\pi_{k}\geq c\log(n)/n$ in our analysis (see, for instance, Lemma 15 of Appendix C) at the cost of more technical proofs and a less transparent presentation.
As we mentioned earlier, we decided to focus on the case where the columns of $B$ are (potentially) sparse and allow for different sparsity patterns across its columns. Therefore, for some positive integer $s\leq p$, we consider the following parameter space of $B$:
$\Theta_{s}:=\left\{B\in\mathbb{R}^{p\times L}:\max_{\ell\in[L]}\|B_{\ell}\|_{0}\leq s\right\}.$
For ease of presentation, we simply use the largest number of non-zero entries among all columns of $B$. However, our analysis can be easily extended to $\|B_{\ell}\|_{0}\leq s_{\ell}$ by allowing different $s_{\ell}$ for $\ell\in[L]$. Moreover, the analysis of $\widehat{B}$ relies on the standard Restricted Eigenvalue condition (RE) on the matrix $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$. Since ${\mathbb{X}}$ is a matrix with random entries, we resort to the following quantity on the within-class covariance matrix $\Sigma_{w}$.
Definition 1.
For any integer $1\leq s\leq p$, define
$\kappa_{s}:=\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{v^{\top}\Sigma_{w}v\over v^{\top}v}$
where $\mathcal{C}(S,3):=\{u\in\mathbb{R}^{p}\setminus\{0\}:\|u_{S^{c}}\|_{1}\leq 3\|u_{S}\|_{1}\}$.
Our analysis needs the following regularity conditions on $\Sigma_{w}$ and the matrix ${\mathbb{M}}$ of conditional means. For any matrix $M$, write $\|M\|_{\infty}=\max_{i,j}|M_{ij}|$.
Assumption 2.
For some $1\leq s\leq p$, there exists some absolute constants $0<c\leq C<\infty$ such that $c<\kappa_{s}\leq\|\Sigma_{w}\|_{\infty}\leq C$.
Assumption 3.
There exists an absolute constant $C<\infty$ such that $\|{\mathbb{M}}\|_{\infty}\leq C$.
A condition akin to 2 is commonly used in sparse linear regression models, see, for instance, Bühlmann and Van de Geer (2011); Giraud (2021). 3 is a mild regularity condition on the largest entry of ${\mathbb{M}}$. We remark that the requirements that $\kappa_{s}$, $\|\Sigma_{w}\|_{\infty}$ and $\|{\mathbb{M}}\|_{\infty}$ are absolutely bounded are not essential, and can be relaxed to allow dependence on $n$.
3.1 Finite sample bounds of the proposed regression matrix estimation
The following theorem states finite sample upper bounds of both the (in-sample and out-of-sample) prediction error and the estimation error of the regression matrix estimator $\widehat{B}$ in (1.14) with the lasso-penalty in (2.7). For any numbers $a,b\in\mathbb{R}$, we write $a\vee b=\max\{a,b\}$. For any sequences $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ if there exists some constant $C>0$ such that $a_{n}\leq Cb_{n}$, and $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$.
Theorem 2.
Under model (1.1) and (1.3) with Assumptions 13, assume that there exists some sufficiently small constant $c>0$ such that $B\in\Theta_{s}$ for some $1\leq s\leq p$ with
$(s\vee L)\log(n\vee p)\leq c~{}n$
Then, with probability at least $1-n^{-1}$, we have
$\max_{\ell\in[L]}{1\over\sqrt{n}}\|{\mathbb{X}}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}$
$\max_{\ell\in[L]}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}$
$\max_{\ell\in[L]}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}~{}\lesssim~{}\kappa_{s}^{-1}\lambda\sqrt{s}$
for the estimator $\widehat{B}$ obtained from (1.14) with any $\lambda$ satisfying
$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\infty}+\|{\mathbb{M}}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (3.2)
and some finite, large enough constant $C>0$.
Proof.
See Section A.2. ∎
The two main difficulties in our proofs are (a) to control $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)\|_{\infty}$, and (b) to bound from below the restricted eigenvalue of the Gram matrix $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$. For (a), as pointed out in Gaynanova (2020), the difficulty in this model is elevated relative to the existing analysis in sparse linear regression models due to the fact ${\mathbb{Y}}$ is not a linear model in ${\mathbb{X}}$. We establish a sharp control of $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)\|_{\infty}$ in Lemma 6. For (b), the existing result in Rudelson and Zhou (2012) is not readily applicable as rows of ${\mathbb{X}}$ are generated from a mixture of Gaussians. In Lemma 7 we establish a uniform bound of $v^{\top}(\widehat{\Sigma}-\Sigma)v$ over $s$-sparse vectors $v$ which, in conjunction with the reduction arguments in Rudelson and Zhou (2012), proves (b). Our analysis in this paper allows the number of classes $L$ to grow with the sample size $n$
Gaynanova (2020) considers a group-lasso regularized regression similar to (1.14) and derives the bound of the $\ell_{2}$-norm $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbb{X}}^{\top}({\mathbb{Y}}^{\prime}-{\mathbb{X}}B^{\prime})\|_{2}$ for some different response matrix ${\mathbb{Y}}^{\prime}$ and different target matrix $B^{\prime}$. Their proof is different and in particular requires that $L$ is fixed, independent of $n$.
In the rest of the paper, we assume that $\lambda$ is chosen as the order specified in (3.2). In practice, one can simply choose $\lambda$ via cross-validation, for instance, via the built-in function of the R-package glmnet.
3.2 Estimation of the discriminant directional matrix
We analyze the estimation of the discriminant matrix $B^{*}$ in this section. Recall that
$\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$
with $\widehat{H}$ obtained from (2.9). The following theorem states upper bounds of the transformed $(\widehat{B}^{*}-B^{*})$. Since $H^{-1}$ depends on ${\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}$ (see Lemma 9 of Appendix B for its proof), our results depend in turn on the quantity
$\Delta_{\infty}=\max_{\ell\in[L]}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}.$
This quantity essentially reflects the largest Mahalanobis distance in the feature space between any two label classes as can be seen from the relation (under (1.2))
${1\over 4}\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\leq\Delta_{\infty}\leq\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}).$
Theorem 3.
Under the conditions in Theorem 2, assume that $\Delta_{\infty}\geq 1$ and assume there exists some sufficiently small constant $c>0$ such that
$\Delta_{\infty}^{2}\sqrt{sL\log(n\vee p)\over n}\leq c.$ (3.3)
Then, with probability at least $1-n^{-1}$, we have
$\max_{\ell\in[L]}~{}\bigl{\|}\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\bigr{\|}_{2}\!\!~{}\lesssim~{}\Delta_{\infty}^{5/2}\sqrt{sL\log(n\vee p)\over n}.$
Proof.
See Section A.3. ∎
Remark 3.
Condition (3.3) ensures that, with high probability, $\lambda_{K}(\widehat{H})\geq\lambda_{K}(H)/2$ holds, and hence $\widehat{H}$ is invertible. In Theorem 4 below we will see that condition (3.3) is also needed for consistent estimation of the discriminant function. The reasonable condition $\Delta_{\infty}\geq 1$ requires that the largest separation among all pairs of conditional distributions does not vanish. It is assumed to simplify the presentation, and our analysis can be easily extended to the case $\Delta_{\infty}\to 0$ as $n\to\infty$.
3.3 Finite sample bounds of the misclassification rates
We first analyze the estimation of the discriminant functions in (2.1). For any $X$ that is distributed according to (1.1), we have
$\widehat{G}_{\ell}(X)-G_{\ell}(X)$ $=(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}+(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})+2\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right).$
As a consequence of Theorem 3, the Gaussian distribution assumption of $X\mid Y$ and the independence between $X$ and $({\mathbb{X}},{\mathbb{Y}})$, we have the following control of the estimated discriminant function $\widehat{G}_{\ell}(X)$ uniformly over $\ell\in[L]$.
Theorem 4.
Under the conditions of Theorem 3, we have
$\mathbb{P}\left\{\max_{\ell\in[L]}|\widehat{G}_{\ell}(X)-G_{\ell}(X)|~{}\gtrsim~{}\Delta_{\infty}^{3}\sqrt{sL\log(n\vee p)\log(n)\over n}\right\}\leq 2n^{-\Delta_{\infty}}+n^{-1}.$
Proof.
See Section A.4. ∎
Armed with Theorem 4, we are ready to analyze the proposed classifier $\widehat{g}$ in (2.12) in terms of its excess risk relative to error rate of the Bayes rule in (1.4). Specifically, for any pair $(X,Y)$ from (1.1) that is independent of $({\mathbb{X}},{\mathbb{Y}})$, the excess risk of $\widehat{g}$ is defined as
$\mathcal{R}(\widehat{g}):=\mathbb{P}\left\{Y\neq{\widehat{g}(X)}\right\}-\mathbb{P}\left\{Y\neq{g^{*}(X)}\right\}.$
Note that we use the identification $\{Y=k\}:=\{Y={\bm{e}}_{k}\}$ in the above definition. In the proof of the following theorem, we show that, for any $t>0$,
$\mathcal{R}(\widehat{g})$ $\leq(L-1)~{}t~{}\max_{k\in[L]}\max_{\ell\in[L]\setminus\{k\}}\mathbb{P}\left\{-t\leq G^{(\ell|k)}(X)\leq 0\mid Y={\bm{e}}_{k}\right\}$
$\quad+(L-1)~{}\mathbb{P}\left\{\max_{\ell\in[L]}\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|>t\right\}$
This fact, in combination with Theorem 4, yields the following result on the rates of convergence of $\mathcal{R}(\widehat{g})$.
Theorem 5.
Under the same conditions of Theorem 3 and $\Delta_{\infty}\asymp 1$, we have
$\mathcal{R}(\widehat{g})~{}\lesssim~{}{sL^{2}\log(n\vee p)\log(n)\over n}.$
Proof.
See Section A.5. ∎
Our results allow both the sparsity level $s$ and the number of classes $L$ to grow with the sample size $n$. The existing literature (Cai and Zhang, 2019; Gaynanova, 2020) that analyze the excess risk only consider the case $L\asymp 1$, while Abramovich and Pensky (2019) does allow $s$, $L$ and $n$ to grow, but studies the risk $\mathbb{P}\{Y\neq g(X)\}$ instead of the excess risk $\mathcal{R}(g)$.
Theorem 5 focuses on the statistically most interesting case $\Delta_{\infty}\asymp 1$. Otherwise, if the pairwise separation between different classes grows to infinity, the classification problem becomes much easier and we have super-fast rates of convergence (see, for instance, Cai and Zhang (2019); Bing and Wegkamp (2023)). Indeed, provided
$\min_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\asymp\max_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}),$
using the arguments in the proof of Bing and Wegkamp (2023, Theorem 12), our analysis can be modified easily to show that
$\mathcal{R}(\widehat{g})~{}\lesssim~{}{sL^{2}\log(n\vee p)\log(n)\over n}e^{-c\Delta_{\infty}}+n^{-a}$
for some arbitrary large constant $a>0$ and some constant $c>0$. The first term on the right in the above rate is exponentially fast in $\Delta_{\infty}\to\infty$.
4 Simulation study
In our simulation study, we evaluate our proposed procedure in Section 2.2 that we term as Linear Discriminant Regularized Regression (LDRR). We also include the procedure in Section 2.3 (called LDRR-F) that is based on Fisher’s discriminant rule after estimating $B$. For estimating $B$ from (1.14), in addition to the lasso penalty (2.7), we also consider the elastic net penalty, $\text{pen}(B)=\lambda\sum_{\ell=1}^{L}[\alpha\|B_{\ell}\|_{2}+(1-\alpha)\|B_{\ell}\|_{1}]$, for some $\alpha\in(0,1)$ and $\lambda>0$ chosen by cross-validation from the R-package glmnet. We compare our proposed methods with the nearest shrunken centroids classifier (PAMR) of Tibshirani et al. (2002), the shrunken centroids regularized discriminant analysis (RDA) method of Guo et al. (2007) and the $\ell_{1}$-penalized linear discriminant (PenalizedLDA) method of Witten and Tibshirani (2011). These methods are available in R-packages pamr, rda, penalizedLDA, and we use their default methods for selecting tuning parameters. Finally, the oracle procedure that uses the true values of ${\mathbb{M}}$, $\Sigma_{w}$, and $\pi$ in the computation of the discriminant functions $G_{k}(x)$ in (2.1) serves as our benchmark.
We generate the data from the following mean-shift scenario. A similar setting is also used in Witten and Tibshirani (2011). Specifically, for any $\ell\in[L]$ and its corresponding mean vector $\mu_{\ell}\in\mathbb{R}^{p}$, we sample $\mu_{\ell,j}$, for $5(\ell-1)\leq j\leq 5\ell$, independently from $N(0,2^{2})$ and set the other entries of $\mu_{\ell}$ to 0. Regarding the within-class covariance matrix, we set $\Sigma_{w}=\sigma^{2}W$ where $W$ is generated by independently sampling its diagonal elements from Uniform(1,3), and setting its off-diagonal elements as $W_{ij}=\sqrt{W_{ii}W_{jj}}\rho^{|i-j|}$, for all $i\neq j$. The quantity $\sigma>0$ controls the overall noise level while the coefficient $\rho\in[0,1]$ controls the within-class correlation among the features. We generate the class probabilities $\pi=(\pi_{1},\ldots,\pi_{L})^{\top}$ as follows. For a given $\alpha\geq 0$, we set $\pi_{\ell}=\nu_{\ell}^{\alpha}/(\sum_{i=1}^{L}\nu_{i}^{\alpha})$ for all $\ell\in[L]$. Here $\nu_{1},\ldots,\nu_{L}$ are independent draws from the Uniform$(0,1)$. The quantity $\alpha$ controls the imbalance of each class size: a larger $\alpha$ corresponds to more imbalanced class sizes. At the other extreme, $\alpha=0$ corresponds to the balanced case $\pi=(1/L)1_{L}$. In the following, we examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{100,300,500,700,900\}$; the feature dimension $p\in\{100,300,500,700,900\}$; the number of classes $L\in\{2,5,8,11,14\}$; the correlation coefficient $\rho\in\{0,0.2,0.4,0.6,0.8\}$; the noise parameter $\sigma\in\{0.6,0.9,1.2,1.5,1.8\}$; the imbalance of the class probabilities $\alpha\in\{0,0.4,0.8,1.2,1.6\}$.
For all settings, we fix $n=300$, $p=500$, $L=5$, $\rho=0.6$, $\sigma=1$ and $\alpha=0$ when they are not varied. We use $500$ test data points to compute the misclassification errors of each procedure, and for each setting, we consider $50$ repetitions and report the averaged misclassification errors in Figure 1.
From Figure 1, it is clear that our procedures consistently outperform other methods across all settings. We observe that the performance of our algorithms gets improved as $n$ increases, $L$ decreases and $\sigma$ decreases. Moreover, neither the imbalance of the class probabilities nor the ambient dimension $p$ seems to affect the misclassification error of our algorithms. These findings are in line with our theory in Section 3. It is a bit surprising to see that the classification error of our benchmark decreases as the correlation coefficient $\rho$ keeps increasing after $0.4$. Nevertheless, both our algorithms and RDA adapt to this situation. Finally, we find that both LDRR and LDRR-F as well as their lasso and elastic net variants have very comparable performance in all scenarios. This is as expected as the data is simulated according to model (1.1) and (1.3), although in practice the Fisher’s discriminant analysis (LDRR-F) with the elastic net penalty could have more robust performance, as revealed in our real data analysis in the next section.
Figure 1: The averaged misclassification errors of each algorithm for varying each parameter one at a time.
5 Real data analysis
In this section we evaluate the performance of our approach and compare it with other algorithms on three high-dimensional biological data sets that are commonly used in the literature for discriminant analysis.
The Ramaswamy dataset has $n=198$ samples and each sample has features of $p=16,063$ gene expressions and belongs to one of the $L=14$ distinct cancer subtypes (Ramaswamy et al., 2001). The SRBCT dataset (Khan et al., 2001) has $n=83$ samples and consists of microarray analyses of childhood cancer tumors, more specifically, small round blue cell tumors (SRBCT). These tumors are categorized into $L=4$ classes: Burkitt lymphoma, Ewing sarcoma, neuroblastoma, and rhabdomyosarcoma. Each sample has $p=2,308$ gene expressions as features. The third dataset has $n=62$ samples and consists of $L=3$ distinct lymphoma subtypes. These lymphoma are categorized into diffuse large B-cell lymphoma, follicular lymphoma and chronic lymphocytic leukemia. Each sample has $p=4026$ gene expressions as its features. We used the preprocessed Lymphoma data in Dettling (2004).
For all three data sets, the features are centered to zero mean. Each dataset is randomly divided with 75% of samples as the training data and 25% as the test data. Such process was repeated 10 times and the averaged misclassification errors of each algorithm were reported in Table 1.
Table 1: The averaged misclassification errors (in percentage).
LDRR PenalizedLDA PAMR RDA
Ramaswamy 14.4 37.6 28.8 15.4
SRBCT 1.0 5.2 1.4 1.0
Lymphoma 0.0 0.6 0.6 0.6
Given the similar performance between LDRR and LDRR-F as well as their variants of using the lasso and the elastic-net penalty, we only compare LDRR-F using the elastic net penalty with PenalizedLDA, PAMR and RDA in our real data analysis due to its robustness against model misspecification. From Table 1, we see that our procedure LDRR-F has the smallest misclassification errors across all three datasets. Furthermore, in Figure 2, we show that LDRR-F with $K=2$ in (2.15) can be used for visualization in the 2D space of the first two discriminant directions.
Figure 2: Visualization of Ramaswamy, SRBCT and Lymphoma data points in the space of the first two discriminant vectors.
Acknowledgements
Wegkamp is supported in part by the National Science Foundation grant NSF DMS-2210557.
Appendix A Proofs
Throughout the proofs, we will use the following notation. Define
$\Delta={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}$
and write
$\Delta_{\infty}=\max_{\ell\in[L]}\Delta_{\ell\ell},\qquad\Delta_{\rm op}=\|\Delta\|_{\rm op}.$
Further write
$\pi_{\min}=\min_{\ell\in[L]}\pi_{\ell},\qquad\pi_{\max}=\max_{\ell\in[L]}\pi_{\ell}.$
The following event will be frequently used in our proofs. Define the event
$\mathcal{E}_{\pi}=\left\{\max_{k\in[L]}\left|\widehat{\pi}_{k}-\pi_{k}\right|\leq C\sqrt{\log(n)\over nL}\right\}$ (A.1)
for some absolute constant $C>0$. Lemma 15 in Appendix C ensures that $\mathbb{P}(\mathcal{E}_{\pi})\geq 1-2n^{-2}$ under 1. Also note that, on $\mathcal{E}_{\pi}$ and under the condition $L\log(n)\leq n$,
$\max_{\ell\in[L]}\left|\log\left({\pi_{\ell}\over\widehat{\pi}_{\ell}}\right)\right|\leq\max_{\ell\in[L]}{|\widehat{\pi}_{\ell}-\pi_{\ell}|\over\pi_{\ell}}\lesssim\sqrt{L\log(n)\over n}$ (A.2)
as well as $\widehat{\pi}_{\ell}\asymp\pi_{\ell}$ for all $\ell\in[L]$.
A.1 Proof of the equivalence between (1.4) and (2.13)
Proof.
Recall that $B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}$. We find that
$B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}$ $=\Sigma_{w}^{-1}{\mathbb{M}}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{+}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}$
so that, for any $\ell\in[L]$,
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}(x-\mu_{\ell})-2\log(\pi_{\ell})$
$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{\mathbb{M}}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{+}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$
$\qquad+\mu_{\ell}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{+}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}).$
Write the singular value decomposition of $\Sigma_{w}^{-1/2}{\mathbb{M}}$ as $U\Lambda V^{\top}$ with $U\in\mathbb{R}^{p\times r}$ containing the first $r$ singular vectors and $r={\rm rank}(\Sigma_{w}^{-1/2}{\mathbb{M}})$. It follows that the above equals to
$\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{\mathbb{M}}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{\mathbb{M}}^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{\mathbb{M}}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{\mathbb{M}}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
which, in view of (1.4) and (2.1), completes the proof. ∎
A.2 Proof of Theorem 2
Proof.
We work on the intersection of the event
$\mathcal{E}_{\lambda}:=\left\{{1\over n}\|{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)\|_{\infty}\leq{1\over 2}\lambda\right\}$ (A.3)
with the event
$\mathcal{E}_{RE}:=\left\{\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{\|{\mathbb{X}}v\|_{2}^{2}\over n\|v\|_{2}^{2}}\geq{\kappa_{s}\over 2}\right\}.$ (A.4)
Recall $\mathcal{C}(S,3):=\{u\in\mathbb{R}^{p}\setminus\{0\}:\|u_{S^{c}}\|_{1}\leq 3\|u_{S}\|_{1}\}$. The main difficulty in proving Theorem 2 is to establish the order of $\lambda$ as well as to show that $\mathcal{E}_{\lambda}\cap\mathcal{E}_{RE}$ holds with overwhelming probability.
Lemma 6 ensures that $\mathcal{E}_{\lambda}$ holds with probability at least $1-n^{-1}$ for any $\lambda$ satisfying (3.2).
We proceed to verify that $\mathbb{P}(\mathcal{E}_{RE})\geq 1-(n\vee p)^{-2}$. To this end, for any set $S\subseteq[p]$, we write $E_{S}:=\text{span}\{{\bm{e}}_{j}:j\in S\}$. Under 2, we define the space
$\Gamma_{s}:=\bigcup_{S\subseteq[p]:\ |S|=\lfloor Cs\rfloor}E_{S}$ (A.5)
for some constant $C=C(\kappa_{s},\|\Sigma_{w}\|_{\infty})>1$. Let $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$. Lemma 7 and the condition $(s\vee L)\log(n\vee p)\leq cn$ imply that
${9\over 10}\|\Sigma^{1/2}v\|_{2}\leq\|\widehat{\Sigma}^{1/2}v\|_{2}\leq{11\over 10}\|\Sigma^{1/2}v\|_{2}\qquad\forall~{}v\in\Gamma_{s}$ (A.6)
holds with probability at least $1-(n\vee p)^{-2}$. According to Rudelson and Zhou (2012, Theorem 3), (A.6) implies that there exists $S\subseteq[p]$ with $|S|\leq s$ such that
$u^{\top}\widehat{\Sigma}u\geq{1\over 2}u^{\top}\Sigma u$ (A.7)
holds for any $u\in\mathcal{C}(S,3)$. Eq. A.7 in conjunction with 2 ensures that $\mathcal{E}_{RE}$ holds.
It is now relatively straightforward to finish the proof. Standard arguments (Bühlmann and Van de Geer, 2011; Giraud, 2021) yield that on the event $\mathcal{E}_{\lambda}$, for any $\ell\in[L]$,
${1\over 2n}\|{\mathbb{X}}\widehat{B}_{\ell}-{\mathbb{X}}B_{\ell}\|_{2}^{2}$ $\leq{\lambda\over 2}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}+\lambda\|B_{\ell}\|_{1}-\lambda\|\widehat{B}_{\ell}\|_{1}$
from which one can deduce that
$\widehat{B}_{\ell}-B_{\ell}\in\mathcal{C}(S_{\ell},3)$ (A.8)
with $S_{\ell}:={\rm supp}(B_{\ell})$. Invoking (A.7), we find on the event $\mathcal{E}_{RE}$
${\kappa_{s}\over 4}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}^{2}\leq{1\over 4}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}^{2}$ $\leq{1\over 2n}\|{\mathbb{X}}\widehat{B}_{\ell}-{\mathbb{X}}B_{\ell}\|_{2}^{2}$
$\leq{3\lambda\over 2}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}$
$\leq 6\lambda\|(\widehat{B}_{\ell}-B_{\ell})_{S_{\ell}}\|_{1}$ by (A.8)
$\leq 6\lambda\sqrt{s}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}.$
The proof is complete. ∎
A.2.1 Two key lemmas used in the proof of Theorem 2
Lemma 6.
Under model (1.1) and (1.3) with 1, assume $L\log(n\vee p)\leq n$. With probability at least $1-n^{-1}$, one has
$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbb{X}}_{j}^{\top}({\mathbb{Y}}_{\ell}-{\mathbb{X}}B_{\ell})\right|\lesssim\left(\sqrt{\|\Sigma_{w}\|_{\infty}}+\|{\mathbb{M}}\|_{\infty}\right)\sqrt{\log(n\vee p)\over nL}.$
Proof.
Note that
${1\over n}{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)$ $={1\over n}{\mathbb{X}}^{\top}{\mathbb{Y}}-\Sigma_{XY}-\widehat{\Sigma}B+\Sigma B$ (A.9)
We first analyze the term $(\widehat{\Sigma}-\Sigma)B$. Pick any $j\in[p]$, $\ell\in[L]$ and fix $t\geq 0$. We observe from Lemma 15 and the inequality $L\leq n$ that $\mathcal{E}_{\pi}$ holds with probability at least $1-2n^{-1}$. Next, we invoke Lemma 13 with $v_{1}={\bm{e}}_{j}$, $v_{2}=B_{\ell}$, $t=C\log(n\vee p)$ and we use (B.5), (B.6) and (B.7) in Lemma 9 to derive that the inequalities
$\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{[\Sigma_{w}]_{jj}B_{\ell}^{\top}\Sigma_{w}B_{\ell}}\sqrt{\log(n\vee p)\over n}$
$\quad+\|{\mathbb{M}}_{j\cdot}\|_{\infty}\sqrt{B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}}\sqrt{\log(n\vee p)\over nL}+\|{\mathbb{M}}_{j\cdot}\|_{\infty}\|{\mathbb{M}}^{\top}B_{\ell}\|_{\infty}{\log(n\vee p)\over n}$
$\quad+\sqrt{[\Sigma_{w}]_{jj}B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}}\sqrt{\log(n\vee p)\over nL}+\|{\mathbb{M}}_{j\cdot}\|_{2}\sqrt{B_{\ell}^{\top}\Sigma_{w}B_{\ell}}\sqrt{\log(n\vee p)\over nL}$
$\lesssim\sqrt{[\Sigma_{w}]_{jj}+\|{\mathbb{M}}_{j\cdot}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}+\|{\mathbb{M}}_{j\cdot}\|_{\infty}{\log(n\vee p)\over n}$
$\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{\mathbb{M}}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$
hold with probability at least $1-(n\vee p)^{-3}$. After we take the union bound over $j\in[p]$ and $\ell\in[L]$, we have that
$\max_{j\in[p],\ \ell\in[L]}\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{\mathbb{M}}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (A.10)
holds with probability $1-n^{-1}$.
To bound the first term in (A.9), by recalling that $\Sigma_{XY}={\mathbb{M}}D_{\pi}$, we find
${1\over n}{\mathbb{X}}_{j}^{\top}{\mathbb{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}$ $={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{L}\left(X_{ij}Y_{i\ell}-{\mathbb{M}}_{j\ell}\pi_{\ell}\right)\mathbbm{1}\{Y_{i}={\bm{e}}_{\ell}\}$
$={1\over n}\sum_{i=1}^{L}\left(X_{ij}-{\mathbb{M}}_{j\ell}\right)\mathbbm{1}\{Y_{i}=\ell\}+\left({n_{\ell}\over n}-\pi_{\ell}\right){\mathbb{M}}_{j\ell}$
$={n_{\ell}\over n}(\widehat{\mathbb{M}}_{j\ell}-{\mathbb{M}}_{j\ell})+(\widehat{\pi}_{\ell}-\pi_{\ell}){\mathbb{M}}_{j\ell}.$
Invoking 1, Lemma 15 and Lemma 16 with $t=\sqrt{\log(p\vee n)}$ and taking a union bound over $j\in[p]$ and $\ell\in[L]$, we obtain
$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbb{X}}_{j}^{\top}{\mathbb{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}\right|\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}\log(p\vee n)\over nL}+\|{\mathbb{M}}\|_{\infty}\sqrt{\log(n)\over nL}$ (A.11)
with probability $1-n^{-1}$. Combining (A.10) and (A.11) completes the proof. ∎
Recall the set
$\Gamma_{s}=\bigcup_{S\subseteq[p]:|S|=\lfloor Cs\rfloor}E_{S}$
with $E_{S}=\text{span}\{{\bm{e}}_{j}:j\in S\}$. The following lemma gives an upper bound for $v^{\top}(\widehat{\Sigma}-\Sigma)v$, uniformly over $v\in\Gamma_{s}$.
Lemma 7.
Under model and (1.3) with Assumptions 1 & 2, assume $(s\vee L)\log(n\vee p)\leq n$ for some integer $1\leq s<p$. Then, on the event $\mathcal{E}_{\pi}$, with probability at least $1-(n\vee p)^{-2}$, the following holds uniformly over $v\in\Gamma_{s}$,
$v^{\top}(\widehat{\Sigma}-\Sigma)v~{}\lesssim~{}v^{\top}\Sigma v\left(\sqrt{s\log(n\vee p)\over n}+\sqrt{L\log(n)\over n}\right).$
Proof.
Without loss of generality, we prove the result for
$v\in\Gamma_{s}^{\prime}:=\Gamma_{s}\cap\mathcal{S}^{p}$
via a discretization argument. For any subset $S\subseteq[p]$ with $|S|=\lfloor Cs\rfloor$, let $\mathcal{N}_{S}$ be the $(1/3)$-net of $E_{S}\cap\mathcal{S}^{p}$ satisfying
$\mathcal{N}_{S}\subset E_{S}\cap\mathcal{S}^{p},\qquad|\mathcal{N}_{S}|\leq 7^{|S|}$
The existence of such net is ensured by Rudelson and Zhou (2012, Lemma 23). Hence, for any fixed $v\in E_{S}\cap\mathcal{S}^{p}$, there exists $u\in\mathcal{N}_{S}$ such that $\|u-v\|_{2}\leq 1/3$, and we find
$v^{\top}(\widehat{\Sigma}-\Sigma)v$ $\leq(v-u)^{\top}(\widehat{\Sigma}-\Sigma)v+u^{\top}(\widehat{\Sigma}-\Sigma)(v-u)+u^{\top}(\widehat{\Sigma}-\Sigma)u$
$\leq 2\|u-v\|_{2}\sup_{v\in E_{S}\cap\mathcal{S}^{p}}v^{\top}(\widehat{\Sigma}-\Sigma)v+u^{\top}(\widehat{\Sigma}-\Sigma)u.$
The second inequality uses the fact that $(u-v)/\|u-v\|_{2}\in E_{S}\cap\mathcal{S}^{p}$ whenever $u\neq v$. Taking the supremum over $v\in E_{S}\cap\mathcal{S}^{p}$ and the maximum over $u\in\mathcal{N}_{S}$ gives
$\sup_{v\in E_{S}\cap\mathcal{S}^{p}}v^{\top}(\widehat{\Sigma}-\Sigma)v\leq 3\max_{u\in\mathcal{N}_{S}}u^{\top}(\widehat{\Sigma}-\Sigma)u$
and
$\sup_{v\in\Gamma_{s}^{\prime}}v^{\top}(\widehat{\Sigma}-\Sigma)v\leq 3\max_{u\in\mathcal{N}}u^{\top}(\widehat{\Sigma}-\Sigma)u.$ (A.12)
Here
$\mathcal{N}:=\bigcup_{S\subseteq[p]:\ |S|=\lfloor Cs\rfloor}\mathcal{N}_{S}$
has cardinality
$|\mathcal{N}|\leq\sum_{|S|=\lfloor Cs\rfloor}|\mathcal{N}_{S}|\lesssim{p}^{Cs}.$
We will bound the right hand side (A.12) via the same arguments used to prove Lemma 13, except for the control of the term $\sum_{k}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v$. Specifically, repeating the proof of Lemma 13 with $v_{1}=v_{2}=v$, for any $v\in\mathcal{N}$, gives that for any $t\geq 0$
$v^{\top}(\widehat{\Sigma}-\Sigma)v$ $\lesssim v^{\top}\Sigma_{w}v\left(\sqrt{t\over n}+{t\over n}\right)+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v+2{\|{\mathbb{M}}^{\top}v\|_{2}\over\sqrt{L}}\sqrt{v^{\top}\Sigma_{w}v}\sqrt{t\over n}.$
holds, with probability at least $1-6e^{-t}$. Note that
${\|{\mathbb{M}}^{\top}v\|_{2}^{2}\over L}\lesssim v^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}v$
under 1. By using the decomposition of $\Sigma$ in (2.6), choosing $t=Cs\log(n\vee p)$ for some large constant $C>0$, we obtain that
$v^{\top}(\widehat{\Sigma}-\Sigma)v\lesssim v^{\top}\Sigma v\left(\sqrt{s\log(n\vee p)\over n}+{s\log(n\vee p)\over n}\right)+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v$
holds uniformly over $v\in\mathcal{N}$, with probability at least $1-(n\vee p)^{-C^{\prime\prime}s}$. On the event $\mathcal{E}_{\pi}$, invoking 1, we conclude that
$\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v\leq\max_{k\in[L]}{|\widehat{\pi}_{k}-\pi_{k}|\over\pi_{k}}v^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}v\lesssim v^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}v\sqrt{L\log(n)\over n}$
holds uniformly over $\mathcal{N}$. The proof is complete in view of (A.12). ∎
A.3 Proof of Theorem 3
Proof.
Define
$\delta\asymp\Delta_{\infty}\sqrt{\log(n)\over nL}+\sqrt{\Delta_{\infty}}\|{\mathbb{M}}\|_{\infty}\sqrt{s\log(n\vee p)\over nL}.$ (A.13)
We work on the event
$\mathcal{E}_{H}:=\left\{\|\widehat{H}-H\|_{\infty,1}\leq\delta\right\}$ (A.14)
which, according to Lemma 8, holds with probability at least $1-n^{-1}$. Moreover, on the event $\mathcal{E}_{H}$, we find that
$\lambda_{K}(\widehat{H})$ $\geq\lambda_{K}(H)-\|\widehat{H}-H\|_{\rm op}$ by Weyl’s inequality
$=\left[\lambda_{1}(\Omega)\right]^{-1}-\|\widehat{H}-H\|_{1,\infty}$ $\text{by writing }\Omega:=H^{-1}$
[…]
Hence, $\widehat{H}$ is invertible on the event $\mathcal{E}_{H}$. For simplicity, we write $\widehat{\Omega}=\widehat{H}^{-1}$. By definition,
$\widehat{B}^{*}-B^{*}$ $=(\widehat{B}-B)\widehat{\Omega}+B(\widehat{\Omega}-\Omega)$
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}H(\widehat{\Omega}-\Omega)$ by Lemma 1
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H\widehat{\Omega}-\mathbb{I}_{L})$ by $\text{by }H\Omega=\mathbb{I}_{L}$
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H-\widehat{H})\widehat{\Omega}$ by $\text{by }\widehat{H}\widehat{\Omega}=\mathbb{I}_{L}.$
Recall that (A.17) and pick any $\ell\in[L]$. We it then follows that
$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\widehat{\Omega}_{\ell}\|_{2}+\|\Sigma_{w}^{1/2}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{2}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{2,\infty}\|\widehat{\Omega}_{\ell}\|_{1}+\max_{k\in[L]}\sqrt{B_{k}^{*\top}\Sigma_{w}B^{*}_{k}}~{}\|(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{1}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{2,\infty}\|\widehat{\Omega}_{\ell}\|_{1}+\sqrt{\Delta_{\infty}}\|\widehat{H}-H\|_{\infty,1}\|\widehat{\Omega}_{\ell}\|_{1}$ by $\text{by }B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}.$
To bound from above $\|\widehat{\Omega}_{\ell}\|_{1}$, we have that
$\|\widehat{\Omega}_{\ell}\|_{1}$ $\leq\|\Omega_{\ell}\|_{1}+\|(\widehat{\Omega}-\Omega){\bm{e}}_{\ell}\|_{1}$
$=\|\Omega_{\ell}\|_{1}+\|\Omega(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{1}$
$\leq\|\Omega_{\ell}\|_{1}+\|\Omega\|_{\infty,1}\|H-\widehat{H}\|_{\infty,1}\|\widehat{\Omega}_{\ell}\|_{1}.$
Since the event $\mathcal{E}_{H}$ and (3.3) ensure that
$\|\Omega\|_{\infty,1}\|H-\widehat{H}\|_{\infty,1}\leq\left(\min_{k\in[K]}{1\over\pi_{k}}+\|\Delta\|_{1,\infty}\right)\delta\lesssim\delta L\Delta_{\infty}\leq{1\over 2},$
we conclude, for any $\ell\in[L]$,
$\|\widehat{\Omega}_{\ell}\|_{1}\leq 2\|\Omega_{\ell}\|_{1}\lesssim L\Delta_{\infty}.$ (A.15)
It then follows that
$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}$
$\lesssim\left(\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{2,\infty}+\delta\sqrt{\Delta_{\infty}}\right)L\Delta_{\infty}$
$\lesssim\Delta_{\infty}\|{\mathbb{M}}\|_{\infty}\sqrt{sL\log(n\vee p)\over n}+\sqrt{\Delta_{\infty}^{5}}\sqrt{L\log(n)\over n}+\Delta_{\infty}^{2}\|{\mathbb{M}}\|_{\infty}\sqrt{sL\log(n\vee p)\over n}$
$\lesssim\Delta_{\infty}^{5/2}\|{\mathbb{M}}\|_{\infty}\sqrt{sL\log(n\vee p)\over n},$ (A.16)
completing the proof. ∎
The following lemma provide upper bounds for $\|\widehat{H}-H\|_{\infty,1}$ under condition of Theorem 3.
Lemma 8.
Under conditions of Theorem 3, with probability at least $1-n^{-1}$,
$\|\widehat{H}-H\|_{\infty,1}\lesssim\Delta_{\infty}\sqrt{\log(n)\over nL}+\sqrt{\Delta_{\infty}}\|{\mathbb{M}}\|_{\infty}\sqrt{s\log(n\vee p)\over nL}.$
Proof.
Recall that
$\begin{split}\widehat{H}-H&=D_{\widehat{\pi}}-D_{\pi}-\left(\widehat{B}^{\top}\widehat{\Sigma}\widehat{B}-B^{\top}\Sigma B\right)\\ &=D_{\widehat{\pi}}-D_{\pi}-B^{\top}(\widehat{\Sigma}-\Sigma)B-(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\\ &\qquad-(\widehat{B}-B)^{\top}\widehat{\Sigma}B-B^{\top}\widehat{\Sigma}(\widehat{B}-B).\end{split}$ (A.17)
We first bound the $\|\cdot\|_{\infty,1}$ norm. By triangle inequality, we have
$\|\widehat{H}-H\|_{\infty,1}$ $\leq\|D_{\widehat{\pi}}-D_{\pi}\|_{\infty,1}+\|(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\infty,1}+\|(\widehat{B}-B)^{\top}\widehat{\Sigma}B\|_{\infty,1}$
$\quad+\|B^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\infty,1}+\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\infty,1}.$
We proceed to bound each term separately. On the event $\mathcal{E}_{\pi}$, we find that
$\|D_{\widehat{\pi}}-D_{\pi}\|_{\infty,1}\leq\max_{\ell\in[L]}|\widehat{\pi}_{\ell}-\pi_{\ell}|\lesssim\sqrt{\log(n)\over nL}.$ (A.18)
Next, pick any $k,\ell\in[L]$. We bound from above
$\textrm{I}=(\widehat{B}_{\ell}-B_{\ell})^{\top}\widehat{\Sigma}(\widehat{B}_{k}-B_{k})\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}^{2},$ (A.19)
$\textrm{II}=(\widehat{B}_{\ell}-B_{\ell})^{\top}\widehat{\Sigma}B_{k}\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}\|\widehat{\Sigma}^{1/2}B_{k}\|_{2},$
$\textrm{III}=B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}.$
Note that
$\|\widehat{\Sigma}^{1/2}B_{k}\|_{2}\leq\left(B_{k}^{\top}\Sigma B_{k}+B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}\right)^{1/2}.$
Meanwhile Lemma 14 ensures that, on the event $\mathcal{E}_{\pi}$ in (A.1),
$\textrm{III}\leq\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\infty}\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL}$ (A.20)
with probability at least $1-n^{-3}$. The above two bounds together with (B.5) in Lemma 9 give that
$\textrm{II}\lesssim{\sqrt{\Delta_{\infty}}\over L}\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}.$ (A.21)
By combining (A.18), (A.19), (A.20) and (A.21) and taking the union bounds over $k,\ell\in[L]$, we conclude that
$\|\widehat{H}-H\|_{\infty,1}$ $\lesssim~{}\sqrt{\log(n)\over nL}+\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{L\log(n)\over n}$
$\qquad+\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}\sqrt{\Delta_{\infty}}+L\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}^{2}$
with probability at least $1-n^{-1}$. Invoking Theorem 2 and using (3.3) to simplify the expression complete the proof. ∎
A.4 Proof of Theorem 4
Proof.
Pick any $\ell\in[L]$. Start with
$|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\leq|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}|+|(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+2|\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right)|.$
We work on $\mathcal{E}_{\pi}$ in (A.1) throughout the rest of the proof. For the first term in $\widehat{G}_{\ell}(X)-G_{\ell}(X)$, invoking Lemma 16 with $k=\ell$ and $v=B^{*}_{\ell}$, using the fact that $B_{\ell}^{*\top}\Sigma_{w}B_{\ell}^{*}=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}=\Delta_{\ell\ell}$ and taking a union bounds over $\ell\in[L]$ gives that
$\mathbb{P}\left\{\left\{\max_{\ell\in[L]}|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}|\leq\sqrt{\Delta_{\infty}}\sqrt{aL\log(n)\over n}\right\}\cap\mathcal{E}_{\pi}\right\}\geq 1-2n^{-a}.$ (A.22)
To bound from above
$|\widehat{\mu}_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\leq|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|,$
using Cauchy-Schwarz inequality gives
$|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\leq\|\mu_{\ell}^{\top}\Sigma_{w}^{-1/2}\|_{2}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\leq\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}}.$ (A.23)
On the other hand, by using the same arguments of proving (A.3), we have that on the event $\mathcal{E}_{H}$
$|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\leq|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}|+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B(\widehat{\Omega}_{\ell}-\Omega_{\ell})|$
$\leq\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}\|\widehat{B}-B\|_{1,\infty}\|\widehat{\Omega}_{\ell}\|_{1}+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H\widehat{\Omega}-\mathbb{I}_{L})_{\ell}|$
$\leq\left(\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}\|\widehat{B}-B\|_{1,\infty}+\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{\infty}\delta\right)2\|\Omega_{\ell}\|_{1}.$
Here $\delta$ is defined in (A.13). Invoking (A.15), (A.22), (A.23) and Lemma 16 with $v={\bm{e}}_{j}$ and union bounds over $j\in[p]$ yields that, on $\mathcal{E}_{H}$ defined in (A.14),
$|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $~{}\lesssim L\Delta_{\infty}\sqrt{\|\Sigma_{w}\|_{\infty}}\sqrt{L\log(n\vee p)\over n}\|\widehat{B}-B\|_{1,\infty}$
$\qquad+L\delta\sqrt{\Delta_{\infty}^{3}}\sqrt{L\log(n)\over n}$
with probability at least $1-4(n\vee p)^{-2}$. Therefore, with probability at least $1-5(n\vee p)^{-2}$,
$\begin{split}|\widehat{\mu}_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|&~{}\lesssim~{}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}}+\delta\sqrt{\Delta_{\infty}^{3}}\sqrt{L^{3}\log(n)\over n}\\ &\qquad+\Delta_{\infty}\sqrt{L^{3}\log(n\vee p)\over n}\|\widehat{B}-B\|_{1,\infty}.\end{split}$ (A.24)
Finally, conditioning on $Y={\bm{e}}_{k}$, the Gaussian tail of $X\mid Y,{\mathbb{X}},{\mathbb{Y}}$ gives that, for all $t\geq 0$,
$\mathbb{P}\left\{|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\geq|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+t\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}~{}\Big{|}~{}Y={\bm{e}}_{k},{\mathbb{X}},{\mathbb{Y}}\right\}\leq 2e^{-t^{2}/2}.$
Choosing $t=\sqrt{4a\log(n)}$, taking the union bounds over $\ell\in[L]$ and unconditioning yield
$\max_{\ell\in[L]}|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\geq\max_{k,\ell\in[L]}|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+\max_{\ell\in[L]}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{4a\log(n)}$
$\geq 3\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{2,\infty}\sqrt{(a\vee\Delta_{\infty})\log(n)}$ (A.25)
with probability less than $2n^{-a}$. The last step uses (A.23). Collecting (A.2), (A.22), (A.24) and (A.4) and invoking Theorem 3 complete the proof. ∎
A.5 Proof of Theorem 5
For any $x\in\mathbb{R}^{p}$, write
$\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad g(x)=\operatorname*{arg\,min}_{\ell\in[L]}G_{\ell}(x).$
For any $t\geq 0$, define the event
$\mathcal{E}_{t}=\bigcap_{\ell\in[L]}\left\{\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|\leq t~{}\Big{|}~{}{\mathbb{X}},{\mathbb{Y}}\right\}.$ (A.26)
Proof.
By definition, we start with
$\quad\mathcal{R}(\widehat{g})$
$=\sum_{k\in[L]}\pi_{k}\Bigl{\{}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k\}\mid Y={\bm{e}}_{k}\right]-\mathbb{E}\left[\mathbbm{1}\{g^{*}(X)\neq k\}\mid Y={\bm{e}}_{k}\right]\Bigr{\}}$
$=\sum_{k\in[L]}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k,g^{*}(X)=k\}\mid Y={\bm{e}}_{k}\right]-\sum_{k\in[L]}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=k,g^{*}(X)\neq k\}\mid Y={\bm{e}}_{k}\right]$
$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y={\bm{e}}_{k}\right]-\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=k,g^{*}(X)=\ell\}\mid Y={\bm{e}}_{k}\right]$
$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\Bigl{\{}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y={\bm{e}}_{k}\right]-\pi_{\ell}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid L=\ell\right]\Bigr{\}}.$
Write $f_{k}(x)$ as the p.d.f. of $X=x\mid Y={\bm{e}}_{k}$ for each $k\in[L]$. We find that
[…]
where, for all $k,\ell\in[L]$,
$G^{(\ell|k)}(x):=\left(x-{\mu_{\ell}+\mu_{k}\over 2}\right)^{\top}\Sigma_{w}^{-1}(\mu_{\ell}-\mu_{k})+\log{\pi_{\ell}\over\pi_{k}}={1\over 2}\left(G_{k}(x)-G_{\ell}(x)\right).$
Observe that the event $\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\cap\mathcal{E}_{t}$ implies
$00>G_{z}^{(\ell|k)}(z)\overset{\mathcal{E}_{t}}{\geq}{1\over 2}\left(\widehat{G}_{k}(X)-\widehat{G}_{\ell}(X)\right)-t\geq-t,\quad\forall~{}t>$ 0 .
By the basic inequality $1+z\leq\exp(z)$ for all $z\in\mathbb{R}$, we obtain that, for any $t>0$,
$\mathcal{R}(\widehat{g})$ $\leq\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\left\{t\pi_{k}\mathbb{E}\left[\mathbbm{1}\{-t\leq G^{(\ell|k)}(X)\leq 0,\widehat{g}(X)=\ell\mid Y={\bm{e}}_{k}\}\right]+\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y={\bm{e}}_{k})\right\}$
$\leq t\sum_{k\in[L]}\pi_{k}\left\{\mathbb{P}\left(-t\leq\max_{\ell\in[L]\setminus\{k\}}G^{(\ell|k)}(X)\leq 0\mid Y={\bm{e}}_{k}\right)+(m-1)\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y={\bm{e}}_{k})\right\}$
$\leq t(m-1)\sum_{k\in[L]}\pi_{k}\max_{\ell\in[L]\setminus\{k\}}\mathbb{P}\left\{-t\leq G^{(\ell|k)}(X)\leq 0\mid Y={\bm{e}}_{k}\right\}+(m-1)\mathbb{P}(\mathcal{E}_{t}^{c})$
Note that
$G^{(\ell|k)}(X)\mid Y={\bm{e}}_{k}~{}\sim~{}N\left(-{1\over 2}\|\mu_{k}-\mu_{\ell}\|_{\Sigma_{w}}^{2}+\log\left(\pi_{\ell}\over\pi_{k}\right),~{}\|\mu_{k}-\mu_{\ell}\|_{\Sigma_{w}}^{2}\right).$
By the mean-value theorem, we conclude
$\mathcal{R}(\widehat{g})$ $\leq(m-1)t^{2}+(m-1)\mathbb{P}(\mathcal{E}_{t}^{c}).$
Choosing
$t=\Delta_{\infty}^{3}\sqrt{sL\log(n\vee p)\log(n)\over n}$
and invoking Theorem 4 complete the proof. ∎
Appendix B Technical lemmas
The following lemma provides a few useful facts on quantities related with $B$ in (1.13). Write
$\Omega:=D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}.$ (B.1)
Lemma 9.
With $B$ defined in (1.13), we have
${\mathbb{M}}^{\top}B={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}~{}\Omega^{-1},$ (B.2)
$H=D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B=\Omega^{-1},$ (B.3)
$B^{\top}\Sigma_{w}B=\Omega^{-1}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}~{}\Omega^{-1}.$ (B.4)
Consequently, for any $k,\ell\in[L]$, we have
$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}\leq\pi_{\ell}\left(1\wedge\pi_{\ell}\Delta_{\ell\ell}\right),$ (B.5)
$B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq\left(1\wedge{\pi_{\ell}^{2}\over\pi_{\min}}\Delta_{\ell\ell}\right),$ (B.6)
$\left|\mu_{k}^{\top}B_{\ell}\right|\leq\left(1\wedge\pi_{\ell}\sqrt{\Delta_{kk}\Delta_{\ell\ell}}\right).$ (B.7)
Proof.
From Lemma 1 and its proof, we know that
$B=B^{*}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)=\Sigma_{w}^{-1}{\mathbb{M}}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)$
so that
${\mathbb{M}}^{\top}B={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}-{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}B.$
Rearranging terms gives
${\mathbb{M}}^{\top}B$ $=\left(\mathbb{I}_{L}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}\right)^{-1}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}$
$=\mathbb{I}_{L}-\left(\mathbb{I}_{L}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}\right)^{-1}$
$=\mathbb{I}_{L}-D_{\pi}^{-1}\left(D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\right)^{-1}$ (B.8)
$={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\left(D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\right)^{-1}$
proving (B.2). By using the above identity (B.8), we have
$D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B=\left(D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\right)^{-1},$
proving (B.3). Since Lemma 1 and (B.3) imply
$B=B^{*}\Omega^{-1},$
(B.4) follows from
$B^{\top}\Sigma_{w}B=\Omega^{-1}B^{*\top}\Sigma_{w}B^{*}\Omega^{-1}=\Omega^{-1}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\Omega^{-1}.$
To prove (B.5) – (B.7), pick any $k,\ell\in[L]$. By using (2.6) twice and (2.5), we find that
$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}=\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}\leq\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$ (B.9)
yielding the second bound in (B.5). The other bound in (B.5) follows by observing that
[…]
For (B.6), on the one hand, similar arguments yield
$B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}\mu_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}.$
On the other hand, using (B.2) and (B.3) proves
$B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq{\bm{e}}_{\ell}^{\top}\Omega^{-1}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{2}\Omega^{-1}{\bm{e}}_{\ell}\leq 1.$
Finally, the last statement follows by noting that
$|\mu_{k}^{\top}B_{\ell}|^{2}\leq B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq 1$
and
$|\mu_{k}^{\top}B_{\ell}|=\pi_{\ell}|\mu_{k}^{\top}\Sigma^{-1}\mu_{\ell}|\leq\pi_{\ell}\sqrt{\mu_{k}^{\top}\Sigma^{-1}\mu_{k}}\sqrt{\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}}.$
The following lemmas contain deviation inequalities related with $\widehat{\mathbb{M}}-{\mathbb{M}}$ and $\widehat{\pi}-\pi$.
Lemma 10.
Let $v\in\mathbb{R}^{p}$ and $u\in\mathbb{R}^{L}$ be any fixed vectors. Under (1.1), for any $t\geq 0$, with probability at least $1-2e^{-t^{2}/2}$, we have
$\left|\sum_{k=1}^{L}{n_{k}\over n}v^{\top}(\widehat{\mu}_{k}-\mu_{k})u_{k}\right|\leq t\|u\|_{2}\sqrt{v^{\top}\Sigma_{w}v\over n}\sqrt{\max_{k}n_{k}\over n}.$
Proof.
Recall that, for any $k\in[L]$,
$\widehat{\mu}_{k}={1\over n_{k}}\sum_{i=1}^{n}X_{i}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}.$
We obtain
$\sum_{k=1}^{L}{n_{k}\over n}v^{\top}(\widehat{\mu}_{k}-\mu_{k})u_{k}$ $=v^{\top}\sum_{k=1}^{L}{1\over n}\sum_{i=1}^{n}\left(X_{i}-\mu_{k}\right)u_{k}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}$
$=v^{\top}\sum_{k=1}^{L}{1\over n}\sum_{i=1}^{n}\left(X_{i}-{\mathbb{M}}Y_{i}\right)u^{\top}Y_{i}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}$
$={1\over n}\sum_{i=1}^{n}v^{\top}\left(X_{i}-{\mathbb{M}}Y_{i}\right)u^{\top}Y_{i}.$
For $i\in[n]$, by conditioning on ${\mathbb{Y}}$, we know that $W_{i}:=X_{i}-{\mathbb{M}}Y_{i}$ are i.i.d. from $\mathcal{N}_{p}(0,\Sigma_{w})$, hence
${1\over n}\sum_{i=1}^{n}v^{\top}W_{i}u^{\top}Y_{i}~{}\Big{|}~{}{\mathbb{Y}}\sim\mathcal{N}\left(0,{v^{\top}\Sigma_{w}v\over n}{u^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}u\over n}\right).$
The claim follows by invoking the standard Gaussian tail probability bounds and after unconditioning. ∎
Lemma 11.
Let $u,v\in\mathbb{R}^{L}$ be any fixed vectors. Under (1.1), for any $t\geq 0$, with probability at least $1-2e^{-t/2}$,
$\left|\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})u_{k}v_{k}\right|\lesssim\sqrt{t\sum_{k=1}^{L}\pi_{k}u_{k}^{2}v_{k}^{2}\over n}+{t\|u\|_{\infty}\|v\|_{\infty}\over n}.$
Proof.
Observe that we can write
$\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})u_{k}v_{k}$ $=\sum_{k=1}^{L}{1\over n}\sum_{i=1}^{n}Y_{ik}u_{k}v_{k}-\sum_{k=1}^{L}\pi_{k}u_{k}v_{k}$
$={1\over n}\sum_{i=1}^{n}Z_{i}-\pi_{k}u_{k}v_{k}$
Here
$Z_{i}=\sum_{k=1}^{L}Y_{ik}u_{k}v_{k}$
are independent, bounded random variables (as $|Z_{i}|\leq\|u\|_{\infty}\|v\|_{\infty}$) with mean
$\mathbb{E}[Z_{i}]=\sum_{k=1}^{L}\pi_{k}u_{k}v_{k}$
and variance
${\rm Var}(Z_{i})={\rm Var}(y_{i}^{\top}{\rm diag}(u)v)\leq\sum_{k=1}^{L}\pi_{k}v_{k}^{2}u_{k}^{2}$
for $i\in[n]$. The proof follows after a straightforward application of Bernstein’s inequality for bounded random variables. ∎
The following lemma controls the difference
$\widehat{\Sigma}_{w}-\Sigma_{w}:={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}-\Sigma_{w}.$
Lemma 12.
Let $u,v\in\mathbb{R}^{p}$ be any fixed vectors. Under model (1.1), for any $t\geq 0$, with probability at least $1-2e^{-t}$, one has
$\left|u^{\top}\left(\widehat{\Sigma}_{w}-\Sigma_{w}\right)v\right|\lesssim\sqrt{u^{\top}\Sigma_{w}u}\sqrt{v^{\top}\Sigma_{w}v}\left(\sqrt{t\over n}+{t\over n}\right)$
Proof.
Start with
$\widehat{\Sigma}_{w}-\Sigma_{w}$ $={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}-\Sigma_{w}$
$={1\over n}\sum_{i=1}^{n}\left[(X_{i}-{\mathbb{M}}Y_{i})(X_{i}-{\mathbb{M}}Y_{i})^{\top}-\Sigma_{w}\right].$
Recall that, conditioning ${\mathbb{Y}}$, $W_{i}=X_{i}-{\mathbb{M}}Y_{i}$ for $i\in[n]$ are independent, $\mathcal{N}_{p}(0,\Sigma_{w})$. The result follows from standard concentration inequalities for the quadratic term of Gaussian random vectors. ∎
B.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
Recall that $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$ and the definition of the event $\mathcal{E}_{\pi}$ is given in (A.1).
Lemma 13.
Under model (1.1) and 1,assume $L\log(n)\leq n$. On the event $\mathcal{E}_{\pi}$, for any fixed $v_{1},v_{2}\in\mathbb{R}^{p}$ and any $t\geq 0$, with probability at least $1-8e^{-t}$, one has
$\left|v_{1}^{\top}(\widehat{\Sigma}-\Sigma)v_{2}\right|$ $\lesssim~{}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}\left(\sqrt{t\over n}+{t\over n}\right)$
$\quad+\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{\mathbb{M}}^{\top}v_{1}\|_{\infty}\|{\mathbb{M}}^{\top}v_{2}\|_{\infty}{t\over n}$
$\quad+\left({\|{\mathbb{M}}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}+{\|{\mathbb{M}}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\right)\sqrt{t\over n}.$
Proof.
By definition, we have
$\widehat{\Sigma}$ $={1\over n}{\mathbb{X}}^{\top}{\mathbb{X}}={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}X_{i}X_{i}^{\top}$
$={1\over n}\sum_{k=1}^{L}\left[\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}+n_{k}(\widehat{\mu}_{k}\mu_{k}^{\top}+\mu_{k}\widehat{\mu}_{k}^{\top})-n_{k}\mu_{k}\mu_{k}^{\top}\right]$
$=\widehat{\Sigma}_{w}+\sum_{k=1}^{L}{n_{k}\over n}\mu_{k}\mu_{k}^{\top}+\sum_{k=1}^{L}{n_{k}\over n}\left[(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}+\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}\right]$
where we write
$\widehat{\Sigma}_{w}:={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}.$
By further using the decomposition of $\Sigma$ in (2.6) and $\widehat{\pi}_{k}=n_{k}/n$ for $k\in[L]$, we find that
$\widehat{\Sigma}-\Sigma$ $=\widehat{\Sigma}_{w}-\Sigma_{w}+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})\mu_{k}\mu_{k}^{\top}+\sum_{k=1}^{L}\widehat{\pi}_{k}\left[(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}+\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}\right].$
The first term $\widehat{\Sigma}_{w}-\Sigma_{w}$ is bounded in Lemma 12. We bound the second term by invoking 1 and Lemma 11 with $u={\mathbb{M}}^{\top}v_{1}$ and $v={\mathbb{M}}^{\top}v_{2}$. This yields
$\left|\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v_{1}^{\top}\mu_{k}\mu_{k}^{\top}v_{2}\right|$ $\lesssim\sqrt{t\sum_{k=1}^{L}\pi_{k}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over n}+{t\|{\mathbb{M}}^{\top}v_{1}\|_{\infty}\|{\mathbb{M}}^{\top}v_{2}\|_{\infty}\over n}$
$\lesssim\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{\mathbb{M}}^{\top}v_{1}\|_{\infty}\|{\mathbb{M}}^{\top}v_{2}\|_{\infty}{t\over n}$ (B.11)
with probability at least $1-2e^{-t}$. Furthermore, on the event $\mathcal{E}_{\pi}$, 1 and $L\log(n)\leq n$ imply
$\widehat{\pi}_{k}\lesssim\pi_{k}+\sqrt{\log(n)\over nL}\lesssim{1\over L},\quad\forall\ k\in[L],$
and after we invoke Lemma 10 twice with $v=v_{1},u={\mathbb{M}}^{\top}v_{2}$ and $v=v_{2},u={\mathbb{M}}^{\top}v_{1}$, respectively, we get that
$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}v_{2}\lesssim\|{\mathbb{M}}^{\top}v_{1}\|_{2}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{\mathbb{M}}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n},$ (B.12)
$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}v_{2}\lesssim\|{\mathbb{M}}^{\top}v_{2}\|_{2}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{\mathbb{M}}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}$ (B.13)
hold with probability at least $1-4e^{-t}$. Collecting all three bounds completes the proof. ∎
As an immediate application of Lemma 13, we have the following bounds on the sup-norm and operator norm of the matrix $B^{\top}(\widehat{\Sigma}-\Sigma)B$.
Lemma 14.
Under conditions of Lemma 13, assume $\Delta_{\infty}\gtrsim 1$. On the event $\mathcal{E}_{\pi}$, the following holds with probability at least $1-n^{-2}$,
$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\infty}\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL},$
$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\lesssim{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}.$
Proof.
For the sup-norm bound, it suffices to bound from above $|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ for any $k\in[L]$. Invoking Lemma 13 with $v_{1}=v_{2}=B_{k}$ and $t=\log(n)$ together with (B.5), (B.6) and (B.7) of Lemma 9 gives that, with probability at least $1-n^{-2}$,
$|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ $\lesssim B_{k}^{\top}\Sigma_{w}B_{k}\sqrt{\log(n)\over n}+\|{\mathbb{M}}^{\top}B_{k}\|_{\infty}\sqrt{B_{k}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
$\quad+\|{\mathbb{M}}^{\top}B_{k}\|_{\infty}^{2}{\log(n)\over n}+2\sqrt{B_{k}^{\top}\Sigma_{w}B_{k}B_{k}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
$\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)\sqrt{\log(n)\over nL^{2}}+\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL}+\left(1\wedge{\Delta_{\infty}\over L}\right)^{2}{\log(n)\over n}$
$\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL}.$
The last step uses $\Delta_{\infty}\gtrsim 1$ to collect terms. This proves the first claim. Regarding the operator norm bound, a standard discretization argument gives
$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\leq 3\max_{u\in\mathcal{N}_{L}(1/2)}u^{\top}B^{\top}(\widehat{\Sigma}-\Sigma)Bu$
where $\mathcal{N}_{L}(1/3)$ denotes the $(1/3)$-epsilon net of $\{u\in\mathbb{R}^{L}:\|u\|_{2}=1\}$. Note that $|\mathcal{N}_{L}(1/3)|\leq 7^{L}$. Then invoking Lemma 13 with $t=CL\log(n)$ and using
$\|B^{\top}\Sigma_{w}B\|_{\rm op}\leq{\Delta_{\rm op}\over L(L+\Delta_{\rm op})},$
$\|B^{\top}{\mathbb{M}}\|_{\rm op}\leq{\Delta_{\rm op}\over L+\Delta_{\rm op}},$
$\|{\mathbb{M}}^{\top}Bu\|_{\infty}\leq\sqrt{\Delta_{\infty}\Delta_{\rm op}\over L(L+\Delta_{\rm op})}$
deduced from Lemma 9 give that, with probability at least $1-8n^{-CL+L\log(7)}\geq 1-n^{-L}$,
$u^{\top}B^{\top}(\widehat{\Sigma}-\Sigma)Bu$ $\lesssim u^{\top}B^{\top}\Sigma_{w}Bu\sqrt{L\log(n)\over n}+\|{\mathbb{M}}^{\top}Bu\|_{\infty}\sqrt{u^{\top}B^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}Bu}\sqrt{\log(n)\over n}$
$\quad+\|{\mathbb{M}}^{\top}Bu\|_{\infty}^{2}{L\log(n)\over n}+2\sqrt{u^{\top}B^{\top}\Sigma_{w}Buu^{\top}B^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}Bu}\sqrt{\log(n)\over n}$
$\lesssim{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}$
holds uniformly over $u\in\mathcal{N}_{L}(1/3)$. We also use $\Delta_{\infty}\gtrsim 1$ to simplify expressions in the last step above. This completes the proof. ∎
Appendix C Auxiliary lemmas
The following lemmas are proved in Bing and Wegkamp (2023).
Lemma 15.
Assume $\pi_{\min}\geq 2\log n/n$ for some sufficiently large constant $C$. Then, for any $k\in[L]$
$\mathbb{P}\left\{|\widehat{\pi}_{k}-\pi_{k}|<\sqrt{16\pi_{k}(1-\pi_{k})\log n\over n}\right\}\geq 1-n^{-2}.$
Furthermore, if $\pi_{\min}\geq C\log n/n$ for some sufficiently large constant $C$, then
$\mathbb{P}\left\{c\pi_{k}\leq\widehat{\pi}_{k}\leq c^{\prime}\pi_{k}\right\}\geq 1-n^{-2}.$
Lemma 16.
Suppose that model (1.1) holds. For any deterministic vector $v\in\mathbb{R}^{p}$ and $k\in[L]$, for all $t>0$,
$\mathbb{P}\left\{\left|v^{\top}(\widehat{\mu}_{k}-\mu_{k})\right|\geq t\sqrt{v^{\top}\Sigma_{w}v\over n_{k}}\right\}\leq 2e^{-{t^{2}/2}}.$