### v2

# Linear Discriminant Regularized Regression

Abstract
Linear Discriminant Analysis (LDA) is an important classification approach. Its simple linear form makes it easy to interpret and it is capable to handle multi-class responses. It is closely related to other classical multivariate statistical techniques, such as Fisher’s discriminant analysis, canonical correlation analysis and linear regression. In this paper we strengthen its connection to multivariate response regression by characterizing the explicit relationship between the discriminant directions and the regression coefficient matrix. This key characterization leads to a new regression-based multi-class classification procedure that is flexible enough to deploy any existing structured, regularized, and even non-parametric, regression methods. Moreover, our new formulation is amenable to analysis: we establish a general strategy of analyzing the excess misclassification risk of the proposed classifier for all aforementioned regression techniques. As applications, we provide complete theoretical guarantees for using the widely used $\ell_{1}$-regularization as well as for using the reduced-rank regression, neither of which has yet been fully analyzed in the LDA context. Our theoretical findings are corroborated by extensive simulation studies and real data analysis.
Keywords: Multi-class classification, discriminant analysis, high-dimensional data, regularization, multivariate response regression, dimension reduction.
1 Introduction
This paper introduces a new regression-based approach to multi-class, high-dimensional linear discriminant analysis (LDA). The main premise that leads to LDA is that we have a sample of $n$ i.i.d. copies $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ of $(X,Y)$, with the label $Y\in\{{\bm{e}}_{1},\ldots,{\bm{e}}_{L}\}$ taking values in $L\geq 2$ classes and the conditional distributions of the feature vector $X\in\mathbb{R}^{p}$ satisfy
$X\mid Y={\bm{e}}_{\ell}~{}\sim~{}\mathcal{N}_{p}\left(\mu_{\ell},\Sigma_{w}\right),\qquad\text{for all }\ell\in[L]:=\{1,\ldots,L\}.$ (1.1)
Here $\{{\bm{e}}_{1},\ldots,{\bm{e}}_{L}\}$ are the canonical basis vectors in $\mathbb{R}^{L}$ and the within-class covariance matrix $\Sigma_{w}$ is assumed to be strictly positive definite. For any $i\in[n]$, we say that observation $X_{i}$ belongs to class $\ell\in[L]$ if its label $Y_{i}={\bm{e}}_{\ell}$, which one may simply write as $Y_{i}=\ell$. Our insistence on the perhaps unusual notation ${\bm{e}}_{\ell}$ will become transparent later when we formulate the estimation of the optimal hyperplane as a multivariate response regression problem. Since we can always subtract the marginal mean of $X$, we assume $\mathbb{E}(X)=0_{p}$ throughout this paper mainly for simplicity, and furthermore that the class probabilities $\pi_{\ell}=\mathbb{P}\{Y={\bm{e}}_{\ell}\}$ are strictly positive for all $\ell\in[L]$. The Bayes rule under model (1.1) is to classify a new feature $x\in\mathbb{R}^{p}$ as
$g^{*}(x):=\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (1.2)
See, for instance, Izenman (2008). Estimation of this classifier based on $(X_{i},Y_{i})$ with $i\in[n]$ becomes challenging in (i) the high-dimensional setting $p>n$ and (ii) the multi-class setting $L>2$ that allows for $L\to\infty$ as $n\to\infty$.
Few approaches consider (ii) but notable exceptions are Abramovich and Pensky (2019); Nibbering and Hastie (2022). We recall various approaches to deal with (i). We will use the following notation. Write
$\pi=(\pi_{1},\ldots,\pi_{L})^{\top},\quad D_{\pi}={\rm diag}(\pi)\in\mathbb{R}^{L\times L},\quad{M}=(\mu_{1},\ldots,\mu_{L})\in\mathbb{R}^{p\times L}$ (1.3)
and denote the unconditional covariance matrix of $X$ by
$\Sigma={\rm Cov}(X)=\Sigma_{w}+{M}D_{\pi}{M}^{\top}.$ (1.4)
We call $\Sigma_{w}$ the within-class covariance matrix and $\Sigma_{b}:=\Sigma-\Sigma_{w}={M}D_{\pi}{M}^{\top}$ the between-class covariance matrix. The latter has rank less than $L$ due to $\mathbb{E}(X)=0_{p}$.
In high-dimensional settings $p>n$, Cai and Liu (2011) assume that the direction $\Sigma_{w}^{-1}(\mu_{2}-\mu_{1})$ is sparse in the binary case $L=2$, and propose a Dantzig-type procedure by approximately solving the equation $\Sigma_{w}\beta^{*}=\mu_{2}-\mu_{1}$ for $\beta^{*}$ to estimate this direction. Other procedures based on the same or similar equations include, for instance, Qiao et al. (2009); Shao et al. (2011); Fan et al. (2012) and Tibshirani et al. (2002); Fan and Fan (2008); Cai and Zhang (2019); Chen and Sun (2022) for $L>2$. However, it isn’t always reasonable to assume that such directions are sparse and the procedure is less attractive compared to regression-based approaches discussed below shortly. Witten and Tibshirani (2011) alternatively use the fact that the Bayes rule (1.2) is equivalent with Fisher’s discriminant rule
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}F{F^{\top}}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (1.5)
Here the columns of the matrix $F=(F_{1},\ldots,F_{K})\in\mathbb{R}^{p\times K}$, for some $K<L$, are implicitly defined via, for $1\leq k\leq K$,
$\begin{split}F_{k}:=\operatorname*{arg\,max}_{\beta\in\mathbb{R}^{p}}~{}\beta^{\top}\Sigma_{b}\beta\quad\text{subject to}&~{}\beta^{\top}\Sigma_{w}\beta=1,~{}\beta^{\top}\Sigma_{w}F_{i}=0,~{}~{}\forall~{}i<k.\end{split}$ (1.6)
Witten and Tibshirani (2011) develop methodology tailored to the high-dimensional scenarios with a diagonal matrix $\Sigma_{w}$ and sparse columns of $F$. However, the implicit formulation (1.6) of the matrix $F$ makes it extremely difficult to obtain the global solution and to derive theoretical properties of the regularized estimator and the resulting classifier.
More closely related to our approach is optimal scoring which dates back to De Leeuw et al. (1976); Young et al. (1978) and regains attention in the LDA setting, see Hastie et al. (1995, 1994); Clemmensen et al. (2011); Gaynanova (2020) and the references therein. By writing the response matrix and feature matrix as
${\mathbf{Y}}=(Y_{1},\ldots,Y_{n})^{\top}\in\{0,1\}^{n\times L},\quad{\mathbf{X}}=(X_{1},\ldots,X_{n})^{\top}\in\mathbb{R}^{n\times p},$ (1.7)
optimal scoring is a regression-based approach that aims to solve a sequence of optimization problems: for $1\leq k\leq K<L$,
$\begin{split}\min_{\ \theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbf{Y}}\theta-{\mathbf{X}}\beta\|_{2}^{2}\quad\text{s.t.}~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\theta=n,~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k.\end{split}$ (1.8)
Here $\widehat{\theta}_{i}$ is the solution to (1.8) in the $i$th iteration. In the low-dimensional case ($p<n$), Hastie et al. (1995) show that the solutions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$ from (1.8) can be computed from canonical correlation analysis (CCA), and are further related, via CCA, with the matrix $F$ obtained from the sample analogue of (1.6). As a result, classification can be based on the dimension reduction directions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$. In the high-dimensional setting, regularization in (1.8) is needed. For instance, Hastie et al. (1995) consider
$\begin{split}\min_{\theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbf{Y}}\theta-{\mathbf{X}}\beta\|_{2}^{2}+\beta^{\top}\Omega\beta\quad\text{s.t.}~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\theta=n,~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k,\end{split}$ (1.9)
for a chosen $p\times p$ symmetric, positive definite matrix $\Omega$, and show that the global solution to (1.9) can be computed. A possible choice of $\Omega=\lambda\mathbf{I}_{p}$ for some $\lambda>0$ leads to the ridge penalty (Campbell, 1980; Friedman, 1989). To accommodate more general regularization, Gaynanova (2020) studies the matrix formulation of (1.9),
$\begin{split}\min_{\Theta,\ B}&~{}\|{\mathbf{Y}}\Theta-{\mathbf{X}}B\|_{F}^{2}+\text{pen}(B)\quad\text{s.t.}~{}\Theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\Theta=n\mathbf{I}_{K},\quad\Theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}1_{n}=0_{K},\end{split}$ (1.10)
and proves that the global solution to (1.10) can be computed, provided pen$(B)$ satisfies $\text{pen}(BR)=\text{pen}(B)$ for any orthogonal $K\times K$ matrix $R$. Gaynanova (2020) focuses on the group-lasso penalty $\|B\|_{1,2}=\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ and analyzes the resulting classifier. However, it is not clear how the analysis can be extended to other penalized estimation. Moreover, the requirement that pen($B$) is orthogonal invariant excludes the familiar lasso penalty $\|B\|_{1}=\sum_{j,k}|B_{jk}|$ as well as its extensions such as the elastic net penalty. That being said, for penalties that are not orthogonal invariant, there is no guarantee of computing the global solution to (1.10), let alone any theoretical property of the resulting estimator.
Key to our approach is the following reformulation of the problem. We show in Lemma 1 of Section 2 that the discriminant directional matrix under model (1.1)
$B^{*}=\Sigma_{w}^{-1}{M}\in\mathbb{R}^{p\times L}$ (1.11)
and the regression matrix
$B:=\Sigma^{-1}\Sigma_{XY}:=[{\rm Cov}(X)]^{-1}{\rm Cov}(X,Y)\in\mathbb{R}^{p\times L}$ (1.12)
not only share the same column space, but have a closed-form connection, to wit,
$B^{*}$ $=BH^{-1}$ (1.13)
for some invertible $L\times L$ matrix $H$ with explicit expression. This connection is novel and distinguishes our approach from the current literature. Although the fact that columns of $B$ and $B^{*}$ share the same span is known (see, for instance, Hastie et al. (1995); Ye (2007); Lee and Kim (2015); Nie et al. (2022) and the references therein), their closed-form connection given by (1.13) has not been discovered, and a thorough theoretical investigation on the corresponding classifier has not been conducted. To the best of our knowledge, only in the binary case ($L=2$), Mai et al. (2012) leverages sparse linear regression between an appropriately encoded label vector and the feature matrix ${\mathbf{X}}$ for classification and proves consistency. Extension of such procedure to handle multi-class responses with high-dimensional features has been a longstanding open problem. In view of (1.13), we propose to estimate the discriminant direction matrix $B^{*}$ by first estimating the regression matrix $B$ as
$\widehat{B}=\operatorname*{arg\,min}_{B\in\mathbb{R}^{p\times L}}~{}{1\over n}\|{\mathbf{Y}}-{\mathbf{X}}B\|_{F}^{2}+\text{pen}(B),$ (1.14)
and then estimating $H^{-1}$ based on $\widehat{B}$ via its explicit expression in Lemma 1. The advantages of our approach as well as our main contributions are three-fold:
  1. (a)
    Akin to optimal scoring, our approach in (1.14) is regression-based, which, as pointed out in Hastie et al. (1995, 1994), is often easier to compute and more amenable to extensions to regularized and non-parametric regression as well as model selection than canonical correlation analysis, Fisher’s discriminant rule and the aforementioned procedures based on approximately solving the equation $\Sigma_{w}B^{*}=M$ for $B^{*}$. We can also easily tap into the large literature on estimation of the high-dimensional regression matrix $B\in\mathbb{R}^{p\times L}$ in (1.14).
  2. (b)
    After computing $\widehat{B}$ from (1.14), we only need to estimate the inverse of an $L\times L$ matrix to estimate the discriminant direction matrix $B^{*}$ rather than that of a $p\times p$ matrix $\Sigma_{w}$ as in the classical LDA context. Our approach in Section 2.2 uses the explicit form of $H$ in Lemma 1 and enjoys computational advantage as well as numerical stability, compared to the existing approaches. In Section 2.3, we further offer an alternative way to estimate the Bayes rule that performs Fisher’s discriminant analysis on low-dimensional transformations $\widehat{B}^{\top}x\in\mathbb{R}^{L}$ of the original features $x\in\mathbb{R}^{p}$. This leads to a practically useful method, which could extract fewer discriminant directions than those in $B^{*}$, for both downstream analysis and visualization.
  3. (c)
    We are able to quantify the errors of estimating both $B$ and $B^{*}$ as well as to provide theoretical guarantees for the resulting classifier, thanks to (1.13). Concretely, in Section 3.1 and Section 3.2 we provide a general strategy for analyzing the excess misclassification risk of the proposed classifier, which is valid for a generic estimator $\widehat{B}$ of $B$. While various choices of $\widehat{B}$ are feasible in our approach, more specifically in (1.14), we apply the general result to two popular estimators: the lasso estimator in Section 3.3 and the reduced rank estimator in Section 3.4. The former also contrasts our approach with Cai and Zhang (2019) and Witten and Tibshirani (2011) which assume sparsity of $B^{*}$ and $F$, respectively. The latter is particularly suitable when the rank of the discriminant direction matrix $B^{*}$ is low and the number of classes $L$ is moderate / large. For instance, our earlier work Bing and Wegkamp (2023) considers a multi-class classification setting with high-dimensional features $X\in\mathbb{R}^{p}$ that follow a factor model $X=AZ+W$ with unobserved (latent) features $Z\in\mathbb{R}^{r}$ with $r<p$ and random noise $W$, that is independent of both $Z$ and the label $Y$. In this case, the discriminant direction matrix $B^{*}$ has rank no greater than $r$ when $r<L$. The procedure in Bing and Wegkamp (2023), however, is different than the one proposed here. We refer to Remark 1 in Section 2.2 for more discussion.
Finally, we provide a thorough simulation study in Section 4 to corroborate our theoretical findings. In Section 5, we report our findings in several real data studies. The Appendix contains most of the proofs and additional simulation study.
Notation.
For any numbers $a,b\in\mathbb{R}$, we write $a\vee b=\max\{a,b\}$. For any vector $v$ and $1\leq q\leq\infty$, we use $\|v\|_{q}$ to denote the standard $\ell_{q}$-norm. For any positive integer $d$, we write $[d]:=\{1,\ldots,d\}$. For any matrix $Q\in\mathbb{R}^{d_{1}\times d_{2}}$, any $i\in[d_{1}]$ and $j\in[d_{2}]$, we write $Q_{i\cdot}$ for its $i$th row and $Q_{j}$ for its $j$th column. For norms, we write $\|Q\|_{\infty}=\max_{i,j}|Q_{ij}|$, $\|Q\|_{F}^{2}=\sum_{i,j}Q_{ij}^{2}$, $\|Q\|_{1}=\sum_{i,j}|Q_{ij}|$, $\|Q\|_{1,2}=\sum_{i}\|Q_{i\cdot}\|_{2}$ and $\|Q\|_{\rm op}=\sup_{v:\|v\|_{2}=1}\|Qv\|_{2}$. For any symmetric, semi-positive definite matrix $Q\in\mathbb{R}^{d\times d}$, we use $\lambda_{1}(Q),\lambda_{2}(Q),\ldots,\lambda_{d}(Q)$ to denote its eigenvalues in non-increasing order. For any sequences $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ if there exists some constant $C>0$ such that $a_{n}\leq Cb_{n}$, and $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$.
2 Methodology
We discuss our classification procedure in this section. Our goal is to estimate the Bayes rule in (1.2) which, for a given $x\in\mathbb{R}^{p}$, requires estimation of the discriminant functions
$G_{\ell}(x):=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2x^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}),\qquad\text{for all }\ell\in[L].$ (2.1)
This, in turn, requires to estimate the discriminant directional matrix $B^{*}=\Sigma_{w}^{-1}{M}$. The following lemma establishes the connection between $B^{*}$ and the regression matrix $B$ in (1.12). Recall the definitions of $\pi$, $D_{\pi}$ and ${M}$ from (1.3) as well as $\Sigma$ from (1.4).
Lemma 1.
Under model (1.1), we have
$B^{*}=BH^{-1}$ (2.2)
for a matrix $H\in\mathbb{R}^{L\times L}$ that can be inverted and satisfies
$H=D_{\pi}-B^{\top}\Sigma B=D_{\pi}-D_{\pi}{M}^{\top}B=D_{\pi}-B^{\top}{M}D_{\pi}.$ (2.3)
Proof.
From the identity
$\Sigma_{XY}=\mathbb{E}(XY^{\top})-\mathbb{E}(X)(\mathbb{E}(Y))^{\top}={M}D_{\pi}$ (2.4)
we see that
$B=\Sigma^{-1}{M}D_{\pi}.$ (2.5)
Next, using
$\Sigma=\Sigma_{w}+\Sigma_{b}=\Sigma_{w}+{M}D_{\pi}{M}^{\top},$ (2.6)
the Woodbury matrix identity yields
$\Sigma_{w}^{-1}=\Sigma^{-1}+\Sigma^{-1}{M}\left(\mathbf{I}_{L}-D_{\pi}{M}^{\top}\Sigma^{-1}{M}\right)^{-1}D_{\pi}{M}^{\top}\Sigma^{-1}.$
It follows that
$B^{*}$ $=\Sigma_{w}^{-1}{M}$
$=\Sigma^{-1}{M}+\Sigma^{-1}{M}\left(\mathbf{I}_{L}-D_{\pi}{M}^{\top}\Sigma^{-1}{M}\right)^{-1}D_{\pi}{M}^{\top}\Sigma^{-1}{M}$
$=\Sigma^{-1}{M}\left(\mathbf{I}_{L}-D_{\pi}{M}^{\top}\Sigma^{-1}{M}\right)^{-1}$
$=\Sigma^{-1}{M}D_{\pi}\left(D_{\pi}-D_{\pi}{M}^{\top}\Sigma^{-1}{M}D_{\pi}\right)^{-1}$
$=B\left(D_{\pi}-D_{\pi}{M}^{\top}B\right)^{-1}$
which completes the proof. ∎
Lemma 1 suggests that estimation of $B^{*}$ could be done by first estimating the matrix $B$ and then estimating $H^{-1}$ according to (2.3). We discuss these two steps in details in the next two sections.
2.1 Estimation of the regression matrix
To accommodate high-dimensional data, we estimate $B$ via regularized regression methods in (1.14). Choices of the penalty term $\text{pen}(B)$ depend on the concrete problem at hand. For instance, when entries of $B$ exhibit certain smoothness structure such as in some imaging or audio applications (Hastie et al., 1995), one could choose $\text{pen}(B)={\rm tr}(B^{\top}\Omega B)$ for some pre-specified “roughness” penalty matrix $\Omega$. In this paper we mainly focus on the following two types of structural estimation of $B$.
Sparsity-based structural estimation.
If we believe that only a subset of the original features are important for predicting the label, then sparsity structures of the regression matrix $B$ are reasonable assumptions. If the subset of features, that are pertinent for predicting the label, varies across different levels, then the entry-wise lasso penalty is appropriate. The corresponding matrix estimator $\widehat{B}$ is defined in (1.14) with the following lasso penalty (Tibshirani, 1996)
$\text{pen}(B)$ $=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{1}$ (2.7)
for some tuning parameter $\lambda>0$. In Section 3.3 we provide a complete analysis of this estimator. If there exist many features that are not predictive for all levels of the response label, then the group-lasso penalty $\text{pen}(B)=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ (Yuan and Lin, 2006) is a more suitable choice. In this case, one can retain a much smaller subset of important features and achieve feature selection. Furthermore, one can also consider either a row-wise or a column-wise elastic net penalty (Zou and Hastie, 2005) if the features tend to (highly) correlate with each other. Our numerical results in Sections 4 and 5 include both the lasso penalty and the elastic net penalty.
Low-rank structural estimation.
Different from sparsity, if the goal is to construct a low-dimensional subset of linear combinations of the original feature for classification purpose, then a low-rank penalty in (1.14) could be more appropriate. For example, if the matrix of conditional means ${M}\in\mathbb{R}^{p\times L}$ has a reduced-rank $r\ll\min\{p,L\}$, then Lemma 1 implies that ${\rm rank}(B)\leq r$. In this case estimation of $B\in\mathbb{R}^{p\times L}$ in (1.14) could be done by using the following reduced-rank penalty (see, for instance, Izenman (1975))
$\text{pen}(B)=\lambda~{}{\rm rank}(B)$ (2.8)
for some tuning parameter $\lambda>0$. A full theoretical analysis of using (2.8) is presented in Section 3.4. Some alternative penalties include, but are not limited to, the reduced-rank penalty plus the matrix ridge penalty in Mukherjee and Zhu (2011), the nuclear norm penalty in Koltchinskii et al. (2011), adaptive nuclear norm penalty in Chen et al. (2013) as well as the generalized cross-validation rank penalty in Bing and Wegkamp (2019).
Remark 1.
We conclude this section by giving another concrete example under which $B$ has low rank. In Bing and Wegkamp (2023) the authors consider the model $X=AZ+W$ with latent factors $Z\in\mathbb{R}^{r}$ for some $r\ll p$ and some general additive noise $W$, that is independent of both $Z$ and the label $Y$. Conditioning on $Y={\bm{e}}_{\ell}$ for all $\ell\in[L]$, the low-dimensional factors $Z$ are assumed to follow Gaussian distributions with different conditional means $\alpha_{\ell}\in\mathbb{R}^{r}$ but the same within-class covariance matrix $\Sigma_{z}\in\mathbb{R}^{r\times r}$. When $W$ is also Gaussian, this model is a sub-model of (1.1) with $\mu_{\ell}=A\alpha_{\ell}$ and $\Sigma_{w}=A\Sigma_{z}A^{\top}+{\rm Cov}(W)$. When $r\leq\min\{p,L\}$, it is easy to verify that ${\rm rank}(B)\leq r$. The two-step approach in Bing and Wegkamp (2023) uses the unsupervised Principal Component Analysis (PCA) to reduce the feature dimension and then base classification on the reduced dimension. It is entirely different from our approach in Section 2.2 but connected to the one in Section 2.3 in the way that our approach in Section 2.3 is also a two-step procedure but reduces the feature space in the first step in a supervised way via the regression step in (1.14). Numerical comparison of both procedures are also included in our simulation study in Appendix A.
2.2 Estimation of the discriminant matrix and the Bayes rule
The discriminant functions in (2.1) that form the basis of the Bayes classification rule (1.2) necessitate that we estimate the matrix of the conditional means ${M}$, the vector of prior probabilities $\pi$ as well as the discriminant directional matrix $B^{*}$. We adopt the following standard estimators
$\widehat{M}={\mathbf{X}}^{\top}{\mathbf{Y}}({\mathbf{Y}}^{\top}{\mathbf{Y}})^{-1},\qquad D_{\widehat{\pi}}={\rm diag}(\widehat{\pi})={1\over n}{\mathbf{Y}}^{\top}{\mathbf{Y}}$ (2.9)
to estimate ${M}$ and $D_{\pi}$, respectively. Recall from Lemma 1 that $B^{*}=BH^{-1}$. Since the estimator $\widehat{B}$ of $B$ is computed in (1.14), in view of the closed-form of $H$ in (2.3), we propose to estimate $H$ by the plug-in estimator
$\widehat{H}=D_{\widehat{\pi}}-{1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}{\mathbf{X}}\widehat{B}={1\over n}\left({\mathbf{Y}}^{\top}{\mathbf{Y}}-\widehat{B}^{\top}{\mathbf{X}}^{\top}{\mathbf{X}}\widehat{B}\right).$ (2.10)
Note that ${\mathbf{X}}\widehat{B}$ is simply the in-sample fit of the regression in (1.14). Computation of ${\mathbf{X}}\widehat{B}$ thus can be done efficiently even in non-parametric regression settings, for instance, when the columns of ${\mathbf{X}}$ consist of basis expansions. Now, using $B^{*}=BH^{-1}$, we can further estimate $B^{*}$ by
$\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}.$ (2.11)
Here for any matrix $M$, $M^{+}$ denotes its moore-penrose inverse. Finally, in view of (2.1), we estimate the discriminant functions by
$\widehat{G}_{\ell}(x)=\widehat{\mu}_{\ell}^{\top}\widehat{B}^{*}_{\ell}-2x^{\top}\widehat{B}^{*}_{\ell}-2\log(\widehat{\pi}_{\ell}),\qquad\text{for all $\ell\in[L]$. }$ (2.12)
The corresponding classifier is
$\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad\text{ for any $x\in\mathbb{R}^{p}$}.$ (2.13)
Theoretically, we show in Theorem 5 of Section 3.2 that $\widehat{H}$ is invertible with overwhelming probability under the same mild condition that is required for consistent classification. In Theorem 6 of Section 3.3 and Theorem 8 of Section 3.4, the condition is verified for the lasso estimator, and the reduced rank estimator, of $B$, respectively. Moreover, it is worth mentioning that the inverse of $\widehat{H}$ always exists when any $\ell_{2}$-norm related penalty is deployed in (1.14).
2.3 Classification based on reduced dimensions
Since in practice it is often desirable to extract fewer discriminant directions than those in $B^{*}$, we discuss in this section an alternative procedure for this purpose, which estimates the Bayes rule by also using $\widehat{B}$ from (1.14) but in a different way.
We start with the fact (its proof is stated in Section B.1 for completeness) that the Bayes rule in (1.2) is equivalent to
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}(x-\mu_{\ell})-2\log(\pi_{\ell})$ (2.14)
which, by Lemma 1, further equals to
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B(B^{\top}\Sigma_{w}B)^{+}B^{\top}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (2.15)
The formulation in (2.15) has the following two-step interpretation:
  1. (1)
    Transform the original feature $x\in\mathbb{R}^{p}$ to $z:=B^{\top}x$, with the latter belonging to a subspace of $\mathbb{R}^{L}$;
  2. (2)
    Perform linear discriminant analysis on the space of $z$.
The second point follows by noticing that
$C_{w}:=B^{\top}\Sigma_{w}B\in\mathbb{R}^{L\times L}$
is indeed the within-class covariance matrix of $Z=B^{\top}X$ under model (1.1).
Due to the fact that $C_{w}$ is rank deficient with ${\rm rank}(C_{w})<L$ and for the purpose of extracting fewer discriminant directions, we adopt the equivalent formulation of (2.15) from the Fisher’s perspective in the sequel. Write the between-class covariance matrix of $Z=B^{\top}X$ as
$C_{b}:=B^{\top}{M}D_{\pi}{M}^{\top}B.$
For a given reduced dimension $1\leq K<L$, we can find $K$ discriminant directions in the space of $Z$ by solving, for each $k\in[K]$, the optimization problem
$\begin{split}\alpha_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~{}\alpha^{\top}C_{b}~{}\alpha\quad\text{ s.t.}~{}\alpha^{\top}C_{w}~{}\alpha=1,~{}\alpha^{\top}C_{w}~{}\alpha_{i}=0,~{}~{}\forall~{}i<k.\end{split}$
The resulting classification rule based on these $K$ discriminant directions is
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B\sum_{k=1}^{K}\alpha_{k}\alpha_{k}^{\top}B^{\top}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (2.16)
In practice, after computing $\widehat{B}$ from (1.14), we propose to estimate $C_{b}$ and $C_{w}$ by
$\widehat{C}_{b}=\widehat{B}^{\top}\widehat{M}D_{\widehat{\pi}}\widehat{M}^{\top}\widehat{B}={1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}P_{{\mathbf{Y}}}{\mathbf{X}}\widehat{B},\qquad\widehat{C}_{w}={1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}(\mathbf{I}_{n}-P_{{\mathbf{Y}}}){\mathbf{X}}\widehat{B},$
with $P_{{\mathbf{Y}}}={\mathbf{Y}}({\mathbf{Y}}^{\top}{\mathbf{Y}})^{-1}{\mathbf{Y}}^{\top}$. Subsequently, the estimated discriminant directions $\widehat{\alpha}_{k}$, for $1\leq k\leq K$, are computed by solving
$\begin{split}\widehat{\alpha}_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~{}\alpha^{\top}\widehat{C}_{b}~{}\alpha\quad\text{ s.t.}~{}\alpha^{\top}\widehat{C}_{w}~{}\alpha=1,~{}\alpha^{\top}\widehat{C}_{w}~{}\widehat{\alpha}_{i}=0,~{}~{}\forall~{}i<k,\end{split}$ (2.17)
By similar reasoning for solving the Fisher’s discriminant analysis in (1.6), we propose to solve the optimization problem (2.17) via eigen-decomposing
$[\widehat{C}_{w}^{+}]^{1\over 2}\widehat{C}_{b}[\widehat{C}_{w}^{+}]^{1\over 2}.$
The eigenvalues of the above matrix can also be used to determine the upper bound for $K$. For instance, $K$ should be chosen no greater than the rank of the above matrix.
Remark 2 (Numerical stability of computing (2.17)).
The procedure in (2.17) computes Fisher’s discriminant directions $(\widehat{\alpha}_{1},\ldots,\widehat{\alpha}_{K})$ in
$\text{span}({\mathbf{X}}\widehat{B})\subset\text{span}({\mathbf{X}})$
of dimension less than $L$. Consequently, as long as
${\rm rank}((\mathbf{I}_{n}-P_{{\mathbf{Y}}}){\mathbf{X}}\widehat{B})={\rm rank}({\mathbf{X}}\widehat{B}),$ (2.18)
we do not suffer numerically in solving (2.17), in stark contrast with the algorithmic issue of solving (1.6) (or a penalized version thereof) in the original $p$-dimensional space. As pointed out by Wu et al. (2015), the latter leads to meaningless solutions with the objective function equal to infinity. Finally, we remark that the requirement in (2.18) is easy to satisfy as ${\rm rank}({\mathbf{X}}\widehat{B})\leq L$ and ${\rm rank}(\mathbf{I}_{n}-P_{{\mathbf{Y}}})=n-L$ when we observe at least one data point per class.
3 Theoretical guarantees
In this section we provide a unified theory for analyzing the classifier $\widehat{g}$ in (2.13) based on estimating the regression coefficient matrix. In Section 3.1 we start with full generality and bound the excess risk of $\widehat{g}$ by a quantity that is related with the error of estimating $B^{*}$. The results there hold for any estimator $\widehat{B}^{*}$ of $B^{*}$. In Section 3.2 we focus on the proposed estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ and provide a unified analysis of its estimation error that is valid for any generic estimator $\widehat{B}$. Finally we apply the general result to two particular choices of $\widehat{B}$, the lasso estimator in Section 3.3 and the reduced-rank estimator in Section 3.4.
3.1 A reduction scheme for bounding the excess risk
Let $(X,Y)$ be a new pair from (1.1) that is independent of the training data ${\bm{D}}:=\{{\mathbf{X}},{\mathbf{Y}}\}$. For any classifier $\widehat{g}:\mathbb{R}^{p}\to[L]$ in (2.13), its excess risk relative to the error rate of the Bayes rule in (1.2) is defined as
$\mathcal{R}(\widehat{g}):=\mathbb{P}\{Y\neq\widehat{g}(X)\mid{\bm{D}}\}-\mathbb{P}\{Y\neq g^{*}(X)\}.$
Note that we use the identification $\{Y=k\}:=\{Y={\bm{e}}_{k}\}$ in the above definition. Our analysis of $\mathcal{R}(\widehat{g})$ requires the following condition on the priors $\pi=(\pi_{1},\ldots,\pi_{L})$.
Assumption 1.
There exists some absolute constants $0<c\leq C<\infty$ such that
${c\over L}\leq\min_{k\in[L]}\pi_{k}\leq\max_{k\in[L]}\pi_{k}\leq{C\over L}.$ (3.1)
This is a common regularity assumption even in binary classification problems. We may relax Assumption 1 to $\min_{k\in[L]}\pi_{k}\geq c\log(n)/n$ in our analysis (see, for instance, Lemma 21 of Appendix D) at the cost of more technical proofs and a less transparent presentation.
The following proposition states that for the purpose of bounding $\mathcal{R}(\widehat{g})$, it suffices to study the estimation error of each discriminant function, that is, $\widehat{G}_{\ell}(X)-G_{\ell}(X)$ for $\ell\in[L]$. The proof essentially follows the argument of proving Theorem 12 in Bing and Wegkamp (2023) but with modifications to capture the sum of squared errors of estimating all discriminant functions.
Proposition 2.
Under model (1.1) with 1, for any sequences $t_{1},\ldots,t_{L}$ (that possibly depend on ${\bm{D}}$), with probability equal to one, one has
$\mathcal{R}(\widehat{g})~{}\leq~{}C\sum_{\ell=1}^{L}t_{\ell}^{2}+L\sum_{\ell=1}^{L}\mathbb{P}\left\{|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\geq t_{\ell}\mid{\bm{D}}\right\}.$
Proof.
See Section B.2. ∎
2 is valid for any estimator $\widehat{G}_{\ell}$ of the discriminant function, in particular, for our proposed $\widehat{G}_{\ell}$ in (2.12). The choice of $(t_{1},\ldots,t_{L})$ depends on the estimation error of $|\widehat{G}_{\ell}(X)-G_{\ell}(X)|$. We therefore proceed to analyze
$\widehat{G}_{\ell}(X)-G_{\ell}(X)$ $=(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}+(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})+2\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right)$ (3.2)
for each $\ell\in[L]$. It is worth mentioning that in this section we allow $\widehat{B}^{*}$ in the above display to be any estimator of $B^{*}$ whereas $\widehat{\mu}_{\ell}$ and $\widehat{\pi}_{\ell}$ are given in (2.9). Since the variance of the first term in (3.2) depends on $\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$, we define
$\Delta_{\infty}:=\max_{\ell\in[L]}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}.$
This quantity essentially reflects the largest Mahalanobis distance in the feature space between any two label classes as can be seen from the relation (under $\mathbb{E}(X)=0_{p}$)
${1\over 4}\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\leq\Delta_{\infty}\leq\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}).$
1, the Gaussian tail of $X\mid Y$ and the independence between $X$ and $({\mathbf{X}},{\mathbf{Y}})$ in (3.2) entail the following upper bound of $|\widehat{G}_{\ell}(X)-G_{\ell}(X)|$.
Theorem 3.
Under model (1.1) with 1, assume $\Delta_{\infty}\geq 1$ and $L\log(n)\leq n$. With probability at least $1-3n^{-1}$, we have, for all $\ell\in[L]$,
$\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|~{}\lesssim~{}\sqrt{\Delta_{\infty}}\sqrt{L\log(n)\over n}$ $+\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+\log(n)}$
$+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|.$
Proof.
See Section B.3. ∎
Condition $L\log(n)\leq n$ together with 1 ensures that $\widehat{\pi}_{\ell}\asymp\pi_{\ell}$ for all $\ell\in[L]$. The reasonable condition $\Delta_{\infty}\geq 1$ requires that the largest separation among all pairs of conditional distributions does not vanish. It is assumed to simplify the presentation, and our analysis can be easily extended to the case $\Delta_{\infty}\to 0$ as $n\to\infty$.
Combining 2 with Theorem 3 immediately yields the following corollary. Denote by $\mathbb{E}_{{\bm{D}}}$ the expectation taken with respect to the training data ${\bm{D}}$ only. For any estimator $\widehat{B}^{*}$ of $B^{*}$, define the following (random) quantity
$\mathcal{Q}(\widehat{B}^{*}):=\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}^{2}(\Delta_{\infty}+\log n)+\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}.$ (3.3)
Corollary 4.
Under model (1.1) with 1, assume $\Delta_{\infty}\geq 1$. For any estimator $\widehat{B}^{*}$ of $B^{*}$, the corresponding classifier $\widehat{g}$ in (2.13) satisfies
$\mathbb{E}_{\bm{D}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}\mathbb{E}_{{\bm{D}}}\left[\min\left\{{\Delta_{\infty}}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*}),~{}1\right\}\right].$
Proof.
See Section B.4. ∎
4 ensures that bounding $\mathcal{R}(\widehat{g})$ can be reduced to controlling $\mathcal{Q}(\widehat{B}^{*})$ for any estimator $\widehat{B}^{*}$ of $B^{*}$. In particular, 4 is valid for any existing approach based on estimating $B^{*}$. In the next section we provide a unified analysis of $\mathcal{Q}(\widehat{B}^{*})$ for our proposed estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ based on a generic $\widehat{B}$. As shown later, the first term in (3.3) is typically the dominating term.
3.2 A unified analysis of the estimation of $B^{*}$ based on generic regression coefficient estimation
In this section we provide more explicit bounds of $\mathcal{Q}(\widehat{B}^{*})$ in (3.3) for the estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ with $\widehat{H}$ obtained from (2.10). Our results in this section are valid for any estimator $\widehat{B}$ of $B$ that satisfies the following assumption.
Assumption 2.
There exists some deterministic sequences $\omega_{1}$ and $\omega_{2}$ such that the following holds with probability at least $1-n^{-1}$,
$\max\left\{n^{-1/2}\|{\mathbf{X}}(\widehat{B}-B)\|_{F},~{}\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}\right\}\leq\omega_{2},\quad\|\widehat{B}-B\|_{1,2}\leq\omega_{1}.$
We need the following regularity conditions on $\Sigma_{w}$. It assumes the largest entry of $\Sigma_{w}$ is bounded from above and the smallest eigenvalue of $\Sigma_{w}$ is bounded away from zero.
Assumption 3.
There exists some positive constants $0<c\leq C<\infty$ such that $c<\lambda_{p}(\Sigma_{w})\leq\|\Sigma_{w}\|_{\infty}\leq C$.
Theorem 5.
Under model (1.1) with Assumptions 1 & 3, assume $\Delta\asymp 1$ and $L\log(n)\leq cn$ for some small constant $c>0$. For any $\widehat{B}$ satisfying 2 with $\omega_{2}\sqrt{L}\leq c,$ with probability at least $1-2n^{-1}$,
the matrix $\widehat{H}$ in (2.10) is invertible,
and
$\mathcal{Q}(\widehat{B}^{*})\lesssim~{}L^{2}{\log^{2}(n)\over n}+\omega_{2}^{2}L^{2}\log(n)+\omega_{1}^{2}L^{3}{\log(n\vee p)\over n}.$ (3.4)
Proof.
See Section B.5. ∎
Theorem 5 relates $\mathcal{Q}(\widehat{B}^{*})$ to the estimation error of $\widehat{B}$ in both the Frobenius norm and the $\ell_{1}/\ell_{2}$-norm. In the bound of (3.4), the second term on the right-hand-side is usually the leading term and its rate depends on specific choice of $\widehat{B}$. We provide concrete examples in Section 3.3 for the lasso estimator and in Section 3.4 for the reduced-rank estimator.
Condition $\omega_{2}\sqrt{L}\leq c$ ensures that, with high probability, $\lambda_{K}(\widehat{H})\geq\lambda_{K}(H)/2$ holds so that $\widehat{H}$ is invertible. Furthermore, the bound in (3.4) suggests that the condition $\omega_{2}\sqrt{L}\leq c$ is also needed for consistent classification.
3.3 Application to estimation of $B$ with entry-wise $\ell_{1}$-regularization
In this section we focus on the case where the columns of $B$ are (potentially) sparse and allow for different sparsity patterns across its columns. Therefore, for some positive integer $s\leq p$, we consider the following parameter space of $B$:
$\Theta_{s}:=\left\{B\in\mathbb{R}^{p\times L}:\max_{\ell\in[L]}\|B_{\ell}\|_{0}\leq s\right\}.$
For ease of presentation, we simply use the largest number of non-zero entries among all columns of $B$. However, our analysis can be easily extended to $\|B_{\ell}\|_{0}\leq s_{\ell}$ by allowing different $s_{\ell}$ for $\ell\in[L]$. Our analysis also needs the following mild regularity condition on the matrix ${M}$ of conditional means.
Assumption 4.
There exists an absolute constant $C<\infty$ such that $\|{M}\|_{\infty}\leq C$.
The following theorem establishes explicit rates of $\omega_{2}$ and $\omega_{1}$ in 2 for the regression matrix estimator $\widehat{B}$ in (1.14) using the lasso-penalty in (2.7).
Theorem 6.
Under model (1.1) with Assumptions 1, 3 and 4, assume that there exists some sufficiently small constant $c>0$ such that $B\in\Theta_{s}$ for some $1\leq s\leq p$ with $(s\vee L)\log(n\vee p)\leq c~{}n.$ Let the estimator $\widehat{B}$ be obtained from (1.14) with any $\lambda$ satisfying
$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (3.5)
and some finite, large enough constant $C>0$. Then, with probability at least $1-3n^{-1}$, 2 holds for
$\omega_{2}\asymp\lambda\sqrt{sL},\qquad\omega_{1}\asymp\lambda sL.$
Proof.
See Section B.6. ∎
Our proof is in fact based on a weaker condition than 3 by replacing $\lambda_{p}(\Sigma_{w})$ with the Restricted Eigenvalue condition (RE) on $\Sigma_{w}$. The two main difficulties in our proofs are (a) to control $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}$, and (b) to bound from below the restricted eigenvalue of the Gram matrix $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$. For (a), as pointed out in Gaynanova (2020), the difficulty in this model is elevated relative to the existing analysis in sparse linear regression models due to the fact ${\mathbf{Y}}$ is not a linear model in ${\mathbf{X}}$. We establish a sharp control of $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}$ in Lemma 11. For (b), the existing result in Rudelson and Zhou (2012) is not readily applicable as rows of ${\mathbf{X}}$ are generated from a mixture of Gaussians. In Lemma 12 we establish a uniform bound of $v^{\top}(\widehat{\Sigma}-\Sigma)v$ over $s$-sparse vectors $v$ which, in conjunction with the reduction arguments in Rudelson and Zhou (2012), proves (b). Our analysis in this paper allows the number of classes $L$ to grow with the sample size $n$
Gaynanova (2020) considers a group-lasso regularized regression similar to (1.14) and derives the bound of the $\ell_{2}$-norm $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}^{\prime}-{\mathbf{X}}B^{\prime})\|_{2}$ for some different response matrix ${\mathbf{Y}}^{\prime}$ and different target matrix $B^{\prime}$. Their proof is different and in particular requires that $L$ is fixed, independent of $n$.
In the rest of the paper, we assume that $\lambda$ is chosen as the order specified in (3.5). In practice, one can simply choose $\lambda$ via cross-validation, for instance, via the built-in function of the R-package glmnet.
As an immediate corollary of 4, Theorem 5 and Theorem 6, we have the following upper bound of the misclassification rate of $\widehat{g}$ based on the entry-wise lasso estimator with $\lambda$ chosen as the rate in (3.5).
Corollary 7.
Under the conditions in Theorem 6, assume that $\Delta_{\infty}\asymp 1$. Then
$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}\min\left\{{sL^{2}\over n}\log^{2}(n\vee p),~{}1\right\}.$
Our results allow both the sparsity level $s$ and the number of classes $L$ to grow with the sample size $n$. The existing literature (Cai and Zhang, 2019; Gaynanova, 2020) that analyze the excess risk only consider the case $L\asymp 1$, while Abramovich and Pensky (2019) does allow $s$, $L$ and $n$ to grow, but studies the risk $\mathbb{P}\{Y\neq g(X)\}$ instead of the excess risk $\mathcal{R}(g)$. Guarantees on the excess risk are stronger since they together with the Bayes error, which is oftentimes easy to calculate, imply guarantees of the risk.
7 focuses on the statistically most interesting case $\Delta_{\infty}\asymp 1$. Otherwise, if the pairwise separation between different classes grows to infinity, the classification problem becomes much easier and we have super-fast rates of convergence (see, for instance, Cai and Zhang (2019); Bing and Wegkamp (2023)). Indeed, provided $\min_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\asymp\max_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}),$ using the arguments in the proof of Bing and Wegkamp (2023, Theorem 12), our analysis can be modified easily to show that
$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}{sL^{2}\over n}\log^{2}(n\vee p)e^{-c\Delta_{\infty}}+n^{-C_{0}}$
for some arbitrary large constant $C_{0}>0$ and some constant $c>0$. The first term on the right in the above rate is exponentially fast in $\Delta_{\infty}\to\infty$.
3.4 Application to estimation of $B$ via reduced-rank regression
We provide analysis for the classifier that estimates the regression coefficient matrix $B$ by reduced-rank regression. Specifically, let $\widehat{B}$ be obtained from (1.14) with $\text{pen}(B)=\lambda~{}{\rm rank}(B)$ for some tuning parameter $\lambda>0$. The following theorem establishes the rates of $\omega_{2}$ and $\omega_{1}$ for which 2 holds. Write $r={\rm rank}(B)$.
Theorem 8.
Under model (1.1) with Assumptions 1 and 3, assume $(p+L)\log(n)\leq c~{}n$ for some sufficiently small constant $c>0$. For the estimator $\widehat{B}$ in (1.14) with any
$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\rm op}(1+\Delta_{\infty})}\sqrt{(p+L)\log(n)\over nL}$ (3.6)
and some finite, large enough constant $C>0$, with probability at least $1-2n^{-1}$, 2 holds for
$\omega_{2}\asymp\sqrt{\lambda r},\qquad\omega_{1}\asymp\sqrt{\lambda rp}.$
Proof.
The proof is deferred to Section B.7. ∎
For the reduced-rank estimator, establishing its in-sample prediction risk can be done without any condition on the design matrix (see, for instance, Bunea et al. (2011); Giraud (2011)). However, since 2 is also related with the out-of-sample prediction risk, we need the smallest eigenvalue of the Gram matrix $\widehat{\Sigma}$ to be bounded away from zero. The inequality $(p+L)\log(n)\leq cn$ is assumed for this purpose. In case $p$ is large, it is recommended that an additional ridge penalty be added in $\text{pen}(B)$ to alleviate the singularity issue of $\widehat{\Sigma}$ (Mukherjee and Zhu, 2011). Another main ingredient of our proof of Theorem 8 is to control $\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}$ in Lemma 13. It requires non-standard analysis for the same reason mentioned before that ${\mathbf{Y}}$ is not linearly related with ${\mathbf{X}}$.
Combining 4, Theorem 5 and Theorem 8 immediately yields the following guarantee on the excess risk of the classifier using the reduced-rank estimator of $B$.
Corollary 9.
Under the conditions in Theorem 8, assume $\Delta_{\infty}\asymp 1$ and $\|\Sigma_{w}\|_{\rm op}\lesssim 1$. Then
$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}\min\left\{{(p+L)rL\over n}\log^{2}(n),~{}1\right\}.$
Note that for the $p\times L$ matrix $B$ with rank $r$, its effective number of parameters is exactly $(p+L)r$ (Izenman, 1975) which could be much smaller than $pL$ when $r\ll(p\wedge L)$. 9 suggests that the classifier using the reduced-rank estimator has benefit when $r$ is relatively small comparing to $p$ and $L$.
The same remark as the one at the end of Section 3.3 applies, and our analysis can be extended to $\Delta_{\infty}\to\infty$, where the rate in 9 becomes exponentially fast in $\Delta_{\infty}$.
4 Simulation study
In our simulation study, we evaluate our proposed procedure in Section 2.2 that we term as Linear Discriminant Regularized Regression (LDRR). We also include the procedure in Section 2.3 that is based on Fisher’s discriminant rule after estimating $B$ (called LDRR-F). For estimating $B$ from (1.14), in the sparse scenarios of Section 4.1, we consider both the lasso penalty (2.7) (L1) and the elastic net penalty (L1+L2), $\text{pen}(B)=\lambda\sum_{\ell=1}^{L}[\alpha\|B_{\ell}\|_{1}+(1-\alpha)/2\|B_{\ell}\|_{2}^{2}]$, for some $\alpha\in[0,1]$ and $\lambda>0$ chosen by cross-validation from the R-package glmnet. In the low-rank scenarios of Section 4.2, we consider both the reduced-rank penalty (2.8) (RR) and the reduced-rank penalty plus the ridge penalty (RR+L2), $\text{pen}(B)=\lambda[\alpha~{}{\rm rank}(B)+(1-\alpha)/2\|B\|_{F}^{2}]$, with tuning parameters $\alpha\in[0,1]$ and $\lambda>0$ chosen by cross-validation.
We compare our proposed methods with the nearest shrunken centroids classifier (PAMR) of Tibshirani et al. (2002), the shrunken centroids regularized discriminant analysis (RDA) method of Guo et al. (2007) and the $\ell_{1}$-penalized linear discriminant (PenalizedLDA) method of Witten and Tibshirani (2011). These methods are available in R-packages pamr, rda, penalizedLDA, and we use their default methods for selecting tuning parameters. Finally, the oracle procedure that uses the true values of ${M}$, $\Sigma_{w}$, and $\pi$ in computing the discriminant functions $G_{k}(x)$ in (2.1) serves as our benchmark.
4.1 Sparse scenarios
We generate the data from the following mean-shift scenario. A similar setting is also used in Witten and Tibshirani (2011). Specifically, for any $\ell\in[L]$ and its corresponding mean vector $\mu_{\ell}\in\mathbb{R}^{p}$, we sample $[\mu_{\ell}]_{j}$, for $5(\ell-1)\leq j\leq 5\ell$, independently from $N(0,2^{2})$ and set the other entries of $\mu_{\ell}$ to 0. Regarding the within-class covariance matrix, we set $\Sigma_{w}=\sigma^{2}W$ where $W$ is generated by independently sampling its diagonal elements from Uniform(1,3), and setting its off-diagonal elements as $W_{ij}=\sqrt{W_{ii}W_{jj}}\rho^{|i-j|}$, for all $i\neq j$. The quantity $\sigma>0$ controls the overall noise level while the coefficient $\rho\in[0,1]$ controls the within-class correlation among the features. We generate the class probabilities $\pi=(\pi_{1},\ldots,\pi_{L})^{\top}$ as follows. For a given $\alpha\geq 0$, we set $\pi_{\ell}=\nu_{\ell}^{\alpha}/(\sum_{i=1}^{L}\nu_{i}^{\alpha})$ for all $\ell\in[L]$. Here $\nu_{1},\ldots,\nu_{L}$ are independent draws from the Uniform$(0,1)$. The quantity $\alpha$ controls the imbalance of each class size: a larger $\alpha$ corresponds to more imbalanced class sizes. At the other extreme, $\alpha=0$ corresponds to the balanced case $\pi=(1/L)1_{L}$. In the following, we examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{100,300,500,700,900\}$; the feature dimension $p\in\{100,300,500,700,900\}$; the number of classes $L\in\{2,5,8,11,14\}$; the correlation coefficient $\rho\in\{0,0.2,0.4,0.6,0.8\}$; the noise parameter $\sigma\in\{0.6,0.9,1.2,1.5,1.8\}$; the imbalance of the class probabilities $\alpha\in\{0,0.4,0.8,1.2,1.6\}$.
For all settings, we fix $n=300$, $p=500$, $L=5$, $\rho=0.6$, $\sigma=1$ and $\alpha=0$ when they are not varied. We use $500$ test data points to compute the misclassification errors of each procedure, and for each setting, we consider $50$ repetitions and report their averaged misclassification errors in Figure 1.
From Figure 1, it is clear that our procedures consistently outperform other methods across all settings. We observe that the performance of our algorithms improves as $n$ increases, $L$ decreases and $\sigma$ decreases. Moreover, neither the imbalance of the class probabilities nor the ambient dimension $p$ seems to affect the misclassification error of our algorithms. These findings are in line with our theory in Section 3.3. It is a bit surprising to see that the classification error of our benchmark decreases as the correlation coefficient $\rho$ keeps increasing after $0.4$. Nevertheless, both our algorithms and RDA adapt to this situation. Finally, we find that both LDRR and LDRR-F as well as their lasso and elastic net variants have very comparable performance in all scenarios. This is as expected as the data is simulated according to model (1.1), although in practice the Fisher’s discriminant analysis (LDRR-F) with the elastic net penalty could have more robust performance, as revealed in our real data analysis in the next section.
Figure 1: The averaged misclassification errors of each algorithm for varying each parameter one at a time.
4.2 Low-rank scenarios
In this section we evaluate our proposed procedure when the rank of the conditional mean matrix $M$ is low and hence $B$ is also low-rank. We consider a general within-class covariance matrix $\Sigma_{w}$ here, and defer to Appendix A for the factor model mentioned in Remark 1 where $\Sigma_{w}$ has approximately low-rank.
Specifically, let the within-class covariance matrix $[\Sigma_{w}]_{ij}=\rho^{|i-j|}$ with $\rho=0.6$. To generate $M$, for some scalar $\eta>0$ and some randomly generated orthogonal matrix $A\in\mathbb{R}^{p\times r}$ with $A^{\top}A=\mathbf{I}_{r}$, we set $M=\eta A\alpha$ with entries of $\alpha\in\mathbb{R}^{r\times L}$ generated i.i.d. from $N(0,32/r)$. The quantity $\eta$ controls the magnitude of separation between the conditional means.
We examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{300,500,700,900,1100\}$; the feature dimension $p\in\{100,200,300,400,500\}$; the number of classes $L\in\{5,15,25,35,45\}$; the scalar $\eta\in\{0.3,0.6,0.9,1.2,1.5\}$. For each setting, we fix $n=1000,L=10,r=3,p=100$ and $\eta=1$ when they are not varied. We set $n=500$ when we vary $\eta$. Figure 2 depicts the misclassification errors of each algorithm averaged over 50 repetitions.
From Figure 2 we see that the proposed approaches of using the reduced-rank penalty generally outperform the other methods. As $L$ gets larger with $r$ fixed, the benefit of using reduced-rank penalty becomes more visible, in line with 9. When the dimension ($p$) increases and gets closer to the sample size, the performance of LDRR(RR) deteriorates quickly while LDRR(RR+L2) still performs the best. For relative large $p$, incorporating the ridge penalty is beneficial as it improves the estimation of $B$ when the sample covariance matrix $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$ becomes ill-conditioned.
Figure 2: The averaged misclassification errors of each algorithm.
5 Real data analysis
In this section we evaluate the performance of our approach and compare it with other algorithms on three high-dimensional biological data sets and one low-dimensional football dataset.
The Ramaswamy dataset has $n=198$ samples and each sample has features of $p=16,063$ gene expressions and belongs to one of the $L=14$ distinct cancer subtypes (Ramaswamy et al., 2001). The SRBCT dataset (Khan et al., 2001) has $n=83$ samples and consists of microarray analyses of childhood cancer tumors, more specifically, small round blue cell tumors (SRBCT). These tumors are categorized into $L=4$ classes: Burkitt lymphoma, Ewing sarcoma, neuroblastoma, and rhabdomyosarcoma. Each sample has $p=2,308$ gene expressions as features. The third dataset has $n=62$ samples and consists of $L=3$ distinct lymphoma subtypes. These lymphoma are categorized into diffuse large B-cell lymphoma, follicular lymphoma and chronic lymphocytic leukemia. Each sample has $p=4026$ gene expressions as its features. We used the preprocessed Lymphoma data in Dettling (2004). The Football dataset from kaggle [^1] https://www.kaggle.com/datasets/vivovinco/20222023-football-player-stats/data contains 2022 – 2023 football player stats per 90 minutes. We excluded categorical data such as the player’s nationality, name, team name, and league from our prediction. We used age, year of birth and performance data $(p=118)$ to predict the $(n=2689)$ football player’s position $(L=10)$.
For all four data sets, the features are centered to have zero mean. Each dataset is randomly divided with 75% of the samples to serve as the training data and the remaining 25% serves as the test data. This randomization was repeated 50 times and the averaged misclassification errors of each algorithm are reported in Table 1.
Table 1: The averaged misclassification errors (in percentage) and standard errors
LDRR-F(L1+L2) LDRR-F(L12+L2) LDRR-F(RR+L2) PenalizedLDA PAMR RDA
Ramaswamy 30.7(0.051) 14.4(0.054) 22.6(0.048) 40.0(0.058) 30.4(0.058) 16.7(0.05)
SRBCT 0.7(0.022) 0.3(0.011) 0.1(0.007) 4.4(0.052) 1.0(0.019) 0.0(0.0)
Lymphoma 0.3(0.012) 0.5(0.017) 0.4(0.015) 1.9(0.034) 1.3(0.025) 1.3(0.025)
Football 31.5(0.013) 32.2(0.011) 30.5(0.008) 52.6(0.008) 43.8(0.016) 31.9(0.014)
Given the similar performance in Section 4 between LDRR and LDRR-F as well as their variants of using the lasso and the elastic-net penalty, we only compare LDRR-F using the elastic net penalty (L1+L2) and the reduced rank plus ridge penalty (RR+L2) with PenalizedLDA, PAMR and RDA in our real data analysis due to its robustness against model misspecification. We also include LDRR-F that uses the group-lasso penalty plus the ridge penalty (L12+L2) for comparison. From Table 1, we see that our procedure LDRR-F is close to have the best performance across all datasets. Choice of the best penalty depends on the data structure. For instance, when the data has the group-sparsity structure such as in Ramaswamy data set, the classifier using the group-lasso penalty has better performance; when the sparsity structure varies across the response levels such as in Lymphoma, the classifier using the elastic-net penalty performs better; when the data meets the low rank structure, the benefit of the classifier using the reduced-rank estimator is apparent. For the Football dataset with $p=118$ and $L=10$, the selected rank is found to be $\widehat{r}=3$. Furthermore, in Figure 3, we show that LDRR-F with $K=2$ in (2.16) can be used for visualization in the 2D space of the first two discriminant directions. The algorithm RDA has comparable performance in all settings, but it cannot be used for low-dimensional representation / visualization.
Figure 3: Visualization of Ramaswamy, SRBCT, Lymphoma and Football data points in the space of the first two discriminant vectors.
Acknowledgements
Wegkamp is supported in part by the National Science Foundation grant NSF DMS-2210557.
Appendix A Additional simulation study when the conditional mean matrix is low-rank and $\Sigma_{w}$ is approximately low-rank
In this setting, we consider the setting in Remark 1 and also compare with the method (PCLDA) in Bing and Wegkamp (2023). As mentioned in Remark 1, PCLDA is a two-step procedure similar to LDRR-F but with the first step done in a unsupervised way by using PCA. We generate the data according to model (1.1) with $M=\eta A\alpha$ as in Section 4.2 and $\Sigma_{w}=\eta^{2}A\Sigma_{z}A^{\top}+{\rm Cov}(W)$ where $[\Sigma_{z}]_{ij}=(0.6)^{|i-j|}$ for $i,j\in[r]$ and ${\rm Cov}(W)$ is generated the same as $\Sigma_{w}$ in Section 4.1 with $\rho=0.2$ and its diagonal elements from Uniform(0,1). The quantity $\eta$ represents the signal-to-noise ratio (SNR) between the low-dimensional signal $\eta AZ$ and the noise $W$. As pointed out in Bing and Wegkamp (2023), classification becomes easier as $\eta$ increases.
We examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{300,500,700,900,1100\}$; the feature dimension $p\in\{100,200,300,400,500\}$; the number of classes $L\in\{5,15,25,35,45\}$; the scalar $\eta\in\{0.5,1.0,2.0,4.0,6.0\}$. For all settings, we fix $n=1000,L=10,r=3,p=100$ and $\eta=2$ when they are not varied. We set $n=500$ when we vary $\eta$. Figure 4 depicts the misclassification errors of each algorithm averaged over 50 repetitions.
From Figure 4 we have similar findings as in the low-rank setting in Section 4.2 that the proposed approaches of using the reduced rank penalty outperform other methods when $n$ is much larger than $p$. In the last setting when we vary the SNR $\eta$, it is interesting to notice that PCLDA, the procedure that reduces the dimension by PCA in its first step, has comparable performance to the LDRR-F(RR+L2) when $\eta$ is sufficiently large ($\eta\geq 2$ in these settings). This aligns with the minimax optimality of PCLDA derived in Bing and Wegkamp (2023) in the regime of moderate / large $\eta$. But when $\eta$ is small such as $\eta<2$, LDRR-F has benefit by reducing the dimension in a supervised way.
Figure 4: The averaged misclassification errors of each algorithm.
Appendix B Proofs
Throughout the proofs, we will use the following notation. Define
$\Delta={M}^{\top}\Sigma_{w}^{-1}{M}$
and write
$\Delta_{\infty}=\max_{\ell\in[L]}\Delta_{\ell\ell},\qquad\Delta_{\rm op}=\|\Delta\|_{\rm op}.$
Further write
$\pi_{\min}=\min_{\ell\in[L]}\pi_{\ell},\qquad\pi_{\max}=\max_{\ell\in[L]}\pi_{\ell}.$
The following event will be frequently used in our proofs. Define the event
$\mathcal{E}_{\pi}=\left\{\max_{k\in[L]}\left|\widehat{\pi}_{k}-\pi_{k}\right|\leq C\sqrt{\log(n)\over nL}\right\}$ (B.1)
for some absolute constant $C>0$. Lemma 21 in Appendix D ensures that $\mathbb{P}(\mathcal{E}_{\pi})\geq 1-2n^{-2}$ under 1. Also note that, on $\mathcal{E}_{\pi}$ and under the condition $L\log(n)\leq n$,
$\max_{\ell\in[L]}\left|\log\left({\pi_{\ell}\over\widehat{\pi}_{\ell}}\right)\right|\leq\max_{\ell\in[L]}{|\widehat{\pi}_{\ell}-\pi_{\ell}|\over\pi_{\ell}}\lesssim\sqrt{L\log(n)\over n}$ (B.2)
as well as $\widehat{\pi}_{\ell}\asymp\pi_{\ell}$ for all $\ell\in[L]$.
B.1 Proof of the equivalence between (1.2) and (2.14)
Proof.
Recall that $B^{*}=\Sigma_{w}^{-1}{M}$. We find that
$B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}$ $=\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}$
so that, for any $\ell\in[L]$,
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}(x-\mu_{\ell})-2\log(\pi_{\ell})$
$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$
$\qquad+\mu_{\ell}^{\top}\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}).$
Write the singular value decomposition of $\Sigma_{w}^{-1/2}{M}$ as $U\Lambda V^{\top}$ with $U\in\mathbb{R}^{p\times r}$ containing the first $r$ singular vectors and $r={\rm rank}(\Sigma_{w}^{-1/2}{M})$. It follows that the above equals to
$\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{M}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{M}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
which, in view of (1.2) and (2.1), completes the proof. ∎
B.2 Proof of 2
For any $x\in\mathbb{R}^{p}$, write
$\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad g(x)=\operatorname*{arg\,min}_{\ell\in[L]}G_{\ell}(x).$
For any $t_{1},\ldots,t_{L}\geq 0$, define the event
$\mathcal{E}_{t}=\bigcap_{\ell\in[L]}\left\{\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|\leq t_{\ell}\right\}.$ (B.3)
In the proof of 2, we use the notation $\{Y=k\}:=\{Y={\bm{e}}_{k}\}$.
Proof.
We condition on ${\bm{D}}$ throughout the proof. The law of $\mathbb{E}$ and $\mathbb{P}$ is with respect to the randomness of $(X,Y)$, which is independent of ${\bm{D}}$. By definition, the excess risk $\mathcal{R}(\widehat{g})$ equals to
$\sum_{k\in[L]}\pi_{k}\Bigl{\{}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k\}\mid Y=k\right]-\mathbb{E}\left[\mathbbm{1}\{g^{*}(X)\neq k\}\mid Y=k\right]\Bigr{\}}$
$=\sum_{k\in[L]}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k,g^{*}(X)=k\}\mid Y=k\right]-\sum_{k\in[L]}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=k,g^{*}(X)\neq k\}\mid Y=k\right]$
$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y=k\right]-\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=k,g^{*}(X)=\ell\}\mid Y=k\right]$
$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\Bigl{\{}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y=k\right]-\pi_{\ell}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y=\ell\right]\Bigr{\}}.$
Write $f_{k}(x)$ as the p.d.f. of $X=x\mid Y={\bm{e}}_{k}$ for each $k\in[L]$. We find that
[…]
where, for all $k,\ell\in[L]$,
$G^{(\ell|k)}(x):=\left(x-{\mu_{\ell}+\mu_{k}\over 2}\right)^{\top}\Sigma_{w}^{-1}(\mu_{\ell}-\mu_{k})+\log{\pi_{\ell}\over\pi_{k}}={1\over 2}\left(G_{k}(x)-G_{\ell}(x)\right).$
Fix any $t_{1},\ldots,t_{L}\geq 0$. Observe that the event $\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\cap\mathcal{E}_{t}$ implies
$0>2G^{(\ell|k)}(X)\overset{\mathcal{E}_{t}}{\geq}\widehat{G}_{k}(X)-\widehat{G}_{\ell}(X)-(t_{k}+t_{\ell})\geq-(t_{k}+t_{\ell}).$
By the basic inequality $1+z\leq\exp(z)$ for all $z\in\mathbb{R}$, we obtain that,
$\mathcal{R}(\widehat{g})$ $\leq\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)$
$+{1\over 2}\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}(t_{k}+t_{\ell})\mathbb{E}\left[\mathbbm{1}\left\{-t_{k}-t_{\ell}\leq 2G^{(\ell|k)}(X)\leq 0,\widehat{g}(X)=\ell\mid Y=k\right\}\right].$
For the first term, we have
$\sum_{k\neq\ell}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)$ $\leq(L-1)\sum_{k=1}^{L}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)=(L-1)\mathbb{P}(\mathcal{E}_{t}^{c}).$ (B.4)
Regarding the second term, it is bounded from above by
${1\over 2}\sum_{k=1}^{L}\pi_{k}\sum_{\ell\in[L]\setminus\{k\}}(t_{k}+t_{\ell})\mathbb{P}\left\{-t_{k}-t_{\ell}\leq 2G^{(\ell|k)}(X)\leq 0\mid Y=k\right\}.$ (B.5)
Note that $G^{(\ell|k)}(X)\mid Y=k$ is normally distributed
$N\left(-{1\over 2}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})+\log\left(\pi_{\ell}/\pi_{k}\right),~{}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\right).$
By the mean-value theorem, the quantity in (B.5) is no greater than
${1\over 4}\sum_{k=1}^{L}\pi_{k}\sum_{\ell\in[L]\setminus\{k\}}(t_{k}+t_{\ell})^{2}$ $\leq{1\over 2}\sum_{k=1}^{L}\pi_{k}\left[(L-1)t_{k}^{2}+\sum_{\ell\in[L]\setminus\{k\}}t_{\ell}^{2}\right]={1\over 2}\sum_{k=1}^{L}(L\pi_{k}+1)t_{k}^{2}.$
Together with (B.4) and $\pi_{k}\leq C/L$, the proof is complete. ∎
B.3 Proof of Theorem 3
Proof.
Pick any $\ell\in[L]$. Recall that
$|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\leq|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}|+|(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+2|\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right)|.$ (B.6)
Throughout the proof, we work on $\mathcal{E}_{\pi}$ in (B.1) which implies (B.2), an upper bound for the last term on the right of (B.6). Recall that $\mathbb{P}\{\mathcal{E}_{\pi}\}\geq 1-2n^{-2}$ and note that we consider both the randomness of $(X,Y)$ and that of ${\bm{D}}$ in this proof.
For the first term in (B.6), we invoke Lemma 22 with $k=\ell$ and $v=B^{*}_{\ell}$, use the fact that $B_{\ell}^{*\top}\Sigma_{w}B_{\ell}^{*}=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}=\Delta_{\ell\ell}$ and take a union bound over $\ell\in[L]$ to conclude that the event
$\mathcal{E}_{\mu}:=\left\{\left|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}\right|\leq\sqrt{3\Delta_{\ell\ell}}\sqrt{L\log(n)\over n},~{}\text{for all }\ell\in[L]\right\}$ (B.7)
has probability at least $1-2Ln^{-3}\geq 1-n^{-2}$. Regarding the second term in (B.6)
$|\widehat{\mu}_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\leq|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|,$
the Cauchy-Schwarz inequality gives
$|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\leq\|\Sigma_{w}^{-1/2}\mu_{\ell}\|_{2}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}=\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\ell\ell}}$ (B.8)
Finally, conditioning on $Y={\bm{e}}_{k}$, the Gaussian tail of $X\mid Y,{\bm{D}}$ gives that, for all $t\geq 0$,
$\mathbb{P}\left\{|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\geq|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+t\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}~{}\Big{|}~{}Y={\bm{e}}_{k},{\bm{D}}\right\}\leq 2e^{-t^{2}/2}.$
Choosing $t=2\sqrt{\log(n)}$, taking the union bound over $k\in[L]$ and unconditioning yield that for all $\ell\in[L]$,
$|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\leq\sum_{k=1}^{L}\pi_{k}|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+2\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\log(n)}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+4\log(n)}$ (B.9)
holds with probability at least $1-2Ln^{-2}$. The last step uses (B.8). Collecting (B.2), (B.7), (B.8) and (B.3) completes the proof. ∎
B.4 Proof of 4
Proof.
For any $\ell\in[L]$, let
${t_{\ell}\over C}=\sqrt{\Delta_{\infty}}\sqrt{L\log(n)\over n}+\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+\log(n)}+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|.$
Recall from the proof of Theorem 3 that $1-\mathbb{P}(\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi})\leq 2n^{-2}+n^{-2}=3n^{-2}$ and
$\sum_{\ell=1}^{L}\mathbb{P}\left\{\left\{\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|\geq t_{\ell}\right\}\cap\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\right\}\leq{1\over n}.$
By the fact $\mathcal{R}(\widehat{g})\leq 1$ almost surely and invoking 2, we find that
$\mathbb{E}_{{\bm{D}}}[\mathcal{R}(\widehat{g})]$ $\leq\mathbb{E}_{{\bm{D}}}\left[\min\left\{C\sum_{\ell=1}^{L}t_{\ell}^{2}+L\sum_{\ell=1}^{L}\mathbb{P}\left\{|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\geq t_{\ell}\mid{\bm{D}}\right\},~{}1\right\}1\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}\right]$
$\quad+1-\mathbb{P}_{{\bm{D}}}\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}$
$\lesssim\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta_{\infty}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*})+{L\over n},~{}1\right\}1\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}\right]+{1\over n}$
$\lesssim\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta_{\infty}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*}),~{}1\right\}\right],$
completing the proof. ∎
B.5 Proof of Theorem 5
The proof of Theorem 5 uses the following lemma on $\|\widehat{H}-H\|_{\rm op}$. The proof of Lemma 10 appears at the end of this section.
Lemma 10.
Under the conditions of Theorem 5, with probability at least $1-n^{-1}$,
$\|\widehat{H}-H\|_{\rm op}\lesssim\sqrt{\log(n)\over nL}+{\omega_{2}\over\sqrt{L}}.$
Proof of Theorem 5.
Define
$\delta=C\sqrt{\log(n)\over nL}+C{\omega_{2}\over\sqrt{L}}$ (B.10)
for some constant $C>0$. We work on the event
$\mathcal{E}_{H}:=\left\{\|\widehat{H}-H\|_{\rm op}\leq\delta\right\}.$
Lemma 10 ensures that $\mathcal{E}_{H}$ holds with probability at least $1-n^{-1}$. Moreover, on the event $\mathcal{E}_{H}$, we find that
$\lambda_{K}(\widehat{H})$ $\geq\lambda_{K}(H)-\|\widehat{H}-H\|_{\rm op}$ by Weyl’s inequality
$=\left[\lambda_{1}(\Omega)\right]^{-1}-\|\widehat{H}-H\|_{\rm op}$ by $\text{by }\Omega=H^{-1}$
[…]
Hence, on the event $\mathcal{E}_{H}$, $\widehat{H}$ is invertible. We proceed to prove
$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}^{2}~{}\lesssim~{}\left(\omega_{2}^{2}+\delta^{2}L\Delta_{\infty}\right)L^{2}\Delta_{\infty}^{2},$ (B.11)
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}\lesssim\left(\omega_{1}^{2}{\log(n\vee p)\over n}+\delta^{2}L\Delta_{\infty}{\log(n)\over n}\right)L^{3}\Delta_{\infty}^{2}$ (B.12)
which, in conjunction with (B.10), $\Delta\asymp 1$, $L\log(n)\leq cn$ and $\omega_{2}\sqrt{L}\leq c$, yield the claim.
To prove (B.11), we write $\widehat{\Omega}=\widehat{H}^{-1}$ for simplicity. By definition,
$\widehat{B}^{*}-B^{*}$ $=(\widehat{B}-B)\widehat{\Omega}+B(\widehat{\Omega}-\Omega)$
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}H(\widehat{\Omega}-\Omega)$ by Lemma 1
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H\widehat{\Omega}-\mathbf{I}_{L})$ by $\text{by }H\Omega=\mathbf{I}_{L}$
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H-\widehat{H})\widehat{\Omega}$ by $\text{by }\widehat{H}\widehat{\Omega}=\mathbf{I}_{L}.$ (B.13)
It then follows that
$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\widehat{\Omega}\|_{F}+\|\Sigma_{w}^{1/2}B^{*}(H-\widehat{H})\widehat{\Omega}\|_{F}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}\|\widehat{\Omega}\|_{\rm op}+\|\Sigma_{w}^{1/2}B^{*}\|_{F}~{}\|(H-\widehat{H})\widehat{\Omega}\|_{\rm op}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}\|\widehat{\Omega}\|_{\rm op}+\sqrt{L\Delta_{\infty}}\|\widehat{H}-H\|_{{\rm op}}\|\widehat{\Omega}\|_{\rm op}.$
In the last step, we use $\|\Sigma_{w}^{1/2}B^{*}\|_{F}^{2}={\rm tr}(B^{*\top}\Sigma_{w}B^{*})={\rm tr}({M}^{\top}\Sigma_{w}^{-1}{M})\leq L\Delta_{\infty}$. We bound $\|\widehat{\Omega}\|_{\rm op}$ as follows:
$\|\widehat{\Omega}\|_{\rm op}$ $\leq\|\Omega\|_{\rm op}+\|\widehat{\Omega}-\Omega\|_{\rm op}$
$=\|\Omega\|_{\rm op}+\|\Omega(H-\widehat{H})\widehat{\Omega}\|_{\rm op}$
$\leq\|\Omega\|_{\rm op}+\|\Omega\|_{\rm op}\|H-\widehat{H}\|_{\rm op}\|\widehat{\Omega}\|_{\rm op}.$
Since the event $\mathcal{E}_{H}$ and $\omega_{2}\sqrt{L}\leq c$ ensure that
$\|\Omega\|_{\rm op}\|H-\widehat{H}\|_{\rm op}\leq\left(\max_{k\in[K]}{1\over\pi_{k}}+\Delta_{\rm op}\right)\delta\lesssim\delta L\Delta_{\infty}\leq{1\over 2},$
we conclude
$\|\widehat{\Omega}\|_{\rm op}\leq 2\|\Omega\|_{\rm op}\lesssim L\Delta_{\infty}.$ (B.14)
It now follows from 2 that
$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}$ $\lesssim\left(\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}+\delta\sqrt{L\Delta_{\infty}}\right)L\Delta_{\infty}$
$\leq\left(\omega_{2}+\delta\sqrt{L\Delta_{\infty}}\right)L\Delta_{\infty},$ (B.15)
proving (B.11).
We now prove (B.12). Using the identity (B.5), we have, on the event $\mathcal{E}_{H}$,
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}$
$\leq 2\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\right]^{2}+2\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}.$ (B.16)
For the first term on the right, we find
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\right]^{2}$ $=\sum_{\ell=1}^{L}(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\widehat{\Omega}_{\ell}^{\top}(\widehat{B}-B)^{\top}(\widehat{\mu}_{\ell}-\mu_{\ell})$
$\leq\max_{\ell\in[L]}\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}^{2}\sum_{\ell=1}^{L}\|(\widehat{B}-B)\widehat{\Omega}_{\ell}\widehat{\Omega}_{\ell}^{\top}(\widehat{B}-B)^{\top}\|_{1}.$
On the one hand, on the event $\mathcal{E}_{\pi}$, Lemma 22 with $k=\ell$ and $v={\bm{e}}_{j}$, in conjunction with union bounds over $j\in[p]$ and $\ell\in[L]$, yields that
$\max_{\ell\in[L]}\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}^{2}\lesssim{\|\Sigma_{w}\|_{\infty}}{L\log(n\vee p)\over n}$ (B.17)
with probability at least $1-n^{-2}$. On the other hand, we notice that
$\sum_{\ell=1}^{L}\|(\widehat{B}-B)\widehat{\Omega}_{\ell}\widehat{\Omega}_{\ell}^{\top}(\widehat{B}-B)^{\top}\|_{1}$ $=\sum_{\ell=1}^{L}\sum_{i,j\in[p]}\left|{\bm{e}}_{i}^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\widehat{\Omega}_{\ell}^{\top}(\widehat{B}-B)^{\top}{\bm{e}}_{j}\right|$
$\leq\sum_{i,j\in[p]}\left(\sum_{\ell=1}^{L}({\bm{e}}_{i}^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell})^{2}\right)^{1/2}\left(\sum_{\ell=1}^{L}({\bm{e}}_{j}^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell})^{2}\right)^{1/2}$
$=\sum_{i,j\in[p]}\|\widehat{\Omega}(\widehat{B}-B)^{\top}{\bm{e}}_{i}\|_{2}\|\widehat{\Omega}(\widehat{B}-B)^{\top}{\bm{e}}_{j}\|_{2}$
$\leq\|\widehat{\Omega}\|_{\rm op}^{2}\sum_{i,j\in[p]}\|(\widehat{B}-B)^{\top}{\bm{e}}_{i}\|_{2}\|(\widehat{B}-B)^{\top}{\bm{e}}_{j}\|_{2}$
$\leq\|\widehat{\Omega}\|_{\rm op}^{2}\|\widehat{B}-B\|_{1,2}^{2}.$
By invoking 2 and applying inequalities (B.14) and (B.17), we obtain
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}\lesssim\omega_{1}^{2}{\|\Sigma_{w}\|_{\infty}}{\log(n\vee p)\over n}L^{3}\Delta_{\infty}^{2}$ (B.18)
with probability at least $1-n^{-2}$. Regarding the second term on the right in (B.5), we have
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}$ $\leq\sum_{\ell=1}^{L}\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{2}^{2}\|(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{2}^{2}$
$\leq\max_{\ell\in[L]}\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{2}^{2}\|H-\widehat{H}\|_{\rm op}^{2}\sum_{\ell=1}^{L}\|\widehat{\Omega}_{\ell}\|_{2}^{2}$
$\leq\delta^{2}\max_{\ell\in[L]}\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{2}^{2}\|\widehat{\Omega}\|_{F}^{2}$
$\lesssim\delta^{2}\max_{\ell\in[L]}\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{2}^{2}~{}L\|\widehat{\Omega}\|_{\rm op}^{2}$
on the event $\mathcal{E}_{H}$. Further invoking (B.7) and (B.14) yields that
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}$ $~{}\lesssim\delta^{2}{\log(n)\over n}L^{4}\Delta_{\infty}^{3}$ (B.19)
with probability at least $1-n^{-2}$. Combining (B.18) with (B.19) proves (B.12), and hence completes the proof. ∎
Proof of Lemma 10.
Recall that
$\begin{split}\widehat{H}-H&=D_{\widehat{\pi}}-D_{\pi}-\left(\widehat{B}^{\top}\widehat{\Sigma}\widehat{B}-B^{\top}\Sigma B\right)\\ &=D_{\widehat{\pi}}-D_{\pi}-B^{\top}(\widehat{\Sigma}-\Sigma)B-(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\\ &\qquad-(\widehat{B}-B)^{\top}\widehat{\Sigma}B-B^{\top}\widehat{\Sigma}(\widehat{B}-B).\end{split}$ (B.20)
By triangle inequality, we have
$\|\widehat{H}-H\|_{\rm op}$ $\leq\|D_{\widehat{\pi}}-D_{\pi}\|_{\rm op}+\|(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\rm op}+2\|(\widehat{B}-B)^{\top}\widehat{\Sigma}B\|_{\rm op}$
$\quad+\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}.$
We proceed to bound each term on the right-hand-side separately. On the event $\mathcal{E}_{\pi}$, we first find that
$\|D_{\widehat{\pi}}-D_{\pi}\|_{\rm op}\leq\max_{\ell\in[L]}|\widehat{\pi}_{\ell}-\pi_{\ell}|\lesssim\sqrt{\log(n)\over nL}.$ (B.21)
Regarding the second term, 2 ensures that
$\|(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\rm op}\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{F}^{2}\leq\omega_{2}^{2}.$ (B.22)
For the last two terms, invoking Lemma 19 ensures with probability at least $1-n^{-2}$,
$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\lesssim{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}.$ (B.23)
It also follows that
$\|(\widehat{B}-B)^{\top}\widehat{\Sigma}B\|_{\rm op}$ $\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{\rm op}\|\widehat{\Sigma}^{1/2}B\|_{\rm op}$
$\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{\rm op}\left(\|B^{\top}\Sigma B\|_{\rm op}+\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\right)^{1/2}$
$\lesssim\omega_{2}\sqrt{\Delta_{\rm op}\over L(L+\Delta_{\rm op})}$
where the last step uses
[…]
deduced from Lemma 14. Combining the above bounds with (B.21), (B.22) and (B.23) gives
$\|\widehat{H}-H\|_{\rm op}$ $\lesssim\sqrt{\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}$
$\qquad+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}+\omega_{2}\sqrt{\Delta_{\rm op}\over L(L+\Delta_{\rm op})}+\omega_{2}^{2}$
with probability at least $1-n^{-1}$. The result follows by using $\Delta_{\infty}\asymp 1$ and $\omega_{2}\sqrt{L}\leq c$ to collect terms. ∎
B.6 Proof of Theorem 6: the lasso estimator of $B$
The proof of Theorem 6 uses the following Restricted Eigenvalue condition (RE) on the within-class covariance matrix $\Sigma_{w}$.
Definition 1.
For any integer $1\leq s\leq p$, define
$\kappa_{s}:=\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{v^{\top}\Sigma_{w}v\over v^{\top}v}$
where $\mathcal{C}(S,3):=\{u\in\mathbb{R}^{p}\setminus\{0\}:\|u_{S^{c}}\|_{1}\leq 3\|u_{S}\|_{1}\}$.
We prove Theorem 6 under the following weaker condition than 3.
Assumption 5.
For some $1\leq s\leq p$, there exists some absolute constants $0<c\leq C<\infty$ such that $c<\kappa_{s}\leq\|\Sigma_{w}\|_{\infty}\leq C$.
Proof of Theorem 6.
We first prove that
$\begin{split}&\max_{\ell\in[L]}{1\over\sqrt{n}}\|{\mathbf{X}}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}\\ &\max_{\ell\in[L]}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}\\ &\max_{\ell\in[L]}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}~{}\lesssim~{}\kappa_{s}^{-1}\lambda\sqrt{s}\end{split}$ (B.24)
hold with probability at least $1-2n^{-1}$. We work on the intersection of the event
$\mathcal{E}_{\lambda}:=\left\{{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}\leq{1\over 2}\lambda\right\}$ (B.25)
with the event
$\mathcal{E}_{RE}:=\left\{\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{\|{\mathbf{X}}v\|_{2}^{2}\over n\|v\|_{2}^{2}}\geq{\kappa_{s}\over 2}\right\}.$ (B.26)
Recall $\mathcal{C}(S,3):=\{u\in\mathbb{R}^{p}\setminus\{0\}:\|u_{S^{c}}\|_{1}\leq 3\|u_{S}\|_{1}\}$. The main difficulty in proving Theorem 6 is to establish the order of $\lambda$ as well as to show that $\mathcal{E}_{\lambda}\cap\mathcal{E}_{RE}$ holds with overwhelming probability.
Lemma 11 ensures that $\mathcal{E}_{\lambda}$ holds with probability at least $1-n^{-1}$ for any $\lambda$ satisfying (3.5).
We proceed to verify that $\mathbb{P}(\mathcal{E}_{RE})\geq 1-(n\vee p)^{-2}$. To this end, for any set $S\subseteq[p]$, we write $E_{S}:=\text{span}\{{\bm{e}}_{j}:j\in S\}$. Under 5, we define the space
$\Gamma_{s}:=\bigcup_{S\subseteq[p]:\ |S|=\lfloor Cs\rfloor}E_{S}$ (B.27)
for some constant $C=C(\kappa_{s},\|\Sigma_{w}\|_{\infty})>1$. Recall that $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$. Lemma 12 and the condition $(s\vee L)\log(n\vee p)\leq c~{}n$ imply that
${9\over 10}\|\Sigma^{1/2}v\|_{2}\leq\|\widehat{\Sigma}^{1/2}v\|_{2}\leq{11\over 10}\|\Sigma^{1/2}v\|_{2}\qquad\forall~{}v\in\Gamma_{s}$ (B.28)
holds with probability at least $1-(n\vee p)^{-2}$. According to Rudelson and Zhou (2012, Theorem 3), (B.28) implies that there exists $S\subseteq[p]$ with $|S|\leq s$ such that
$u^{\top}\widehat{\Sigma}u\geq{1\over 2}u^{\top}\Sigma u$ (B.29)
holds for any $u\in\mathcal{C}(S,3)$. Eq. B.29 in conjunction with 5 ensures that $\mathcal{E}_{RE}$ holds.
It is now relatively straightforward to finish the proof of (B.24). Standard arguments (Bühlmann and Van de Geer, 2011; Giraud, 2021) yield that on the event $\mathcal{E}_{\lambda}$, for any $\ell\in[L]$,
${1\over 2n}\|{\mathbf{X}}\widehat{B}_{\ell}-{\mathbf{X}}B_{\ell}\|_{2}^{2}$ $\leq{\lambda\over 2}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}+\lambda\|B_{\ell}\|_{1}-\lambda\|\widehat{B}_{\ell}\|_{1}$
from which one can deduce that
$\widehat{B}_{\ell}-B_{\ell}\in\mathcal{C}(S_{\ell},3)$ (B.30)
with $S_{\ell}:={\rm supp}(B_{\ell})$. Invoking (B.29), we find on the event $\mathcal{E}_{RE}$
${\kappa_{s}\over 4}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}^{2}\leq{1\over 4}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}^{2}$ $\leq{1\over 2n}\|{\mathbf{X}}\widehat{B}_{\ell}-{\mathbf{X}}B_{\ell}\|_{2}^{2}$
$\leq{3\lambda\over 2}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}$
$\leq 6\lambda\|(\widehat{B}_{\ell}-B_{\ell})_{S_{\ell}}\|_{1}$ by (B.30)
$\leq 6\lambda\sqrt{s}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}.$
We have proved (B.24).
Finally, the claim follows by noting that
$\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{F}\leq\sqrt{L}\max_{\ell\in[L]}\|\widehat{\Sigma}^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2},$
$\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}\leq\sqrt{L}\max_{\ell\in[L]}\|\Sigma_{w}^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}\leq\sqrt{L}\max_{\ell\in[L]}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2},$
$\|\widehat{B}-B\|_{1,2}=\sum_{j=1}^{p}\|\widehat{B}_{j\cdot}-B_{j\cdot}\|_{2}\leq\sum_{j=1}^{p}\|\widehat{B}_{j\cdot}-B_{j\cdot}\|_{1}=\sum_{\ell=1}^{L}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}.$
The proof is complete. ∎
B.6.1 Two key lemmas used in the proof of Theorem 6
Lemma 11.
Under model (1.1) with 1, assume $L\log(n\vee p)\leq n$. With probability at least $1-n^{-1}$, one has
$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbf{X}}_{j}^{\top}({\mathbf{Y}}_{\ell}-{\mathbf{X}}B_{\ell})\right|\lesssim\left(\sqrt{\|\Sigma_{w}\|_{\infty}}+\|{M}\|_{\infty}\right)\sqrt{\log(n\vee p)\over nL}.$
Proof.
Note that
${1\over n}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)$ $={1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-\Sigma_{XY}-\widehat{\Sigma}B+\Sigma B$ (B.31)
We first analyze the term $(\widehat{\Sigma}-\Sigma)B$. Pick any $j\in[p]$, $\ell\in[L]$ and fix $t\geq 0$. We observe from Lemma 21 and the inequality $L\leq n$ that $\mathcal{E}_{\pi}$ holds with probability at least $1-2n^{-1}$. Next, we invoke Lemma 18 with $v_{1}={\bm{e}}_{j}$, $v_{2}=B_{\ell}$, $t=C\log(n\vee p)$ and we use (C.5), (C.6) and (C.7) in Lemma 14 to derive that the inequalities
$\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{[\Sigma_{w}]_{jj}B_{\ell}^{\top}\Sigma_{w}B_{\ell}}\sqrt{\log(n\vee p)\over n}$
$\quad+\|{M}_{j\cdot}\|_{\infty}\sqrt{B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}}\sqrt{\log(n\vee p)\over nL}+\|{M}_{j\cdot}\|_{\infty}\|{M}^{\top}B_{\ell}\|_{\infty}{\log(n\vee p)\over n}$
$\quad+\sqrt{[\Sigma_{w}]_{jj}B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}}\sqrt{\log(n\vee p)\over nL}+\|{M}_{j\cdot}\|_{2}\sqrt{B_{\ell}^{\top}\Sigma_{w}B_{\ell}}\sqrt{\log(n\vee p)\over nL}$
$\lesssim\sqrt{[\Sigma_{w}]_{jj}+\|{M}_{j\cdot}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}+\|{M}_{j\cdot}\|_{\infty}{\log(n\vee p)\over n}$
$\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$
hold with probability at least $1-(n\vee p)^{-3}$. After we take the union bound over $j\in[p]$ and $\ell\in[L]$, we have that
$\max_{j\in[p],\ \ell\in[L]}\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (B.32)
holds with probability $1-n^{-1}$.
To bound the first term in (B.31), by recalling that $\Sigma_{XY}={M}D_{\pi}$, we find
${1\over n}{\mathbf{X}}_{j}^{\top}{\mathbf{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}$ $={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{L}\left(X_{ij}Y_{i\ell}-{M}_{j\ell}\pi_{\ell}\right)\mathbbm{1}\{Y_{i}={\bm{e}}_{\ell}\}$
$={1\over n}\sum_{i=1}^{L}\left(X_{ij}-{M}_{j\ell}\right)\mathbbm{1}\{Y_{i}=\ell\}+\left({n_{\ell}\over n}-\pi_{\ell}\right){M}_{j\ell}$
$={n_{\ell}\over n}(\widehat{M}_{j\ell}-{M}_{j\ell})+(\widehat{\pi}_{\ell}-\pi_{\ell}){M}_{j\ell}.$
Invoking 1, Lemma 21 and Lemma 22 with $t=\sqrt{\log(p\vee n)}$ and taking a union bound over $j\in[p]$ and $\ell\in[L]$, we obtain
$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbf{X}}_{j}^{\top}{\mathbf{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}\right|\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}\log(p\vee n)\over nL}+\|{M}\|_{\infty}\sqrt{\log(n)\over nL}$ (B.33)
with probability $1-n^{-1}$. Combining (B.32) and (B.33) completes the proof. ∎
Recall the set
$\Gamma_{s}=\bigcup_{S\subseteq[p]:|S|=\lfloor Cs\rfloor}E_{S}$
with $E_{S}=\text{span}\{{\bm{e}}_{j}:j\in S\}$. The following lemma gives an upper bound for $v^{\top}(\widehat{\Sigma}-\Sigma)v$, uniformly over $v\in\Gamma_{s}$.
Lemma 12.
Under model (1.1) with Assumptions 1 & 5, assume $(s\vee L)\log(n\vee p)\leq n$ for some integer $1\leq s<p$. Then, on the event $\mathcal{E}_{\pi}$, with probability at least $1-(n\vee p)^{-2}$, the following holds uniformly over $v\in\Gamma_{s}$,
$v^{\top}(\widehat{\Sigma}-\Sigma)v~{}\lesssim~{}v^{\top}\Sigma v\left(\sqrt{s\log(n\vee p)\over n}+\sqrt{L\log(n)\over n}\right).$
Proof.
Without loss of generality, we prove the result for
$v\in\Gamma_{s}^{\prime}:=\Gamma_{s}\cap\mathcal{S}^{p}$
via a discretization argument. For any subset $S\subseteq[p]$ with $|S|=\lfloor Cs\rfloor$, let $\mathcal{N}_{S}$ be the $(1/3)$-net of $E_{S}\cap\mathcal{S}^{p}$ satisfying
$\mathcal{N}_{S}\subset E_{S}\cap\mathcal{S}^{p},\qquad|\mathcal{N}_{S}|\leq 7^{|S|}$
The existence of such net is ensured by Rudelson and Zhou (2012, Lemma 23). Hence, for any fixed $v\in E_{S}\cap\mathcal{S}^{p}$, there exists $u\in\mathcal{N}_{S}$ such that $\|u-v\|_{2}\leq 1/3$, and we find
$v^{\top}(\widehat{\Sigma}-\Sigma)v$ $\leq(v-u)^{\top}(\widehat{\Sigma}-\Sigma)v+u^{\top}(\widehat{\Sigma}-\Sigma)(v-u)+u^{\top}(\widehat{\Sigma}-\Sigma)u$
$\leq 2\|u-v\|_{2}\sup_{v\in E_{S}\cap\mathcal{S}^{p}}v^{\top}(\widehat{\Sigma}-\Sigma)v+u^{\top}(\widehat{\Sigma}-\Sigma)u.$
The second inequality uses the fact that $(u-v)/\|u-v\|_{2}\in E_{S}\cap\mathcal{S}^{p}$ whenever $u\neq v$. Taking the supremum over $v\in E_{S}\cap\mathcal{S}^{p}$ and the maximum over $u\in\mathcal{N}_{S}$ gives
$\sup_{v\in E_{S}\cap\mathcal{S}^{p}}v^{\top}(\widehat{\Sigma}-\Sigma)v\leq 3\max_{u\in\mathcal{N}_{S}}u^{\top}(\widehat{\Sigma}-\Sigma)u$
and
$\sup_{v\in\Gamma_{s}^{\prime}}v^{\top}(\widehat{\Sigma}-\Sigma)v\leq 3\max_{u\in\mathcal{N}}u^{\top}(\widehat{\Sigma}-\Sigma)u.$ (B.34)
Here
$\mathcal{N}:=\bigcup_{S\subseteq[p]:\ |S|=\lfloor Cs\rfloor}\mathcal{N}_{S}$
has cardinality
$|\mathcal{N}|\leq\sum_{|S|=\lfloor Cs\rfloor}|\mathcal{N}_{S}|\lesssim{p}^{Cs}.$
We will bound the right hand side (B.34) via the same arguments used to prove Lemma 18, except for the control of the term $\sum_{k}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v$. Specifically, repeating the proof of Lemma 18 with $v_{1}=v_{2}=v$, for any $v\in\mathcal{N}$, gives that for any $t\geq 0$
$v^{\top}(\widehat{\Sigma}-\Sigma)v$ $\lesssim v^{\top}\Sigma_{w}v\left(\sqrt{t\over n}+{t\over n}\right)+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v+2{\|{M}^{\top}v\|_{2}\over\sqrt{L}}\sqrt{v^{\top}\Sigma_{w}v}\sqrt{t\over n}.$
holds, with probability at least $1-6e^{-t}$. Note that
${\|{M}^{\top}v\|_{2}^{2}\over L}\lesssim v^{\top}{M}D_{\pi}{M}^{\top}v$
under 1. By using the decomposition of $\Sigma$ in (2.6), choosing $t=Cs\log(n\vee p)$ for some large constant $C>0$, we obtain that
$v^{\top}(\widehat{\Sigma}-\Sigma)v\lesssim v^{\top}\Sigma v\left(\sqrt{s\log(n\vee p)\over n}+{s\log(n\vee p)\over n}\right)+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v$
holds uniformly over $v\in\mathcal{N}$, with probability at least $1-(n\vee p)^{-C^{\prime\prime}s}$. On the event $\mathcal{E}_{\pi}$, invoking 1, we conclude that
$\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v\leq\max_{k\in[L]}{|\widehat{\pi}_{k}-\pi_{k}|\over\pi_{k}}v^{\top}{M}D_{\pi}{M}^{\top}v\lesssim v^{\top}{M}D_{\pi}{M}^{\top}v\sqrt{L\log(n)\over n}$
holds uniformly over $\mathcal{N}$. The proof is complete in view of (B.34). ∎
B.7 Proof of Theorem 8: the reduced-rank estimator of $B$
Proof.
Let $C>0$ be some constant to be specified later. We work on the event
$\mathcal{E}_{\lambda}^{\prime}=\left\{{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}\leq\sqrt{\lambda/C}\right\}$
intersecting with
$\mathcal{E}_{X}=\left\{\lambda_{p}(\widehat{\Sigma})\geq{1\over 2}\lambda_{p}(\Sigma)\geq{c\over 2}\right\}.$
Here $c$ is defined in 3. Lemma 20, Lemma 13 and 3 together with (3.6) ensure that $\mathcal{E}_{X}\cap\mathcal{E}_{\lambda}^{\prime}$ hold with probability at least $1-n^{-1}$. Write $\widehat{r}={\rm rank}(\widehat{B})$ and $r={\rm rank}(B)$. By the optimality of $\widehat{B}$, we have
${1\over n}\|{\mathbf{Y}}-{\mathbf{X}}\widehat{B}\|_{F}^{2}+\lambda\widehat{r}$ $~{}\leq~{}{1\over n}\|{\mathbf{Y}}-{\mathbf{X}}B\|_{F}^{2}+\lambda r.$
Working out the squares gives
${1\over n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}$ $\leq{2\over n}|\langle{\mathbf{Y}}-{\mathbf{X}}B,{\mathbf{X}}(\widehat{B}-B)\rangle|+\lambda(r-\widehat{r})$
$\leq 2\|\widehat{B}-B\|_{*}{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}+\lambda(r-\widehat{r})$
$\leq 2\sqrt{r+\widehat{r}}~{}\|\widehat{B}-B\|_{F}\sqrt{\lambda/C}+\lambda(r-\widehat{r})$ by $\text{by $\mathcal{E}_{\lambda}^{\prime}$}.$
Using the properties on $\mathcal{E}_{X}$, we find
$\|\widehat{B}-B\|_{F}^{2}$ $\leq{2\over c}2\sqrt{\widehat{r}+r}\|\widehat{B}-B\|_{F}\sqrt{\lambda/C}+{2\over c}\lambda(r-\widehat{r})$
$\leq{8\lambda\over c^{2}C}(\widehat{r}+r)+{1\over 2}\|\widehat{B}-B\|_{F}^{2}+{2\over c}\lambda(r-\widehat{r})$ by $2xy\leq x^{2}/2+2y^{2}$
$={1\over 2}\|\widehat{B}-B\|_{F}^{2}+{4\over c}\lambda r$ for $\text{for $C=4/c$}.$
In the same way, on the event $\mathcal{E}_{X}$,
${1\over n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}$ $\leq\sqrt{{8\over cC}\lambda(r+\widehat{r})}~{}{1\over\sqrt{n}}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}+\lambda(r-\widehat{r})$
$\leq\frac{1}{2n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}+2\lambda r$ for $\text{for $C=4/c$}.$
This implies the rates for both $\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}$ and $\|\widehat{B}-B\|_{F}^{2}$. They imply in turn the rates for $\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}$ and $\|\widehat{B}-B\|_{2,1}$ invoking the inequalities
$\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}^{2}$ $\leq\|\Sigma^{1/2}(\widehat{B}-B)\|_{F}^{2}\leq{2\over c}\|\widehat{B}-B\|_{F}^{2}$ by (2.6)
and
$\|\widehat{B}-B\|_{1,2}^{2}$ $\leq{p}\|\widehat{B}-B\|_{F}^{2}.$
This completes the proof. ∎
Lemma 13.
Under model (1.1) and 1, assume $(p+L)\log(n\vee p)\leq n$. With probability at least $1-2n^{-2}$, we have
${1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}\lesssim\sqrt{\|\Sigma_{w}\|_{\rm op}(1+\Delta_{\infty})}\sqrt{(p+L)\log(n)\over nL}.$
Proof.
By (B.31), it suffices to bound from above
$\textrm{I}=\left\|(\widehat{\Sigma}-\Sigma)B\right\|_{\rm op},\qquad\textrm{II}=\left\|{1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-\Sigma_{XY}\right\|_{\rm op}=\left\|{1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-{M}D_{\pi}\right\|_{\rm op}.$
For I, by invoking Lemma 20, we find with probability at least $1-n^{-2}$,
I $\leq\|\Sigma^{1/2}\|_{\rm op}\|\Sigma^{-1/2}(\widehat{\Sigma}-\Sigma)\Sigma^{-1/2}\|_{\rm op}\|\Sigma^{1/2}B\|_{\rm op}$
$\lesssim\sqrt{\|\Sigma_{w}\|_{\rm op}(1+\Delta_{\infty})}\sqrt{(p+L)\log(n)\over nL}$
where in the last step we also use
$\|\Sigma\|_{\rm op}=\|\Sigma_{w}^{1/2}(\mathbf{I}_{p}+\Sigma_{w}^{-1/2}{M}D_{\pi}{M}^{\top}\Sigma_{w}^{-1/2})\Sigma_{w}^{1/2}\|_{\rm op}\leq\|\Sigma_{w}\|_{\rm op}\left(1+{\Delta_{\rm op}\over L}\right)\leq\|\Sigma_{w}\|_{\rm op}(1+\Delta_{\infty})$
as well as
$\|\Sigma^{1/2}B\|_{\rm op}^{2}=\|B^{\top}\Sigma B\|_{\rm op}=\|D_{\pi}{M}^{\top}B\|_{\rm op}\lesssim{\Delta_{\rm op}\over L(L+\Delta_{\rm op})}\leq{1\over L}$
from Lemma 14.
Regarding II, by standard discretization argument, it suffices to bound from above
$\max_{v\in\mathcal{N}_{p}(1/3),u\in\mathcal{N}_{L}(1/3)}v^{\top}\left({1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-{M}D_{\pi}\right)u$
with $\mathcal{N}_{p}(1/3)$ and $\mathcal{N}_{L}(1/3)$ being the $(1/3)$-epsilon net of $\mathcal{S}_{p}$ and $\mathcal{S}_{m}$, respectively. Fix any $v\in\mathcal{N}_{p}(1/3)$ and $u\in\mathcal{N}_{L}(1/3)$. Observe that
$v^{\top}\left({1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-{M}D_{\pi}\right)u$ $=v^{\top}(\widehat{M}D_{\widehat{\pi}}-{M}D_{\pi})u$
$=v^{\top}(\widehat{M}-{M})D_{\widehat{\pi}}u+v^{\top}{M}\left(D_{\widehat{\pi}}-D_{\pi}\right)u$
$=\sum_{k=1}^{L}u_{k}{n_{k}\over n}v^{\top}(\widehat{\mu}_{k}-\mu_{k})+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})u_{k}v^{\top}\mu_{k}.$
Invoking Lemma 15 and Lemma 16 yields that, for any $t\geq 0$, with probability $1-4e^{-t/2}$,
II $\lesssim\|u\|_{2}\sqrt{tv^{\top}\Sigma_{w}v\over n}\sqrt{\max_{k}n_{k}\over n}+\sqrt{t\sum_{k=1}^{m}\pi_{k}u_{k}^{2}(v^{\top}\mu_{k})^{2}\over n}+{t\|u\|_{\infty}\|v^{\top}{M}\|_{\infty}\over n}$
$\lesssim\sqrt{t\|\Sigma_{w}\|_{\rm op}\over nL}+\sqrt{t\|\Sigma_{w}\|_{\rm op}\Delta_{\infty}\over nL}+\sqrt{\|\Sigma_{w}\|_{\rm op}\Delta_{\infty}}{t\over n}$
where in the last step we use
$\|v^{\top}{M}\|_{\infty}=\|v^{\top}\Sigma_{w}^{1/2}\|_{2}\|\Sigma_{w}^{-1/2}{M}\|_{2,\infty}\leq\|\Sigma_{w}\|_{\rm op}^{1/2}\sqrt{\Delta_{\infty}}.$
Recall that $|\mathcal{N}_{p}(1/3)|\leq 7^{p}$ and $|\mathcal{N}_{L}(1/3)|\leq 7^{L}$. Choosing $t=C(p+L)\log(n)$, taking the union bounds over $\mathcal{N}_{p}(1/3)$ and $\mathcal{N}_{L}(1/3)$ and combining with the bound of I complete the proof. ∎
Appendix C Technical lemmas
The following lemma provides a few useful facts on quantities related with $B$ in (1.12). Write
$\Omega:=D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}.$ (C.1)
Lemma 14.
With $B$ defined in (1.12), we have
${M}^{\top}B={M}^{\top}\Sigma_{w}^{-1}{M}~{}\Omega^{-1},$ (C.2)
$H=D_{\pi}-D_{\pi}{M}^{\top}B=\Omega^{-1},$ (C.3)
$B^{\top}\Sigma_{w}B=\Omega^{-1}{M}^{\top}\Sigma_{w}^{-1}{M}~{}\Omega^{-1}.$ (C.4)
Consequently, for any $k,\ell\in[L]$, we have
$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}\leq\pi_{\ell}\left(1\wedge\pi_{\ell}\Delta_{\ell\ell}\right),$ (C.5)
$B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq\left(1\wedge{\pi_{\ell}^{2}\over\pi_{\min}}\Delta_{\ell\ell}\right),$ (C.6)
$\left|\mu_{k}^{\top}B_{\ell}\right|\leq\left(1\wedge\pi_{\ell}\sqrt{\Delta_{kk}\Delta_{\ell\ell}}\right).$ (C.7)
Proof.
From Lemma 1 and its proof, we know that
$B=B^{*}\left(D_{\pi}-D_{\pi}{M}^{\top}B\right)=\Sigma_{w}^{-1}{M}\left(D_{\pi}-D_{\pi}{M}^{\top}B\right)$
so that
${M}^{\top}B={M}^{\top}\Sigma_{w}^{-1}{M}\left(D_{\pi}-D_{\pi}{M}^{\top}B\right)={M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}-{M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}{M}^{\top}B.$
Rearranging terms gives
${M}^{\top}B$ $=\left(\mathbf{I}_{L}+{M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}\right)^{-1}{M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}$
$=\mathbf{I}_{L}-\left(\mathbf{I}_{L}+{M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}\right)^{-1}$
$=\mathbf{I}_{L}-D_{\pi}^{-1}\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1}$ (C.8)
$={M}^{\top}\Sigma_{w}^{-1}{M}\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1}$
proving (C.2). By using the above identity (C.8), we have
$D_{\pi}-D_{\pi}{M}^{\top}B=\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1},$
proving (C.3). Since Lemma 1 and (C.3) imply
$B=B^{*}\Omega^{-1},$
(C.4) follows from
$B^{\top}\Sigma_{w}B=\Omega^{-1}B^{*\top}\Sigma_{w}B^{*}\Omega^{-1}=\Omega^{-1}{M}^{\top}\Sigma_{w}^{-1}{M}\Omega^{-1}.$
To prove (C.5) – (C.7), pick any $k,\ell\in[L]$. By using (2.6) twice and (2.5), we find that
$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}=\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}\leq\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$ (C.9)
yielding the second bound in (C.5). The other bound in (C.5) follows by observing that
[…]
For (C.6), on the one hand, similar arguments yield
$B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}{M}D_{\pi}{M}^{\top}\Sigma^{-1}\mu_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}.$
On the other hand, using (C.2) and (C.3) proves
$B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq{\bm{e}}_{\ell}^{\top}\Omega^{-1}({M}^{\top}\Sigma_{w}^{-1}{M})^{2}\Omega^{-1}{\bm{e}}_{\ell}\leq 1.$
Finally, the last statement follows by noting that
$|\mu_{k}^{\top}B_{\ell}|^{2}\leq B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq 1$
and
$|\mu_{k}^{\top}B_{\ell}|=\pi_{\ell}|\mu_{k}^{\top}\Sigma^{-1}\mu_{\ell}|\leq\pi_{\ell}\sqrt{\mu_{k}^{\top}\Sigma^{-1}\mu_{k}}\sqrt{\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}}.$
The following lemmas contain deviation inequalities related with $\widehat{M}-{M}$ and $\widehat{\pi}-\pi$.
Lemma 15.
Let $v\in\mathbb{R}^{p}$ and $u\in\mathbb{R}^{L}$ be any fixed vectors. Under (1.1), for any $t\geq 0$, with probability at least $1-2e^{-t^{2}/2}$, we have
$\left|\sum_{k=1}^{L}{n_{k}\over n}v^{\top}(\widehat{\mu}_{k}-\mu_{k})u_{k}\right|\leq t\|u\|_{2}\sqrt{v^{\top}\Sigma_{w}v\over n}\sqrt{\max_{k}n_{k}\over n}.$
Proof.
Recall that, for any $k\in[L]$,
$\widehat{\mu}_{k}={1\over n_{k}}\sum_{i=1}^{n}X_{i}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}.$
We obtain
$\sum_{k=1}^{L}{n_{k}\over n}v^{\top}(\widehat{\mu}_{k}-\mu_{k})u_{k}$ $=v^{\top}\sum_{k=1}^{L}{1\over n}\sum_{i=1}^{n}\left(X_{i}-\mu_{k}\right)u_{k}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}$
$=v^{\top}\sum_{k=1}^{L}{1\over n}\sum_{i=1}^{n}\left(X_{i}-{M}Y_{i}\right)u^{\top}Y_{i}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}$
$={1\over n}\sum_{i=1}^{n}v^{\top}\left(X_{i}-{M}Y_{i}\right)u^{\top}Y_{i}.$
For $i\in[n]$, by conditioning on ${\mathbf{Y}}$, we know that $W_{i}:=X_{i}-{M}Y_{i}$ are i.i.d. from $\mathcal{N}_{p}(0,\Sigma_{w})$, hence
${1\over n}\sum_{i=1}^{n}v^{\top}W_{i}u^{\top}Y_{i}~{}\Big{|}~{}{\mathbf{Y}}\sim\mathcal{N}\left(0,{v^{\top}\Sigma_{w}v\over n}{u^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}u\over n}\right).$
The claim follows by invoking the standard Gaussian tail probability bounds and after unconditioning. ∎
Lemma 16.
Let $u,v\in\mathbb{R}^{L}$ be any fixed vectors. Under (1.1), for any $t\geq 0$, with probability at least $1-2e^{-t/2}$,
$\left|\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})u_{k}v_{k}\right|\lesssim\sqrt{t\sum_{k=1}^{L}\pi_{k}u_{k}^{2}v_{k}^{2}\over n}+{t\|u\|_{\infty}\|v\|_{\infty}\over n}.$
Proof.
Observe that we can write
$\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})u_{k}v_{k}$ $=\sum_{k=1}^{L}{1\over n}\sum_{i=1}^{n}Y_{ik}u_{k}v_{k}-\sum_{k=1}^{L}\pi_{k}u_{k}v_{k}$
$={1\over n}\sum_{i=1}^{n}Z_{i}-\pi_{k}u_{k}v_{k}$
Here
$Z_{i}=\sum_{k=1}^{L}Y_{ik}u_{k}v_{k}$
are independent, bounded random variables (as $|Z_{i}|\leq\|u\|_{\infty}\|v\|_{\infty}$) with mean
$\mathbb{E}[Z_{i}]=\sum_{k=1}^{L}\pi_{k}u_{k}v_{k}$
and variance
${\rm Var}(Z_{i})={\rm Var}(y_{i}^{\top}{\rm diag}(u)v)\leq\sum_{k=1}^{L}\pi_{k}v_{k}^{2}u_{k}^{2}$
for $i\in[n]$. The proof follows after a straightforward application of Bernstein’s inequality for bounded random variables. ∎
The following lemma controls the difference
$\widehat{\Sigma}_{w}-\Sigma_{w}:={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}-\Sigma_{w}.$
Lemma 17.
Let $u,v\in\mathbb{R}^{p}$ be any fixed vectors. Under model (1.1), for any $t\geq 0$, with probability at least $1-2e^{-t}$, one has
$\left|u^{\top}\left(\widehat{\Sigma}_{w}-\Sigma_{w}\right)v\right|\lesssim\sqrt{u^{\top}\Sigma_{w}u}\sqrt{v^{\top}\Sigma_{w}v}\left(\sqrt{t\over n}+{t\over n}\right)$
Proof.
Start with
$\widehat{\Sigma}_{w}-\Sigma_{w}$ $={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}-\Sigma_{w}$
$={1\over n}\sum_{i=1}^{n}\left[(X_{i}-{M}Y_{i})(X_{i}-{M}Y_{i})^{\top}-\Sigma_{w}\right].$
Recall that, conditioning ${\mathbf{Y}}$, $W_{i}=X_{i}-{M}Y_{i}$ for $i\in[n]$ are independent, $\mathcal{N}_{p}(0,\Sigma_{w})$. The result follows from standard concentration inequalities for the quadratic term of Gaussian random vectors. ∎
C.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
Recall that $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$ and the definition of the event $\mathcal{E}_{\pi}$ is given in (B.1).
Lemma 18.
Under model (1.1) and 1,assume $L\log(n)\leq n$. On the event $\mathcal{E}_{\pi}$, for any fixed $v_{1},v_{2}\in\mathbb{R}^{p}$ and any $t\geq 0$, with probability at least $1-8e^{-t}$, one has
$\left|v_{1}^{\top}(\widehat{\Sigma}-\Sigma)v_{2}\right|$ $\lesssim~{}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}\left(\sqrt{t\over n}+{t\over n}\right)$
$\quad+\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}{t\over n}$
$\quad+\left({\|{M}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}+{\|{M}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\right)\sqrt{t\over n}.$
Proof.
By definition, we have
$\widehat{\Sigma}$ $={1\over n}{\mathbf{X}}^{\top}{\mathbf{X}}={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}X_{i}X_{i}^{\top}$
$={1\over n}\sum_{k=1}^{L}\left[\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}+n_{k}(\widehat{\mu}_{k}\mu_{k}^{\top}+\mu_{k}\widehat{\mu}_{k}^{\top})-n_{k}\mu_{k}\mu_{k}^{\top}\right]$
$=\widehat{\Sigma}_{w}+\sum_{k=1}^{L}{n_{k}\over n}\mu_{k}\mu_{k}^{\top}+\sum_{k=1}^{L}{n_{k}\over n}\left[(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}+\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}\right]$
where we write
$\widehat{\Sigma}_{w}:={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}.$
By further using the decomposition of $\Sigma$ in (2.6) and $\widehat{\pi}_{k}=n_{k}/n$ for $k\in[L]$, we find that
$\widehat{\Sigma}-\Sigma$ $=\widehat{\Sigma}_{w}-\Sigma_{w}+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})\mu_{k}\mu_{k}^{\top}+\sum_{k=1}^{L}\widehat{\pi}_{k}\left[(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}+\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}\right].$
The first term $\widehat{\Sigma}_{w}-\Sigma_{w}$ is bounded in Lemma 17. We bound the second term by invoking 1 and Lemma 16 with $u={M}^{\top}v_{1}$ and $v={M}^{\top}v_{2}$. This yields
$\left|\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v_{1}^{\top}\mu_{k}\mu_{k}^{\top}v_{2}\right|$ $\lesssim\sqrt{t\sum_{k=1}^{L}\pi_{k}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over n}+{t\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}\over n}$
$\lesssim\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}{t\over n}$ (C.11)
with probability at least $1-2e^{-t}$. Furthermore, on the event $\mathcal{E}_{\pi}$, 1 and $L\log(n)\leq n$ imply
$\widehat{\pi}_{k}\lesssim\pi_{k}+\sqrt{\log(n)\over nL}\lesssim{1\over L},\quad\forall\ k\in[L],$
and after we invoke Lemma 15 twice with $v=v_{1},u={M}^{\top}v_{2}$ and $v=v_{2},u={M}^{\top}v_{1}$, respectively, we get that
$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}v_{2}\lesssim\|{M}^{\top}v_{1}\|_{2}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{M}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n},$ (C.12)
$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}v_{2}\lesssim\|{M}^{\top}v_{2}\|_{2}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{M}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}$ (C.13)
hold with probability at least $1-4e^{-t}$. Collecting all three bounds completes the proof. ∎
As an immediate application of Lemma 18, we have the following bounds on the sup-norm and operator norm of the matrix $B^{\top}(\widehat{\Sigma}-\Sigma)B$.
Lemma 19.
Under conditions of Lemma 18, assume $\Delta_{\infty}\gtrsim 1$. On the event $\mathcal{E}_{\pi}$, the following holds with probability at least $1-n^{-2}$,
$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\infty}\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL},$
$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\lesssim{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}.$
Proof.
For the sup-norm bound, it suffices to bound from above $|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ for any $k\in[L]$. Invoking Lemma 18 with $v_{1}=v_{2}=B_{k}$ and $t=\log(n)$ together with (C.5), (C.6) and (C.7) of Lemma 14 gives that, with probability at least $1-n^{-2}$,
$|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ $\lesssim B_{k}^{\top}\Sigma_{w}B_{k}\sqrt{\log(n)\over n}+\|{M}^{\top}B_{k}\|_{\infty}\sqrt{B_{k}^{\top}{M}{M}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
$\quad+\|{M}^{\top}B_{k}\|_{\infty}^{2}{\log(n)\over n}+2\sqrt{B_{k}^{\top}\Sigma_{w}B_{k}B_{k}^{\top}{M}{M}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
$\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)\sqrt{\log(n)\over nL^{2}}+\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL}+\left(1\wedge{\Delta_{\infty}\over L}\right)^{2}{\log(n)\over n}$
$\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL}.$
The last step uses $\Delta_{\infty}\gtrsim 1$ to collect terms. This proves the first claim.
Regarding the operator norm bound, a standard discretization argument gives
$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\leq 3\max_{u\in\mathcal{N}_{L}(1/2)}u^{\top}B^{\top}(\widehat{\Sigma}-\Sigma)Bu$
where $\mathcal{N}_{L}(1/3)$ denotes the $(1/3)$-epsilon net of $\{u\in\mathbb{R}^{L}:\|u\|_{2}=1\}$. Note that $|\mathcal{N}_{L}(1/3)|\leq 7^{L}$. Then invoking Lemma 18 with $t=CL\log(n)$ and using
$\|B^{\top}\Sigma_{w}B\|_{\rm op}\leq{\Delta_{\rm op}\over L(L+\Delta_{\rm op})},$
$\|B^{\top}{M}\|_{\rm op}\leq{\Delta_{\rm op}\over L+\Delta_{\rm op}},$
$\|{M}^{\top}Bu\|_{\infty}\leq\sqrt{\Delta_{\infty}\Delta_{\rm op}\over L(L+\Delta_{\rm op})}$
deduced from Lemma 14 give that, with probability at least $1-8n^{-CL+L\log(7)}\geq 1-n^{-L}$,
$u^{\top}B^{\top}(\widehat{\Sigma}-\Sigma)Bu$ $\lesssim u^{\top}B^{\top}\Sigma_{w}Bu\sqrt{L\log(n)\over n}+\|{M}^{\top}Bu\|_{\infty}\sqrt{u^{\top}B^{\top}{M}{M}^{\top}Bu}\sqrt{\log(n)\over n}$
$\quad+\|{M}^{\top}Bu\|_{\infty}^{2}{L\log(n)\over n}+2\sqrt{u^{\top}B^{\top}\Sigma_{w}Buu^{\top}B^{\top}{M}{M}^{\top}Bu}\sqrt{\log(n)\over n}$
$\lesssim{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}$
holds uniformly over $u\in\mathcal{N}_{L}(1/3)$. We also use $\Delta_{\infty}\gtrsim 1$ to simplify expressions in the last step above. This completes the proof. ∎
The following lemma provides upper bounds of the operator norm of $(\widehat{\Sigma}-\Sigma)$.
Lemma 20.
Under model (1.1) and 3, assume $(p+L)\log(n)\leq n$. Then with probability at least $1-n^{-2}$, the following holds uniformly over $v\in\mathcal{S}^{p}$
$v^{\top}(\widehat{\Sigma}-\Sigma)v~{}\lesssim~{}v^{\top}\Sigma v\left(\sqrt{p\log(n)\over n}+\sqrt{L\log(n)\over n}\right).$
As a result, with the same probability, we have
$\lambda_{p}(\widehat{\Sigma})\geq{1\over 2}\lambda_{p}(\Sigma).$
Proof.
The proof of the first statement follows from the same arguments of proving Lemma 12 and is thus omitted.
For the second statement, for any $v\in\mathcal{S}^{p}$, we have
$v^{\top}\widehat{\Sigma}v=v^{\top}\Sigma^{1/2}\left(\Sigma^{-1/2}\widehat{\Sigma}\Sigma^{-1/2}\right)\Sigma^{1/2}v\geq v^{\top}\Sigma v~{}\lambda_{p}\left(\Sigma^{-1/2}\widehat{\Sigma}\Sigma^{-1/2}\right).$
Since, by Weyl’s inequality and on the event that the first result holds,
$\lambda_{p}\left(\Sigma^{-1/2}\widehat{\Sigma}\Sigma^{-1/2}\right)\geq 1-\|\Sigma^{-1/2}(\widehat{\Sigma}-\Sigma)\Sigma^{-1/2}\|_{\rm op}\geq 1-\left(\sqrt{p\log(n)\over n}+\sqrt{L\log(n)\over n}\right)$
we conclude
$v^{\top}\widehat{\Sigma}v\geq{1\over 2}v^{\top}\Sigma v$
uniformly over $\mathcal{S}^{p}$, completing the proof. ∎
Appendix D Auxiliary lemmas
The following lemmas are proved in Bing and Wegkamp (2023).
Lemma 21.
Assume $\pi_{\min}\geq 2\log n/n$ for some sufficiently large constant $C$. Then, for any $k\in[L]$
$\mathbb{P}\left\{|\widehat{\pi}_{k}-\pi_{k}|<\sqrt{16\pi_{k}(1-\pi_{k})\log n\over n}\right\}\geq 1-n^{-2}.$
Furthermore, if $\pi_{\min}\geq C\log n/n$ for some sufficiently large constant $C$, then
$\mathbb{P}\left\{c\pi_{k}\leq\widehat{\pi}_{k}\leq c^{\prime}\pi_{k}\right\}\geq 1-n^{-2}.$
Lemma 22.
Suppose that model (1.1) holds. For any deterministic vector $v\in\mathbb{R}^{p}$ and $k\in[L]$, for all $t>0$,
$\mathbb{P}\left\{\left|v^{\top}(\widehat{\mu}_{k}-\mu_{k})\right|\geq t\sqrt{v^{\top}\Sigma_{w}v\over n_{k}}\right\}\leq 2e^{-{t^{2}/2}}.$